### v1

# Why Can’t I Open My Drawer? Mitigating Object-Driven Shortcuts in Zero-Shot Compositional Action Recognition

Abstract
We study Compositional Video Understanding (CVU), where models must recognize verbs and objects and compose them to generalize to unseen combinations. We find that existing Zero-Shot Compositional Action Recognition (ZS-CAR) models fail primarily due to an overlooked failure mode: object-driven verb shortcuts. Through systematic analysis, we show that this behavior arises from two intertwined factors—severe sparsity and skewness of compositional supervision, and the asymmetric learning difficulty between verbs and objects. As training progresses, the existing ZS-CAR model increasingly ignores visual evidence and overfits to co-occurrence statistics. Consequently, the existing model does not gain the benefit of compositional recognition in unseen verb-object compositions. To address this, we propose RCORE, a simple and effective framework that enforces temporally grounded verb learning. RCORE introduces (i) a composition-aware augmentation that diversifies verb–object combinations without corrupting motion cues, and (ii) a temporal order regularization loss that penalizes shortcut behaviors by explicitly modeling temporal structure. Across two benchmarks—Sth-com and our newly constructed EK100-com—RCORE significantly improves unseen composition accuracy, reduces reliance on co-occurrence bias, and achieves consistently positive compositional gaps. Our findings reveal object-driven shortcuts as a critical limiting factor in ZS-CAR and demonstrate that addressing them is essential for robust compositional video understanding. The code is available at https://github.com/KHU-VLL/RCORE.
(a) Co-occurrence bias
(b) Asymmetric learning difficulty
Figure 1: Why object-driven shortcuts emerge in compositional video understanding? (a) Co-occurrence bias. Datasets are intrinsically sparse and highly skewed in their verb–object combinations, creating strong co-occurrence priors. Models exploit these priors as a shortcut: once the object is recognized, the model often predicts the most frequent verb paired with it, ignoring temporal evidence. (b) Asymmetric learning difficulty. Objects are visually explicit and easy to recognize from a single frame, whereas verbs require multi-frame temporal reasoning. This imbalance makes object features dominate training signals. Together, these two factors drive models to adopt object-driven shortcuts, hindering genuine compositional generalization. Best viewed with zoom and color.
11footnotetext: This work was done during an internship at NAVER Cloud.22footnotetext: Corresponding author.
1 Introduction
Many human actions can be decomposed into two semantic primitives—verbs and objects—and robust video understanding requires recognizing each component and reasoning over their interaction as a composition [12, 22, 7, 8, 2]. Zero-Shot Compositional Action Recognition (ZS-CAR) formalizes this goal: a model must correctly identify unseen verb–object combinations despite only being exposed to a sparse subset of compositions during training [16].
However, existing ZS-CAR models exhibit a characteristic and persistent failure: they mispredict verbs in unseen compositions by relying on object-driven shortcuts. In other words, rather than learning verb semantics from temporal evidence, models exploit objects as the dominant cue whenever such shortcuts correlate with training labels. To the best of our knowledge, this failure mode has not been explicitly identified or systematically addressed in prior ZS-CAR approaches [16, 14]. We show that this shortcut failure mode arises from two intertwined issues of the ZS-CAR setting: (i) Sparsity and skewness of compositional supervision. The combinatorial nature of verb–object pairs yields extreme under-coverage of the compositional space, e.g. only $\sim 14.0\%$ label coverage ratio in the Something-Else [22] and EPIC-KITCHENS-100 [7] datasets. This induces strong co-occurrence statistics to the deep learning model as depicted in Figure 1 (a). (ii) Asymmetric learning difficulty between verbs and objects. Objects are recognizable from single frames, while verbs require multi-frame temporal reasoning, as shown in Figure 1 (b). This imbalance naturally encourages models to treat objects as easy predictive shortcuts [25, 17, 32, 11, 30], especially under sparse data regimes.
Our analysis reveals that the existing ZS-CAR model amplifies these issues during training. They increasingly bias predictions toward frequently co-occurring compositions, exhibit growing false-seen and false-co-occurrence prediction ratios, and ultimately yield negative compositional gaps—indicating that they fail to extract any benefit from compositional reasoning beyond independent verb and object recognition. (See Section 3 for details.) These observations call into question the robustness of current ZS-CAR approaches and motivate the need for techniques that explicitly target shortcut mitigation.
We propose RCORE (Robust COmpositional REpresentations), a simple and effective framework designed to mitigate object-driven shortcut and promote temporally grounded verb representations. RCORE consists of two components: (i) a composition-aware augmentation that synthesizes plausible unseen verb–object combinations while preserving the temporal structure of the original video; and (ii) a temporal order regularization objective that penalizes reliance on static cues by contrasting original and perturbed temporal representations. Together, these components address the two root causes of failure in ZS-CAR: compositional sparsity and asymmetric learning difficulty.
Across both datasets, RCORE significantly strengthens verb recognition, reduces co-occurrence bias, produces positive compositional gaps, and improves unseen composition accuracy over strong conditional baselines. Our findings indicate that the primary obstacle in compositional video understanding is not model capacity but model behavior: ZS-CAR models fail because they resort to object-driven shortcuts. By directly tackling this failure mode, RCORE provides a simple path toward more reliable and generalizable ZS-CAR.
Figure 2: Controlled experiments demonstrate object-driven shortcut learning in ZS-CAR. We empirically identify a key failure mode in ZS-CAR—object-driven shortcuts. (a) Objects are easier to learn than verbs. We train a randomly initialized ViT [10] on a balanced 10 $\times$ 10 verb-object subset from Sth-com [16]. The learning curves show that object accuracy increases much faster than verb accuracy, indicating that objects are easier to learn. (b) Under a strong co-occurrence bias, models exploit object cues as shortcuts, failing to learn robust verb representations. We prepare a perfectly biased training set and two kinds of test splits utilizing the Sth-com [16] dataset: seen compositions (bias-aligned) and unseen compositions (bias-conflict). Both the randomly initialized ViT [10] and pretrained CLIP [28] exhibit a consistent trend: while unseen object accuracy increases alongside seen performance, unseen verb accuracy concurrently drops below random chance. Best viewed with zoom and color.
In this work, we make the following major contributions:
  • Identifying the core failure mode of ZS-CAR. We show that existing models rely on object-driven verb shortcuts, caused by the sparsity of compositional supervision and the asymmetric learning difficulty between verbs and objects—an issue not explicitly recognized in prior work.
  • Diagnosing limitations of the ZS-CAR evaluation protocol. We reveal that common practices, such as test-set bias calibration, obscure true compositional generalization. To address this, we adopt an open-world unbiased evaluation setting and introduce the compositional gap metric to measure the benefit of compositional reasoning.
  • A simple yet effective framework that mitigates shortcut learning. We propose RCORE, which comprises a composition-aware augmentation and a temporal order regularization loss (TORC) to promote temporally grounded verb representations under compositional sparsity. Through experiments on Sth-com and our new EK100-com dataset, we demonstrate that RCORE shows consistently favorable performance improvements across diverse evaluation scenarios, metrics, and datasets.
2 Related Work
Zero-shot compositional action recognition. Zero-Shot Compositional Action Recognition (ZS-CAR) [16] aims to generalize to novel verb–object combinations not seen during training. Unlike zero-shot (or open-set) action recognition [5, 4, 40], ZS-CAR retains the same verb and object classes across train and test splits and evaluates a model’s ability to compose known primitives. It addresses the performance degradation issue on novel unseen combinations of known concepts, which arises from the model’s lack of compositional generalization ability. Conditional learning [16, 13] has been effective for estimating joint verb–object probabilities, while other work [14] pursues improved feature disentanglement. However, as our analysis reveals in Section 3 and Section 5.2, these approaches remain susceptible to object-driven shortcut behavior and progressively overfit co-occurrence statistics.
Compositional zero-shot Learning in image understanding.
Compositional Zero-Shot Learning (CZSL) has been widely studied in the image domain [23, 24, 21, 34, 36, 26]. CZSL enables models to recognize novel combinations of previously learned primitive concepts, such as objects and their states. Prior works have approached this problem through feature disentanglement [29, 42], modular architectures [27], and CLIP-based prompt learning [26, 18]. Unlike image-based CZSL, ZS-CAR requires modeling temporal dynamics for verbs, making the problem fundamentally different. Our work focuses specifically on the video domain and highlights a failure mode—object-driven verb shortcuts—that does not arise in image-only CZSL.
Shortcut learning.
Shortcut learning is a pervasive issue in visual recognition. Several studies show that models tend to rely on easy-example biases [35, 15, 30] or easier-to-learn cues—such as color or texture—when these cues correlate with labels [25, 11]. A line of work attributes this behavior to co-occurrence statistics in the training data [1, 32]. In video understanding, similar shortcut behaviors occur when action models rely on backgrounds rather than motion [17, 6, 33, 9, 2]. Building on these observations, we argue that ZS-CAR introduces an additional structural imbalance: verbs require temporal reasoning, whereas objects can be recognized from a single frame. This learning asymmetry amplifies shortcut tendencies and leads models to predict verbs primarily from object cues—a failure mode we diagnose and directly address in this work.
3 Diagnosis: Why ZS-CAR Models Fail?
3.1 Zero-Shot Compositional Action Recognition
We study Zero-Shot Compositional Action Recognition (ZS-CAR) [16], where the goal is to recognize verb–object compositions that might never appear in the training set. Let $\mathbb{Y}^{V}$ and $\mathbb{Y}^{O}$ denote the sets of verb and object labels, and let each action label be a pair $\mathbf{y}=(\mathbf{y}^{V},\mathbf{y}^{O})\in\mathbb{Y}^{V}\times\mathbb{Y}^{O}$. The training set is $\mathbb{D}_{\text{train}}=\{(\mathbf{X}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$, and its set of observed (seen) compositions is $\mathbb{Y}_{\text{seen}}=\{\mathbf{y}_{i}\mid(\mathbf{X}_{i},\mathbf{y}_{i})\in\mathbb{D}_{\text{train}}\}$. The space of unseen compositions is $\mathbb{Y}_{\text{unseen}}=(\mathbb{Y}^{V}\times\mathbb{Y}^{O})\setminus\mathbb{Y}_{\text{seen}}$. We define the validation and test sets analogously. The validation and test sets contain both seen and unseen compositions, enabling evaluation of compositional generalization. ZS-CAR therefore measures whether a model can correctly compose familiar verbs and objects into novel pairs at test time.
3.2 Diagnostic Metrics
For better understanding the underlying behavior of ZS-CAR models, we present two diagnostic metrics: (1) training bias metrics to measure a model’s overfitting and (2) Compositional Gap to assess the benefits of compositional modeling.
Training bias metrics
To measure how the model overfits to the training set statistics, we introduce two types of misclassification ratios: False Seen Prediction and False Co-occurrence Prediction ratios. False Seen Prediction (FSP) ratio denotes the ratio of unseen composition samples incorrectly predicted to seen composition labels. Out of these mispredicted samples, False Co-occurrence Prediction (FCP) ratio is likewise defined as the proportion of those assigned to frequently co-occurring labels. Here, a composition label is considered ‘frequently co-occurring’ if its occurrence is higher than one standard deviation above the mean frequency in the training data.
Compositional Gap
We introduce Compositional Gap to examine whether a model genuinely benefits from understanding videos in a compositional manner. Compositional Gap ($\Delta_{\text{CG}}$) is defined as
$\Delta_{\text{CG}}=\mathrm{Acc}^{C}-(\mathrm{Acc}^{V}\times\mathrm{Acc}^{O}),$ (1)
where $\text{Acc}^{C}$ is the composition accuracy (%) computed from the logits over the composition space, whereas $\text{Acc}^{V}$ and $\text{Acc}^{O}$ are the verb and object accuracies (%) derived from the logits over their individual component spaces. A positive $\Delta_{\text{CG}}$ indicates that the model truly gains an advantage from its compositional architecture.
Figure 3: Learning curve of the SOTA model with our diagnostic metrics. We plot the learning curve of C2C [16] trained on Sth-com [16]. We measure the False Seen Prediction (FSP) and False Co-occurrence Prediction (FCP) ratios, and observe that the seen–unseen accuracy gap ($\Delta_{SU}$) correlates strongly with both metrics. There observations indicate that the current SOTA model exhibits overfitting to seen compositions. Best viewed with zoom and color.
Figure 4: Overview of RCORE. (a) Overview of our proposed RCORE framework. (b) VOCAMix synthesizes plausible yet unseen verb–object compositions while preserving the temporal structure of the primary video. (c) TORC penalizes alignment between original and temporally perturbed feature vectors, enforcing explicit temporal order modeling and reducing object-driven shortcuts.
3.3 Empirical Diagnosis of Object-driven Shortcuts in ZS-CAR
In this section, we systematically diagnose the key failure mode in ZS-CAR—object-driven shortcuts. We demonstrate that these shortcuts prevent ZS-CAR models from learning robust verb representations, consequently leading to poor generalization to unseen compositions. We present additional evidence of object-driven shortcut learning in ZS-CAR models in Section 5 and the supplementary material.
Objects are easier to learn than verbs.
To investigate the learning difficulty of verbs and objects, we train a randomly initialized model on a subset of Sth-com [16]. Based on the observation that the easier-to-learn cue is captured earlier and exhibits higher accuracy [25], we can analyze the learning difficulty of each component. In Figure 2 (a), the results show a clear trend: the model learns objects faster and consistently achieves higher accuracy for objects than for verbs. While intuitively expected, these results empirically verify that objects are inherently easier to learn than verbs in compositional action recognition.
Object-driven shortcuts in ZS-CAR
Under strong co-occurrence priors, models choose shortcuts emerged from easier-to-learn cues, rather than learn robust representations [25, 30, 11, 3, 17, 6]. We argue this shortcut learning behavior accelerates in ZS-CAR due to the asymmetric learning difficulty. To verify this, inspired by prior work [30], we train models on a perfectly biased training set and test on two distinct splits—bias-aligned seen compositions and bias-conflict unseen compositions— utilizing Sth-com [16]. As shown in Figure 2 (b), in the both cases of the randomly initialized ViT [10] and CLIP [28], the object accuracies for unseen compositions increased as the seen composition accuracies also increased, while the verb accuracies for unseen compositions drop under the random accuracy. Paradoxically, CLIP [16], well-known for its superior representational power, exhibits the highest object accuracy (94.0%) and the lowest verb accuracy (1.9%) concurrently on unseen compositions. These results demonstrate models tend to fall into object-driven shortcuts to exploit co-occurrence priors in the training data, resulting in poor verb representations.
Object-driven shortcuts directly harm unseen generalization.
Next, we investigate whether the state-of-the-art (SOTA) model also fails to generalize to unseen compositions as a consequence of these object-driven shortcuts. As shown in Figure 3, we trace the validation accuracies along with FSP and FCP ratios—metrics designed to quantify the model’s reliance on the training statistics—of C2C [16] throughout the training progress on Sth-com [16]. The clear correlation between increased FSP/FCP ratios and widen seen-unseen accuracy gap indicates that the model overfits to the training statistics.
Table 1: The SOTA model exhibits a negative Compositional Gap on unseen compositions. We measure Compositional Gap ($\Delta_{\mathrm{CG}}$) of the SOTA ZS-CAR model. We employ C2C [16] trained on Sth-com [16] with 30 epochs and evaluate the model in open-world setting on the Sth-com [16] test set. The model shows a negative $\Delta_{\mathrm{CG}}$ on unseen compositions, indicating its composition accuracy falls below the level expected from independent verb–object predictions.
Split $\mathrm{Acc}^{V}$ $\mathrm{Acc}^{O}$ $\mathrm{Acc}^{V}\times\mathrm{Acc}^{O}$ $\mathrm{Acc}^{C}$ $\Delta_{\text{CG}}$
Seen 63.60 67.72 43.07 46.31 +3.24
Unseen 54.36 56.10 30.50 30.08 -0.42
Compositional Gap reveals the SOTA model fails to benefit from compositional modeling.
We evaluate the Compositional Gap ($\Delta_{\text{CG}}$) performance of current SOTA ZS-CAR model [16]. As shown in Table 1, the model shows a negative $\Delta_{\text{CG}}$ on unseen compositions in Sth-com [16]. This demonstrates that, without addressing these shortcuts, the SOTA model gains no advantage from compositional modeling compared to independent modeling for unseen compositions.
4 RCORE
We introduce RCORE, a learning framework that mitigates object-driven shortcut learning in zero-shot compositional action recognition. RCORE improves compositional generalization by strengthening verb–object representations under sparse supervision through two complementary components: (i) VOCAMix, a composition-aware augmentation that expands the variety of verb–object combinations, and (ii) TORC, a temporal order regularization loss that promotes temporally grounded verb learning over static object cues. Following prior work [16], we use AIM [37], a CLIP-based [28] parameter-efficient spatio-temporal encoder, as our backbone. Figure 4 (a) gives an overview of RCORE.
Feature extraction.
Given an augmented video $\hat{\mathbf{X}}$ with $T$ frames, the backbone encoder outputs frame-level features $\mathbf{f}_{t}\in\mathbb{R}^{D}$. We then transform them into verb features $\mathbf{F}^{V}\in\mathbb{R}^{T\times D}$ and an object features $\mathbf{f}^{O}\in\mathbb{R}^{D}$ using dedicated verb and object encoders. We obtain text embeddings for verbs and objects, $\mathbf{e}^{V}$ and $\mathbf{e}^{O}$, by feeding class-specific prompts (e.g., “a photo of {verb/object}) into the text encoder.
4.1 Composition-Aware Augmentation
To alleviate the sparsity of compositional training data, we propose Verb–Object Composition-Aware Augmentation (VOCAMix), which synthesizes plausible but unseen combinations while preserving the temporal structure of the primary video.
As shown in Figure 4 (b), VOCAMix replaces the motion region of a video with a static object region from another sample in the mini-batch. Given a primary video $\mathbf{X}_{i}$ with label $(\mathbf{y}_{i}^{V},\mathbf{y}_{i}^{O})$ and a randomly selected video $\mathbf{X}_{j}$ with object label $\mathbf{y}_{j}^{O}$ ($i\neq j$), we construct an augmented video $\hat{\mathbf{X}}_{i}$ and a soft object label:
$h(\hat{\mathbf{X}}_{i}(k,:))$ $=(1-\lambda)\,h(\mathbf{X}_{i}(k,:))+\lambda\,h(\mathbf{X}_{j}(\lfloor T/2\rfloor,:)),$ (2)
$\tilde{\mathbf{y}}^{O}_{i}$ $=(1-\lambda)\,\mathbf{y}_{i}^{O}+\lambda\,\mathbf{y}_{j}^{O},$ (3)
where $h(\cdot)$ is a high motion-region estimation function, $k\in\{1,\ldots,T\}$ is the frame index, and $\lambda\in[0,1]$ controls the strength of object injection. We adopt a simple, learning-free estimator [9] for $h(\cdot)$. We keep the verb label $\mathbf{y}_{i}^{V}$ unchanged to preserve temporal semantics. By injecting a consistent static object across the temporal axis, VOCAMix preventing disruption of temporal cues [33] while expanding compositional diversity. while synthesizing novel compositions.
4.2 Temporal Order Regularization Loss
To counter object-driven shortcuts in verb learning, we introduce the Temporal Order Regularization for Composition (TORC) loss. TORC loss forces the model to rely on temporal structure: it penalizes alignment with temporally incorrect feature sequences and suppresses confident verb predictions when temporal ordering is corrupted. This regularization steers the model away from static object cues and strengthens temporally grounded verb representations.
Figure 5: Analysis on the effects of RCORE on the Sth-com [16] dataset. (a) RCORE prevents the False Co-occurrence Prediction (FCP) ratio from increasing during training, whereas the baseline shows a clear rise in FCP. As a result, RCORE consistently maintains a smaller seen–unseen accuracy gap ($\Delta_{SU}$) throughout training. (b) The cosine similarity between the original and reversed verb features becomes strongly negative for RCORE as training progresses, indicating improved temporal discriminative capability. In contrast, the baseline maintains a high similarity (0.91), revealing limited temporal sensitivity. (c) On the Temporal subset, RCORE exhibits a substantially larger performance gap between original and temporally shuffled features compared to the baseline, demonstrating that RCORE learns verb representations that depend on temporal dynamics rather than static object cues. Best viewed with zoom and color.
Temporal perturbation.
Given frame-level features $\mathbf{F}=(\mathbf{f}_{1},\dots,\mathbf{f}_{T})$, we construct two perturbed variants: (i) a reversed sequence $\mathbf{F}_{\text{rev}}=(\mathbf{f}_{T},\dots,\mathbf{f}_{1})$, and (ii) a temporally shuffled sequence $\mathbf{F}_{\text{shuffled}}=\pi(\mathbf{F})$, where $\pi$ is a random permutation sampled from $\mathcal{P}$. We then feed $\mathbf{F}_{\text{rev}}$ and $\mathbf{F}_{\text{shuffled}}$ into the verb encoder to obtain perturbed verb features $\mathbf{f}^{V}_{\text{rev}}\in\mathbb{R}^{D}$ and $\mathbf{f}^{V}_{\text{shuffled}}\in\mathbb{R}^{D}$.
TORC loss.
We first push the model to distinguish forward and reversed temporal semantics. We minimize the cosine similarity between the original verb feature $\mathbf{f}^{V}$ and the reversed feature $\mathbf{f}^{V}_{\text{rev}}$:
$L_{\cos}=\cos(\mathbf{f}^{V},\mathbf{f}^{V}_{\text{rev}}),\vskip-5.0pt$ (4)
where $\cos(\mathbf{a},\mathbf{b})=\frac{\mathbf{a}^{\top}\mathbf{b}}{\|\mathbf{a}\|\|\mathbf{b}\|}$. Temporal reversal often flips action meaning (e.g., opening vs. closing), yet the state-of-the-art model [16] yields a high cosine similarity ($0.91$; see Figure 5), indicating weak temporal discrimination.
Next, we regularize the model to avoid confident verb predictions when temporal structure is disrupted. Given the set of verb text embeddings $\mathbb{E}^{V}=\{\mathbf{e}^{V}_{i}\}_{i=1}^{|\mathbb{Y}^{V}|}\in\mathbb{R}^{|\mathbb{Y}^{V}|\times D}$, we define the negative entropy loss $L_{\text{ent}}$ as:
$L_{\text{ent}}$ $=\sum_{i=1}^{|\mathbb{Y}^{V}|}p_{i}\log p_{i},\vskip-5.0pt$ (5)
where $p_{i}$ denotes the predicted probability of the $i$-th verb class based on the cosine similarity between visual and textual features, and $\tau$ is a temperature parameter. Maximizing entropy prevents the model from inferring verbs from static spatial cues alone, enforcing reliance on temporal evidence. We define the TORC loss as: $L_{\text{TORC}}=L_{\cos}+L_{\text{ent}}$.
Component loss.
We compute verb/object logits using cosine similarity between visual features and their corresponding text features. We apply standard cross-entropy losses for verb and object prediction: $L_{\text{com}}=L_{V}+L_{O}$.
Composition loss.
Following prior work [16], we combine marginal scores $(\mathbf{p}_{V},\mathbf{p}_{O})$ with conditional logits $(\mathbf{p}_{V|O},\mathbf{p}_{O|V})$ to obtain the composition prediction $\hat{\mathbf{y}}$ as $\hat{\mathbf{y}}=\texttt{Softmax}(\mathbf{p}_{O}\mathbf{p}_{V|O}+\mathbf{p}_{V}\mathbf{p}_{O|V})$. We employ the cross-entropy loss: $L_{\text{comp}}=L_{\text{CE}}(\hat{\mathbf{y}},\mathbf{y})$.
Margin loss: mitigating co-occurrence bias.
Highly frequent verb–object pairs often dominate predictions, even for unseen compositions. To counter this, we introduce a margin loss that penalizes frequent-but-incorrect compositions whose logits approach the ground-truth logit:
$L_{M}=\sum_{c\in\mathcal{C}_{\text{margin}}}\max\left(0,\log\hat{\mathbf{y}}(c)-\log\hat{\mathbf{y}}(g)+m\right),$ (6)
where $g\in\{1,\dots,|\mathbb{Y}^{V}|\times|\mathbb{Y}^{O}|\}$ denotes the ground-truth index, and $\mathcal{C}_{\text{margin}}$ is the intersection of top-$K$ incorrect predictions and the set of frequently co-occurring compositions. This term explicitly discourages reliance on dataset co-occurrence priors.
Total loss.
We train RCORE with the combined objective $L_{\text{total}}$ defined as $L_{\text{total}}=\alpha L_{\text{com}}+\beta L_{\text{comp}}+\gamma L_{\text{TORC}}+\delta L_{M}$ , where $\alpha$, $\beta$, $\gamma$, and $\delta$ control the contribution of each term.
5 Experimental Results
In this section, we conduct rigorous experiments to address the following research questions: (1) Do conventional ZS-CAR evaluation protocols faithfully reveal true model behavior? If not, what is a more appropriate protocol for exposing real generalization performance? (Section 5.1) (2) Does the proposed compositional gap (CG) metric reveal whether a model actually benefits from compositional understanding beyond independent verb/object predictions? (Section 5.2 and Section 5.3) (3) Does RCORE mitigate the two key failure modes—co-occurrence overfitting and object-driven verb shortcuts—identified in our diagnosis? (Section 5.2) (4) Do these improvements translate to consistent gains across datasets? (Section 5.3) (5) How does each proposed component contribute to robust compositional learning? (Section 5.4) We describe the experimental setup and then report results aligned with these questions.
Table 2: Sth-com [16] results. We show the top-1 verb, object, and composition classification accuracy (%). All models use CLIP-B/16 [28] as the backbone. We report performance on both seen and unseen compositions, along with their harmonic mean (H.M.).
Model Bias Calib. Verb Object Composition
Seen Unseen H.M. Seen Unseen H.M. Open-world Closed-world
Seen $(\Delta_{\mathrm{CG}}{})$ Unseen $(\Delta_{\mathrm{CG}}{})$ H.M. H.M. AUC
AIM [37] $\times$ 54.99 39.19 45.76 67.19 53.43 59.53 40.24 (+3.29) 18.50 (-2.44) 25.35 - -
C2C w/o conditioning $\times$ 62.53 48.67 54.74 66.90 52.66 58.93 44.95 (+3.12) 23.94 (-1.69) 25.35 - -
C2C [16] $\times$ 63.60 54.36 58.62 67.72 56.10 61.36 46.31 (+3.24) 30.08 (-0.42) 36.47 - -
RCORE $\times$ 65.93 57.29 61.31 67.82 55.56 61.08 48.56 (+3.85) 32.23 (+0.40) 38.75 - -
AIM [37] 50.32 43.19 46.48 64.02 54.78 59.04 33.40 (+1.19) 24.60 (+0.94) 28.33 34.44 16.99
C2C w/o conditioning 57.08 50.31 53.48 65.07 53.32 58.61 39.32 (+2.18) 26.70 (+1.89) 31.80 37.62 19.74
C2C [16] 60.87 56.81 58.77 65.83 57.37 58.93 42.27 (+2.20) 34.22 (+1.30) 37.82 42.75 23.77
Jung et al. [14] - - - - - - - - - 42.60 23.50
RCORE 63.57 58.77 61.08 66.73 56.78 61.35 45.34 (+2.91) 35.26 (+1.89) 39.67 45.45 26.52
C2C (Baseline)
RCORE (Ours)
Figure 6: RCORE mitigates object-driven shortcuts in verb learning. We visualize confusion matrices for six representative verbs to compare the ability of RCORE and C2C to distinguish opposite temporal semantics on unseen compositions of the Sth-com [16] test set. All values in the confusion matrices are normalized frequencies across the entire verb classes in the dataset.
5.1 Experimental setup
Limitations of conventional ZS-CAR evaluation.
Existing ZS-CAR works [16, 13, 14] rely on the closed-world protocol, where inference is restricted to a small set of compositions that appear in the validation or test split. This protocol hides a model’s incorrect tendency to over-predict seen compositions and obscures co-occurrence-driven behaviors. A more serious methodological flaw is the use of test-set–tuned bias calibration. Selecting the bias term with access to test labels systematically inflates unseen accuracy and prevents meaningful comparison across methods.
Our evaluation setup.
To expose genuine generalization, we adopt an open-world, unbiased protocol by default: (i) inference spans the full space $\mathbb{Y}^{V}\times\mathbb{Y}^{O}$, and (ii) no test labels are used to tune biases. We also report: (a) open-world biased results using a realistic validation set-tuned bias, and (b) closed-world H.M. and AUC for comparison with existing methods. All results in Section 3 also follow the default open-world, unbiased protocol.
Datasets. We evaluate on two ZS-CAR benchmarks: Sth-com [16] and our curated EK100-com. Sth-com is derived from Something-Something V2 [12] and Something-Else [22], containing 79K videos with 161 verbs and 248 objects. Following prior work [39, 31], we also evaluate on the Temporal subset consisting of verbs requiring temporal reasoning. We additionally introduce the EPIC-KITCHENS-100-composition (EK100-com) benchmark, constructed by repurposing EPIC-KITCHENS-100 [7] following the Sth-com protocol. EK100-com includes 71K egocentric videos, 82 verbs, and 228 objects. Full construction details are provided in the Section A.
Evaluation metrics.
We report top-1 composition, verb, and object accuracy on seen and unseen compositions, along with their harmonic mean (H.M.). We also measure the compositional gap $\Delta_{\mathrm{CG}}$ (1) which indicates whether a model benefits from compositional reasoning beyond independent verb/object predictions. (see Section 3) A higher $\Delta_{\mathrm{CG}}$ reflects stronger compositional synergy.
Baselines.
We primarily compare RCORE against C2C [16], the current SOTA ZS-CAR method. To isolate the effect of conditional learning, we also evaluate two independent baselines. AIM [37] is the CLIP [28]-based spatio-temporal adapter model used as our backbone. C2C w/o conditioning augments AIM with all architectural components of C2C except the conditional modeling itself. We additionally include closed-world results of a recent ZS-CAR approach [14].
Table 3: EK100-com results. We show the top-1 verb, object, and composition classification accuracy (%). All models use CLIP-B/16 [28] as the backbone. We report performance on both seen and unseen compositions, along with their harmonic mean (H.M.).
Model Bias Calib. Verb Object Open-world Composition
Seen Unseen H.M. Seen Unseen H.M. Seen $(\Delta_{\mathrm{CG}}{})$ Unseen $(\Delta_{\mathrm{CG}}{})$ H.M.
AIM [37] $\times$ 60.04 39.78 47.85 56.04 46.33 50.72 38.60 (+4.95) 14.74 (-3.69) 21.34
C2C [16] $\times$ 66.19 49.71 56.78 56.61 47.48 51.64 42.72 (+5.25) 22.38 (-1.22) 29.38
RCORE $\times$ 66.07 52.07 58.24 57.76 47.11 51.89 43.86 (+5.70) 24.24 (-0.29) 31.23
AIM [37] 52.42 49.54 50.94 50.70 51.05 50.87 28.00 (+1.42) 32.07 (+6.78) 29.89
C2C [16] 62.21 57.74 59.89 51.66 54.16 52.88 36.05 (+3.91) 37.96 (+6.69) 36.98
RCORE 62.47 59.77 61.09 53.42 54.79 54.10 36.53 (+3.16) 39.78 (+7.03) 38.09
5.2 Analysis
In this section, we validate that our RCORE framework effectively addresses the two key challenges in ZS-CAR—(i) reliance on co-occurrence statistics and (ii) object-driven shortcuts in verb learning—and achieves improved compositional representations compared to the baseline [16].
Overcoming reliance on co-occurrence statistics.
In Figure 5 (a), we present the False Co-occurrence Prediction (FCP) curves of RCORE and the baseline [16] on Sth-com. While the baseline’s FCP ratio increases substantially during training (11% $\rightarrow$ 17%), the FCP of RCORE decreases (9% $\rightarrow$ 8%), indicating that RCORE is markedly less dependent on co-occurrence statistics.
Addressing object-driven shortcuts.
In Figure 5 (b), we compare the cosine similarity between the original verb features $\mathbf{f}^{V}$ and the reversed verb features $\mathbf{f}^{V}_{\text{rev}}$ for RCORE and the baseline [16] over the course of training. While the baseline [16] maintains a high cosine similarity ($0.91$) due to its reliance on object-driven shortcuts, RCORE drives this value to a strongly negative (0.73-0.73), successfully distinguishing opposite temporal semantics such as ’opening’ vs.’closing’.
In Figure 6, we further visualize this behavior through verb confusion matrices on the Sth-com [16] test set. RCORE cleanly separates pairs of opposite verbs, while C2C [16] frequently confuses them, e.g. misclassifying ‘unfolding’ as ‘folding’. These results collectively demonstrate that C2C fails to overcome the critical challenge of object-driven shortcuts in verb learning, whereas RCORE directly targets this failure mode and learns verb representations that generalize robustly to unseen compositions.
The learned verb representations are indeed effective.
We further assess the quality of the learned verb representations by evaluating temporal modeling ability on the Temporal subset of Sth-com [16], as shown in Figure 5 (c). To measure a model’s dependence on static cues, we compare its performance when using temporally shuffled verb features $\mathbf{f}^{V}_{\text{shuffled}}$ versus the original features $\mathbf{f}^{V}$. Across both seen and unseen compositions, RCORE exhibits a substantially larger performance gap between original and shuffled features. This demonstrates that its verb representations are effectively grounded in the temporal dynamics of the input.
Improved unseen generalization.
In Figure 5 (a), we present the learning curve of our RCORE and the baseline [16] on Sth-com. Addressing the two key challenges in ZS-CAR, RCORE significantly improves unseen composition accuracy upon the baseline. While the baseline significantly increases FCP ratio from 10.8% at 5 epochs to 16.8% at 25 epochs, we observe our RCORE reports FCP ratio values of 10.8% at 5 epochs and 8.7% at 25 epochs, demonstrating stable mitigation of the co-occurrence bias. Notably, we also see a reduction in the seen-unseen accuracy gap.
Table 4: Ablation study on the Sth-com [16]. In every experiment, we use C2C [16] as our baseline. We report performance on both seen and unseen compositions, along with their harmonic mean (H.M.). We report the performances using the unbiased open-world setting. The best H.M. numbers are highlighted.
(a) Effects of RCORE components. VOCAMix TORC Margin Verb Object Composition Seen Unseen Seen Unseen Seen Unseen H.M. 63.60 54.36 67.72 56.10 46.31 30.08 36.47 64.26 54.65 67.32 56.61 46.86 30.35 36.84 65.65 56.80 68.59 55.19 48.65 31.35 38.13 64.60 54.42 66.99 55.68 46.71 29.80 36.38 65.72 56.60 67.91 55.81 47.68 31.96 38.27 65.93 57.29 67.82 55.56 48.56 32.23 38.75
(b) Effects of TORC components. $L_{\cos}$ $L_{\text{ent}}$ Verb Object Composition (OW) Seen Unseen Seen Unseen Seen Unseen H.M. 63.60 54.36 67.72 56.10 46.31 30.08 36.47 64.54 55.01 67.17 55.68 47.06 29.98 36.62 64.84 54.82 67.57 56.23 47.46 30.84 37.39 65.65 56.80 68.59 55.19 48.65 31.35 38.13
(c) Effects of VOCAMix. Method Verb Object Composition Seen Unseen Seen Unseen Seen Unseen H.M. Baseline 63.60 54.36 67.72 56.10 46.31 30.08 36.47 CutMix [38] 63.87 52.40 68.08 57.01 46.63 29.00 35.76 Mixup [41] 62.79 52.42 67.84 56.34 45.82 29.37 35.79 VOCAMix w/o F.E. 64.26 54.35 67.61 56.19 46.82 30.07 36.62 VOCAMix 64.26 54.65 67.32 56.61 46.86 30.35 36.84
(d) Effects of Margin loss. Method Verb Object Composition FCP Seen Unseen Seen Unseen Seen Unseen H.M. Baseline (CE) 63.60 54.36 67.72 56.10 46.31 30.08 36.47 15.30 L-softmax [19] 63.39 52.87 67.56 55.96 46.59 29.11 35.84 10.93 Margin (Ours) 64.60 54.42 66.99 55.68 46.71 29.80 36.38 8.81
5.3 Comparisons with State-of-the-Art
Sth-com results.
As shown in Table 2, RCORE delivers consistent improvements across all metrics under both unbiased and biased evaluation on the Sth-com [16] dataset. Importantly, RCORE achieves a positive compositional gap $(\Delta_{\mathrm{CG}})$ on unseen compositions under the open-world unbiased setting, whereas all baselines exhibit negative gaps. This highlights the strong compositional generalization ability of RCORE. The superior verb representations learned by RCORE—validated in Section 5.2—translate directly into substantial gains on the ZS-CAR dataset.
EK100-com results.
In Table 3, RCORE consistently outperforms C2C [16] on both seen and unseen compositions under both unbiased and biased evaluation settings, despite the egocentric challenges. Notably, RCORE achieves a substantial improvement in unseen verb accuracy compared to C2C (52.07% vs. 49.71%), underscoring its stronger temporal modeling ability and reduced object-driven shortcuts.
5.4 Ablation Studies
Effects of RCORE components.
In Table 4 (a), we investigate the effect of each component in our RCORE framework. When used individually, TORC loss is the most effective by providing explicit temporal modeling objectives. Notably, it significantly improves verb recognition performances across both seen and unseen compositions (63.6% vs. 65.7% on seen compositions and 54.4% vs. 56.8% on unseen compositions). Importantly, we can achieve the highest ZS-CAR performance by employing all three components together during training. This synergistic improvement is achieved by effectively tackling dataset sparsity with VOCAMix, reinforcing temporal order modeling with TORC, and alleviating co-occurrence bias with Margin loss.
Effects of TORC components.
Table 4 (b) evaluates the roles of $L_{\cos}$, which enforces discrimination of opposite temporal semantics, and $L_{\text{ent}}$, which prevents reliance on static cues. We show the each component of TORC can encourage performance improvement. Each term individually improves performance, while using both together yields the largest gains in unseen verb accuracy.
Effects of VOCAMix.
In Table 4 (c), we validate the effectiveness of our proposed composition-aware VOCAMix by comparing it with widely adopted mixing augmentations. We argue that conventional image-based augmentations are insufficient for composition verb-object learning. Specifically, CutMix [38] allows for mixing backgrounds with corresponding labels and Mixup [41] blends verbs and objects concurrently, failing to preserve the temporal difference information that is essential for verb recognition. As a result, these methods yield a lower H.M. score than the baseline. In contrast, our VOCAMix, aiming to generate realistic unseen composition samples, shows modest performance gains over the baseline. Furthermore, we observe that mixing a static frame into the estimated foreground region shows additional performance boost (+0.2 points in the H.M.) over applying it to the entire background (denoted as ‘VOCAMix w/o F.E.’).
Effects of Margin loss.
In Table 4 (d), we validate the effectiveness of our proposed Margin loss. To mitigate the model’s reliance on dataset co-occurrence priors while maintaining baseline performance, we employ our method and validate its effectiveness by comparing it with the widely-used large-margin softmax loss. The large-margin softmax (L-softmax) [19] loss is implemented by multiplying the angle $\theta$ between the input feature vector and the target classifier by a margin $m$ (e.g., $\cos(m\theta)$), thereby forcing the model to learn stricter decision boundaries with narrower angular margins. We apply this L-softmax loss into composition classification loss, consistent with the implementation schedule of our Margin loss. Compared to the L-softmax [19] approach, our proposed Margin loss successfully reduces FCP ratio while preserving performance. This demonstrates the effectiveness of our strategy, which selectively targets frequent-but-incorrect logits not the entire logits.
6 Conclusions
We revisit zero-shot compositional action recognition and show that existing methods fail mainly due to object-driven shortcuts, stemming from compositional sparsity and the asymmetric difficulty of learning verbs versus objects. We also find that the widely used closed-world, test-tuned evaluation protocol conceals this failure, overstating model generalization. We introduce RCORE, a simple framework that targets these issues directly: VOCAMix enlarges the effective compositional support, and TORC loss enforces temporal order sensitivity to reduce reliance on static object cues. Across two benchmarks and under a realistic open-world, unbiased evaluation setting, RCORE consistently improves unseen composition recognition performances, indicating genuine compositional reasoning. Our results highlight that both proper evaluation and shortcut-aware training are essential for progress in compositional video understanding.
Acknowledgments
This work was supported by the NAVER Cloud Corporation and partly supported by the Institute of Information & Communications Technology Planning & Evaluation(IITP) grant funded by the Korea Government (MSIT) under grant RS-2024-00353131 and ITRC(Information Technology Research Center) grant funded by the Korea government(MSIT)(IITP-2025-RS-2023-00259004). Additionally, it was supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT)(RS-2025-02216217) and (RS-2025-22362968).
Appendix
In this appendix, we provide comprehensive implementation/dataset/method details and quantitative/qualitative results to complement the main paper. We organize the appendix as follows:
  1. 1.
    Details on EK100-com dataset (Section A).
  2. 2.
    Complete implementation details (Section B).
  3. 3.
    Additional evidence of object-driven shortcuts (Section C).
  4. 4.
    Additional results (Section D).
Appendix A Details on EK100-com dataset
In this section, we provide details about our curated ZS-CAR benchmark, EPIC-KITCHENS-100-composition (EK100-com). We construct EK100-com by repurposing EPIC-KITCHENS-100 (EK100) [7] following the same protocol of constructing Sth-com [16]. In particular, we use the original training (67217 samples) and validation (9668 samples) split of EK100 [7] and split them as follows: (1) we filter the data to include only compositions with more than five samples each. (2) we ensure that all compositions present in the validation set also exist in the training set. After initial filtering, the training set comprises 62,790 samples with 1,331 unique compositions and the validation set contains 8,657 samples, which consist of 833 seen compositions and 3 unseen compositions. (3) we randomly select half of the compositions from both the training and validation sets and swap these subsets. (4) Finally, we split the validation and test set with a ratio of $3:4$. We provide the statistics of EK100-com in Table 5 and Figure 7, and some seen/unseen composition examples of EK100-com in Figure 8.
Beyond expanding the pool of publicly available ZS-CAR benchmarks, our repurposed EK100-com exhibits two key properties that make it a compelling ZS-CAR benchmark. First, unlike Something-Something V2 [12], where videos may involve multiple objects associated with a verb, videos in EK100 [7] depict actions consisting of a single verb and a single object. Therefore, the training data of EK100-com provides explicit object supervision for the models. Second, since EK100 [7] has long-tailed distribution, the label coverage ratio of EK100-com is 7.5%, which is much lower than that of Sth-com [16] (12.8%). This makes EK100-com a challenging benchmark, suitable for evaluating whether a model can overcome the risk of shortcut learning induced by data sparsity.
(a) Top-30 frequent compositions
(b) Bottom-30 frequent compositions
Figure 7: Top/Bottom-30 frequent compositions in the EK100-com training set.
Table 5: Details of the our introduced EK100-com benchmark. The label coverage ratio of EK100-com is $7.5\%=1320/(81\times 216)$.
EK100-com # Verb # Object # Compositions # Samples
Train 81 216 1133 54691
Val 57 175 462 Seen & 187 Unseen 3085 Seen & 4007 Unseen
Test 61 177 522 Seen & 187 Unseen 4197 Seen & 5258 Unseen
Total 81 216 1320 71238
(a) Seen composition examples
(b) Unseen composition examples
Figure 8: Example of seen/unseen composition samples in the EK100-com dataset.
Appendix B Complete Implementation Details
B.1 Implementation details on our diagnosis
In this section, we provide comprehensive details of experimental setup and implementation details of our diagnosis for Zero-Shot Compositional Action Recognition (ZS-CAR) in Section 3 in the main paper. All notations used in this supplementary document are consistent with those defined in the main paper. As mentioned in the main paper, we adopt the state-of-the-arts model, C2C [16] which uses a conditional learning approach, as the existing ZS-CAR baseline for diagnosis.
Details about our diagnosis in Figure 2 (a).
In Figure 2 (a) in the main paper, we empirically demonstrate the asymmetric learning difficulty between verbs and objects. We train a randomly initialized ViT [10] with joint spatio-temporal attention on our selected subset of the Sth-com [16] dataset. We aim to select a subset that maximizes the number of samples while maintaining a sufficient number of distinct labels. As a result, the selected subset is composed of 10 verb classes: ‘Moving sth and sth away from each other’, ‘Moving sth and sth closer to each other’, ‘Moving sth away from sth’, ‘Moving sth closer to sth’, ‘Moving sth down’, ‘Moving sth up’, ‘Pushing sth from left to right’, ‘Pushing sth from right to left’, ‘Pushing sth so that it falls off the table’, and ‘Pushing sth so that it slightly moves’; and 10 object classes: ‘Ball’, ‘Battery’, ‘Book’, ‘Bottle’, ‘Coin’, ‘Glass’, ‘Knife’, ‘Paper’, ‘Remote’, and ‘Spoon’. We construct the training set consisting of 2,555 samples from the original Sth-Com [16] training and validation samples, while the validation set comprises 260 samples from the original Sth-Com [16] test set. For training, we employ a multi-task learning approach using separate linear classifiers for verbs and objects, taking the ViT’s CLS token as input. In the main paper, we show that objects are easier to learn than verbs by monitoring the learning curves of a randomly initialized ViT. Furthermore, we substantiate in Section C demonstrating that this asymmetric learning difficulty leads the model to exploit object-driven shortcuts in compositional video understanding.
Details about our diagnosis in Figure 2 (b).
To precisely control the sample counts for each verb-object pair, we select a $4\times 4$ toy set from the Sth-Com [16] dataset. This set contains four verb labels (‘Moving sth down’, ‘Moving sth up’, ‘Pushing sth from left to right’, and ‘Pushing sth from right to left’) and four object labels (Book, Remote, Glass, Bottle). The training set consists of the four diagonal pairs (refer to Figure 2 (b)) with 40 samples per pair. The test set comprises two distinct splits: seen compositions (25 samples per pair) and unseen compositions (30 samples per pair). We use 8 frames for input videos.
Details about our diagnosis in Figure 3.
In Figure 3 in the main paper, we present the learning curves of the baseline model [16] trained on the Sth-com [16]. We further present diagnostic results for both the baseline and RCORE on EK100-com in Figure 9. We define two diagnostic metrics, False Seen Prediction (FSP) ratio and False Co-occurrence Prediction (FCP) ratio. FSP measures the proportion of unseen validation samples incorrectly classified as seen compositions:
$\textit{False Seen Prediction}=\frac{1}{|\mathbb{D}_{\text{val}}^{\text{unseen}}|}{\sum_{i=1}^{|\mathbb{D}_{\text{val}}^{\text{unseen}}|}\mathbb{I}(\hat{\mathbf{y}}_{i}\in\mathbb{Y}_{\text{seen}})},$ (7)
where $\mathbb{I}(\cdot)$ is the indicator function. FCP further quantifies the fraction of these false-seen predictions that correspond to frequently co-occurring verb–object pairs in the training data:
$\textit{False Co-oc. Prediction}=\frac{\sum_{i=1}^{|\mathbb{D}_{\text{val}}^{\text{unseen}}|}\mathbb{I}(\hat{\mathbf{y}}_{i}\in\mathbb{Y}_{\text{freq}})}{\sum_{i=1}^{|\mathbb{D}_{\text{val}}^{\text{unseen}}|}\mathbb{I}(\hat{\mathbf{y}}_{i}\in\mathbb{Y}_{\text{seen}})},$ (8)
where $\mathbb{Y}_{\text{freq}}$ denotes the set of frequently observed verb–object pairs in the training set. $\mathbb{Y}_{\text{freq}}$ is defined with the frequency distribution of the training set $P_{\mathbb{D}_{train}}$ as follows:
$\mathbb{Y}_{\text{freq}}$ $=\mathbb{Y}_{\text{freq}}^{O\mid V}\cap\mathbb{Y}_{\text{freq}}^{V\mid O},$ (9)
where $\mathbb{Y}_{\text{freq}}^{O\mid V}$ is defined as the set of pairs where the conditional probability of an object given a verb in the training data is greater than $\mu+\sigma$. Here, $\mu$ and $\sigma$ represent the mean and standard deviation of the conditional probabilities calculated on the training set. The $\mathbb{Y}_{\text{freq}}^{V\mid O}$ is defined similarly for the conditional probability of a verb given an object. $(\mu_{O|V}+\sigma_{O|V},\mu_{V|O}+\sigma_{V|O})$ are calculated as $(0.13,0.19)$ for the Sth-com [16] dataset and $(0.23,0.40)$ for EK100-com.
Figure 9: Learning curve of the baseline and RCORE on the EK100-com dataset. RCORE suppresses the increase of the FCP ratio during training, effectively narrowing the performance gap between seen and unseen composition validation accuracies.
B.2 Implementation details on RCORE
Experimental Setup.
We implement our method using PyTorch and conduct all experiments on 8 NVIDIA Tesla V100 GPUs. For video preprocessing, we uniformly sample frames from each video to generate clips of size $3\times 16\times 224\times 224$ (channels $\times$ frames $\times$ height $\times$ width).
Model Architecture.
We utilize CLIP-B/16 [28] as our visual/text backbone. Following prior work [16], we adopt AIM [37], a parameter-efficient spatio-temporal adapter for CLIP [28], as the video encoder. We also construct the verb encoder with two temporal convolution layers with ReLU activations, while the object encoder employs temporal average pooling followed by a two-layer MLP with ReLU activations. We use CoOp-style [43] learnable text prompts for all models.
Training Strategy.
We train all models, including baselines and RCORE, for 30 epochs with a total batch size of 128. The base learning rate is set to $5\times 10^{-4}$ for Sth-com [16] and $1\times 10^{-4}$ for EK100-com, initialized with a 3-epoch warmup. Following prior work [16], we adopt closed-world training setting. We empirically find that setting the softmax temperature to 0.07, same as the implementation of CLIP [28], results in better ZS-CAR performance than the value of 0.01, used in prior work [16]. Therefore, we adopt this temperature value for both C2C [16] and RCORE, except for the case of AIM [37] with independent modeling. For VOCAMix, we employ the unsupervised foreground estimation algorithm FAME [9] to identify high-motion regions. FAME [9] isolates the moving foreground from background regions based on frame differences and color statistics. To generate the mixed samples, we first sample the mixing coefficient from a $\text{Beta}(2.0,2.0)$ distribution, following prior work [38, 41]. We then scale it by 0.2 for Sth-com [16] and 0.5 for EK100-com to obtain the final blending coefficient $\lambda$. We summarize the hyperparameters used in this work in Table 6.
Loss configuration.
For the total loss, we set the weights for the verb and object components to $\alpha=0.2$ and for the composition loss to $\beta=1.0$, following prior work [16]. As demonstrated in Table 9 (g) and (h), we empirically find it beneficial to linearly warm up the weights for the auxiliary losses: $\gamma$ is increased from 0.0 to 1.0 over 5 epoch to 10 epoch, and $\delta$ from 0.0 to 0.8 over epochs 15 epoch to 20 epoch.
Inference.
During inference, we adopt a single-view, single-crop protocol. For bias calibration, we select the bias term that maximizes the harmonic mean of the accuracies for seen and unseen compositions.
Table 6: Hyperparameters used for training RCORE on each dataset.
Config Sth-com EK100-com
Optimizer AdamW [20]
Base learning rate (visual) 5e-4 1e-4
Base learning rate (text) 1e-4 1e-5
Weight decay (visual) 1e-4
Weight decay (text) 1e-5
Optimizer momentum $\beta_{1},\beta_{2}=0.9,0.999$
Per GPU batch size 16
Warmup epochs 3
$\alpha,\beta,\gamma,\delta$ 0.2, 1.0, 1.0, 0.8
$p_{aug}$ 0.5 0.3
(start, end) epoch for scaling $\gamma$ (5, 10)
(start, end) epoch for scaling $\delta$ (15, 20)
Scale factor for $\lambda$ in VOCAMix 0.2 0.5
K for top-k selection in Margin loss 10
Softmax temperature 0.07
Training epochs 30
Table 7: Top-10 failure cases of the existing ZS-CAR model on unseen compositions of the Sth-com test set. We employ C2C [16] trained on Sth-com [16] as baseline for analysis. We use ‘sth’ instead of ‘something’ to shorten. We report the results with unbiased open-world inference scores. ‘Prediction Rate’ denotes the proportion of mispredictions classified as the specific composition. ‘Verb/Object-side’ indicates the most frequently co-occurring counterpart for each mispredicted component in the training set, along with its co-occurrence ratio. The top-10 cases listed below account for 9.5% of the total mispredictions on unseen compositions.
Rank Ground Truth (Verb, Object) Error Top-2 Misprediction (seen/unseen) Prediction Most Frequent Co-occurring Component During Training
Count / Rate (%) Count / Rate (%) Verb-side Object-side
1 (‘Pretending to be tearing [sth that is not tearable]’, ‘Paper’) 192 / 99.0% (‘Tearing [sth] just a little bit’, ‘Paper’) (seen) 116 / 60.4% w/ ‘Paper’ (71.6%) w/ ‘Tearing [sth] just a little bit’ (26.7%)
(‘Folding [sth]’, ‘Paper’) (seen) 47 / 24.5% w/ ‘Paper’ (43.9%) w/ ‘Tearing [sth] just a little bit’ (26.7%)
2 (‘Unfolding [sth]’, ‘Paper’) 140 / 44.7% (‘Folding [sth]’, ‘Paper’) (seen) 80 / 57.1% w/ ‘Paper’ (43.9%) w/ ‘Tearing [sth] just a little bit’ (26.7%)
(‘Tearing [sth] into two pieces’, ‘Paper’) (unseen) 12 / 8.6% w/ ‘Envelope’ (17.4%) w/ ‘Tearing [sth] just a little bit’ (26.7%)
3 (‘Opening [sth]’, ‘Box’) 106 / 82.8% (‘Showing that [sth] is inside [sth]’, ‘Box’) (seen) 27 / 25.5% w/ ‘Box’ (22.7%) w/ ‘Showing that [sth] is empty’ (10.1%)
(‘Showing that [sth] is empty’, ‘Box’) (seen) 23 / 21.7% w/ ‘Cup’ (30.8%) w/ ‘Showing that [sth] is empty’ (10.1%)
4 (‘Squeezing [sth]’, ‘Plastic bag’) 94 / 40.9% (‘Squeezing [sth]’, ‘Bag’) (seen) 77 / 81.9% w/ ‘Sponge’ (11.0%) w/ ‘Stuffing [sth] into [sth]’ (15.4%)
(‘Folding [sth]’, ‘Plastic bag’) (unseen) 13 / 13.8% w/ ‘Paper’ (43.9%) w/ ‘[sth] falling like a feather or paper’ (57.4%)
5 (‘Putting [sth] into [sth]’, ‘Box’) 88 / 60.3% (‘Stuffing [sth] into [sth]’, ‘Box’) (seen) 10 / 11.4% w/ ‘Box’ (19.2%) w/ ‘Showing that [sth] is empty’ (10.1%)
(‘Putting [sth] into [sth]’, ‘Plastic box’) (seen) 8 / 9.1% w/ ‘Bowl’ (12.4%) w/ ‘Closing [sth]’ (25.9%)
6 (‘Lifting up one end of [sth], then letting it drop down’, ‘Card’) 79 / 54.1% (‘Lifting [sth] up completely without letting it drop down’, ‘Card’) (unseen) 20 / 25.3% w/ ‘Crayon’ (26.8%) w/ ‘Uncovering [sth]’ (19.8%)
(‘Lifting up one end of [sth], then letting it drop down’, ‘Pencil’) (unseen) 19 / 24.1% w/ ‘Crayon’ (26.8%) w/ ‘Uncovering [sth]’ (13.7%)
7 (‘Putting [sth] on a surface’, ‘Bottle’) 77 / 68.1% (‘Putting [sth] upright on the table’, ‘Bottle’) (seen) 33 / 42.9% w/ ‘Bottle’ (42.1%) w/ ‘Turning [sth] upside down’ (5.3%)
(‘Hitting [sth] with [sth]’, ‘Bottle’) (seen) 15 / 19.5% w/ ‘Bottle’ (14.8%) w/ ‘Turning [sth] upside down’ (5.3%)
8 (‘Showing that [sth] is empty’, ‘Bottle’) 70 / 85.4% (‘Turning [sth] upside down’, ‘Bottle’) (seen) 13 / 18.6% w/ ‘Bottle’ (21.7%) w/ ‘Turning [sth] upside down’ (5.3%)
(‘Pretending to turn [sth] upside down’, ‘Bottle’) (seen) 8 / 11.4% w/ ‘Bottle’ (30.6%) w/ ‘Turning [sth] upside down’ (5.3%)
9 (‘Throwing [sth]’, ‘Book’) 69 / 84.2% (‘[sth] falling like a rock’, ‘Book’) (seen) 16 / 23.2% w/ ‘Bottle’ (13.0%) w/ ‘Pushing [sth] from left to right’ (4.9%)
(‘Throwing [sth]’, ‘Hat’) (seen) 9 / 13.0% w/ ‘Tissue’ (7.7%) w/ ‘Throwing [sth]’ (35.0%)
10 (‘Closing [sth]’, ‘Drawer’) 60 / 69.0% (‘Opening [sth]’, ‘Drawer’) (seen) 31 / 51.7% w/ ‘Door’ (18.9%) w/ ‘Opening [sth]’ (36.8%)
(‘Stuffing [sth] into [sth]’, ‘Drawer’) (seen) 6 / 10.0% w/ ‘Box’ (19.2%) w/ ‘Opening [sth]’ (36.8%)
Table 8: Top-10 failure cases of the existing ZS-CAR model on unseen compositions of the EK100-com test set. We employ C2C [16] trained on EK100-com as baseline for analysis. We report the results with unbiased open-world inference scores. ‘Prediction Rate’ denotes the proportion of mispredictions classified as the specific composition. ‘Verb/Object-side’ indicates the most frequently co-occurring counterpart for each mispredicted component in the training set, along with its co-occurrence ratio. The top-10 cases listed below account for $\sim$ 39% of the total mispredictions on unseen compositions.
Rank Ground Truth (Verb, Object) Error Top-2 Misprediction (seen/unseen) Prediction Most Frequent Co-occurring Component During Training
Count / Rate (%) Count / Rate (%) Verb-side Object-side
1 (‘put’, ‘knife’) 340 / 80.8% (‘take’, ‘knife’) (seen) 81 / 23.8% w/ ‘spoon’ (6.8%) w/ ‘take’ (56.2%)
(‘put’, ‘fork’) (seen) 41 / 12.1% w/ ‘spoon’ (8.6%) w/ ‘take’ (40.3%)
2 (‘put’, ‘lid’) 303 / 76.7% (‘take’, ‘lid’) (seen) 44 / 14.5% w/ ‘spoon’ (6.8%) w/ ‘take’ (45.4%)
(‘remove’, ‘lid’) (seen) 22 / 7.3% w/ ‘lid’ (16.1%) w/ ‘take’ (45.4%)
3 (‘take’, ‘plate’) 259 / 46.7% (‘put’, ‘plate’) (seen) 70 / 27.0% w/ ‘spoon’ (8.6%) w/ ‘put’ (52.1%)
(‘take’, ‘bowl’) (seen) 32 / 12.4% w/ ‘spoon’ (6.8%) w/ ‘take’ (32.6%)
4 (‘wash’, ‘sink’) 143 / 100% (‘wash’, ‘top’) (seen) 68 / 47.6% w/ ‘hand’ (12.7%) w/ ‘wash’ (89.2%)
(‘wash’, ‘cloth’) (seen) 14 / 9.8% w/ ‘hand’ (12.7%) w/ ‘take’ (35.2%)
5 (‘put’, ‘spatula’) 131 / 69.0% (‘take’, ‘spatula’) (seen) 42 / 32.1% w/ ‘spoon’ (6.8%) w/ ‘take’ (58.2%)
(‘put’, ‘spoon’) (seen) 18 / 13.7% w/ ‘spoon’ (6.8%) w/ ‘take’ (36.9%)
6 (‘put’, ‘cup’) 115 / 56.1% (‘take’, ‘cup’) (seen) 25 / 21.7% w/ ‘spoon’ (6.8%) w/ ‘take’ (48.8%)
(‘put’, ‘glass’) (seen) 11 / 9.6% w/ ‘spoon’ (8.6%) w/ ‘take’ (36.4%)
7 (‘shake’, ‘hand’) 94 / 97.9% (‘turn-off’, ‘tap’) (seen) 38 / 40.4% w/ ‘tap’ (93.6%) w/ ‘turn-on’ (49.5%)
(‘wash’, ‘hand’) (seen) 18 / 19.1% w/ ‘hand’ (12.7%) w/ ‘wash’ (59.8%)
8 (‘put’, ‘board or chopping’) 89 / 56.0% (‘take’, ‘board or chopping’) (seen) 11 / 12.4% w/ ‘spoon’ (6.8%) w/ ‘take’ (45.1%)
(‘put’, ‘knife’) (unseen) 9 / 10.1% w/ ‘spoon’ (8.6%) w/ ‘take’ (56.2%)
9 (‘wash’, ‘hob’) 83 / 70.9% (‘wash’, ‘top’) (seen) 38 / 45.8% w/ ‘hand’ (12.7%) w/ ‘wash’ (89.2%)
(‘wash’, ‘oven’) (seen) 7 / 10.3% w/ ‘hand’ (12.7%) w/ ‘open’ (35.9%)
10 (‘take’, ‘salt’) 74 / 70.5% (‘put’, ‘salt’) (seen) 9 / 12.2% w/ ‘spoon’ (8.6%) w/ ‘put’ (38.0%)
(‘pour’, ‘salt’) (seen) 7 / 9.5% w/ ‘water’ (25.2%) w/ ‘put’ (37.9%)
Appendix C Additional Evidence of Object-Driven Shortcuts
In this section, we provide further evidence demonstrating that existing ZS-CAR learning approaches are susceptible to object-driven shortcut learning.
Failure case analysis of the existing ZS-CAR model.
In Table 7 and Table 8, we investigate top-10 failure cases of the model trained with the existing ZS-CAR learning approach [16]. The top-10 failure cases shown in the Table 7 constitute 9.5% of the misclassifications on unseen compositions for Sth-com [16], compared to 49% for EK100-com as shown in Table 8. For the analysis of Sth-com [16], we exclude ambiguous cases arising from annotation issues to obtain meaningful results. A prime example is the top-1 failure case, (‘Covering [something] with [something]’, ‘Paper’), where 66% of the errors are misclassified as ‘Pencil’. Upon visual inspection, we confirm that nearly all these videos actually depict the action of covering a pencil with paper. In Table 7 and Table 8, we observe a common tendency across both datasets: the model frequently misclassifies verbs as their temporally reversed or opposing counterparts. Specifically, we demonstrate that these mispredictions stem from the model’s reliance on co-occurrence statistics. For instance, the model misclassifies unseen ‘(Closing, Drawer)’ samples as ‘(Opening, Drawer)’, which accounts for 36% of the errors. Notably, ‘Opening’ is the verb most frequently paired with ‘Drawer’ in the training set. These findings suggest that the model has failed to learn robust verb representations that generalize to unseen compositions, largely due to object-driven shortcut learning.
Table 9: Additional ablation study of RCORE on the Sth-com. We provide the additional ablation study results on Sth-com [16]. In every experiment, we use C2C [16] with CLIP-B/16 [28] as our baseline. We report performance on both seen and unseen compositions, along with their harmonic mean (H.M.).
(a) Effects of TORC with different temporal aggregation.
(b) Effects top-$K$ selection for Margin loss.
Method TORC Verb Object Composition
Seen Unseen Seen Unseen Seen ($\Delta_{\mathrm{CG}}$) Unseen ($\Delta_{\mathrm{CG}}$) H.M.
Average $\times$ 63.60 54.36 67.72 56.10 46.31 (+3.24) 30.08 (-0.42) 36.47
Attn. pool $\times$ 64.28 54.76 67.50 55.73 47.11 (+3.72) 30.09 (-0.43) 36.72
Average 65.65 56.80 68.59 55.19 48.65 (+3.62) 31.35 (-0.00) 38.13
Attn. pool 65.51 56.50 67.80 55.00 48.07 (+3.65) 30.78 (-0.30) 37.53
$K$ Verb Object Composition FCP
Seen Unseen Seen Unseen Seen Unseen H.M.
10 64.60 54.42 66.99 55.68 46.71 29.80 36.38 8.81
100 63.63 53.93 67.58 55.66 46.43 29.18 35.84 9.89
3451 62.75 53.63 66.89 55.20 45.32 29.16 35.49 8.02
(c) Effects of hyperparameters of VOCAMix.
(d) Effects of hyperparameters of TORC.
$p_{aug}$ Verb Object Composition
Seen Unseen Seen Unseen Seen Unseen H.M.
0.0 63.60 54.36 67.72 56.10 46.31 30.08 36.47
0.3 64.67 54.83 68.18 55.37 47.77 29.93 36.80
0.5 64.26 54.65 67.32 56.61 46.86 30.35 36.84
1.0 64.02 53.80 67.04 55.26 46.30 29.07 35.72
$\gamma$ scaling scaling Verb Object Composition
start epoch end epoch Seen Unseen Seen Unseen Seen Unseen H.M.
0.0 0 0 63.60 54.36 67.72 56.10 46.31 30.08 36.47
0.5 1 1 65.59 55.63 68.49 54.80 48.33 30.19 37.17
1.0 1 1 64.88 55.81 67.21 55.12 47.06 30.69 37.15
1.0 5 5 64.93 56.03 67.67 55.81 47.20 31.38 37.70
1.0 5 10 65.65 56.80 68.59 55.19 48.65 31.35 38.13
(e) Effects of hyperparameters of Margin loss.
$\delta$ scaling scaling Verb Object Composition
start epoch end epoch Seen Unseen Seen Unseen Seen Unseen H.M.
0.3 10 10 63.09 52.94 67.90 54.86 46.31 28.48 35.27
0.3 15 15 63.96 53.70 67.70 55.68 46.57 29.28 35.96
0.3 15 20 64.01 53.76 67.53 55.61 46.87 29.42 36.15
0.5 15 20 63.85 53.61 67.90 56.32 46.79 29.49 36.18
0.8 15 20 64.60 54.42 66.99 55.68 46.71 29.80 36.38
Appendix D Additional Results
D.1 Comprehensive ablation results
In Table 9, we provide the comprehensive ablation study results to examine each design choice of RCORE, using the Sth-com [16] dataset.
Effects of TORC loss with different temporal aggregation methods.
In Table 9 (d), we demonstrate that the weak temporal order modeling caused by object-driven shortcut learning cannot be resolved merely by employing more sophisticated temporal aggregation methods. We compare the baseline’s temporal average pooling with attention pooling (Attn. pool). In attention pooling, a linear layer computes frame-wise attention weights from the input verb feature sequences, aggregating them via a weighted sum to obtain the final verb feature for classification. While attention pooling yields a slight improvement in verb accuracy, it fails to reduce the compositional gap ($\Delta_{CG}$) on unseen compositions. In contrast, our TORC loss achieves significant improvements in H.M. with both aggregation methods. Notably, the combination of temporal average pooling and TORC loss enables the model to overcome the negative compositional gap.
Effects of top-$K$ selection for Margin loss.
In Table 9 (e), we investigate the impact of the hyperparameter $K$, which determines the number of top-ranked logits to which the Margin loss is applied. We find that applying the margin selectively to the frequent-but-incorrect logits within the top-10 ($K=10$) reports the highest H.M., outperforming the approach of applying it to all seen composition logits ($K=3451$) (36.4% vs. 35.5%). This suggests that penalizing only the most confusing candidates—specifically those that frequently co-occur with the input components—effectively mitigates co-occurrence bias without compromising the model’s overall performance. For instance, given an input like ‘(Pretending to tear, Paper)’, our intention is to penalize plausible but incorrect distractors such as ‘(Tearing, Paper)’ or ‘(Folding, Paper)’, while avoiding penalties on irrelevant classes like ‘(Throwing, Paper)’ that are already ranked low.
(a) Our reconstructed Temporal/Static splits
(b) Temporal/Static splits from Sevilla et al. [31]
Figure 10: Performances on Temporal/Static split of Sth-com. We evaluate the models on Sth-com [16] using both (a) our reconstructed splits and (b) the splits from Sevilla et al [31]. We utilize both original and temporally shuffled inputs to assess the model’s temporal modeling capability and its reliance on static cues. A larger performance gap between original and shuffled inputs indicates that the model predicts verbs more based on temporal dynamics rather than on static cues.
Effects of hyperparameters of RCORE.
In Table 9 (f), we demonstrate that applying augmentation with a probability of 0.5 within a batch is the most effective, achieving a 0.4 point improvement in H.M. over the baseline. In Table 9 (g), we observe that introducing the explicit temporal order modeling task via TORC starting from 5 epoch, after initial training, is effective. In particular, the strategy of linearly increasing the loss weight from epoch 5 to 10 is the most effective, resulting in 1.7 point boost in H.M. compared to the baseline. Finally, in Table 9 (h), we analyze the impact of scaling the weight of our Margin loss, revealing a trade-off between improved unseen composition accuracy and decreased seen composition accuracy. We observe that applying Margin loss starting from epoch 15, the point at which the model begins to overfit the training data as shown in Figure 3 of the main paper, is more effective than starting from the earlier epoch. We also find that linearly scaling the loss weight over 5 epochs is the most effective strategy to maintain the model’s performance.
D.2 Analysis of confidence distribution in composition classification
(a) Sth-com results
(b) EK100-com results
Figure 11: Conditional modeling overfits to co-occurrence statistics. We track how the composition prediction confidence of C2C [16] evolves during training both on Sth-com [16] and EK100-com datasets. As training progresses, C2C [16] increasingly ignores input evidence and misclassifies unseen compositions as seen ones.
In Figure 11, we track how the composition prediction confidence of C2C [16] evolves during training both on Sth-com [16] and EK100-com datasets. Using the test set of both datasets, we compare two cases of the model’s confidence values: Seen $\to$ Seen and Unseen $\to$ Seen. Given the seen/unseen composition set $\mathbb{Y}_{\text{seen}}$$\mathbb{Y}_{\text{unseen}}$ and the ground-truth/model-predicted composition label $\mathbf{y}$$\hat{\mathbf{y}}$, Seen $\to$ Seen refers to the case where $\mathbf{y}_{i}\in\mathbb{Y}_{\text{seen}}\land\hat{\mathbf{y}}_{i}\in\mathbb{Y}_{\text{seen}}$ and Unseen $\to$ Seen refers to the case where $\mathbf{y}_{i}\in\mathbb{Y}_{\text{unseen}}\land\hat{\mathbf{y}}_{i}\in\mathbb{Y}_{\text{seen}}$. Across both datasets, we observe that the model increasingly classifies unseen compositions as seen compositions as training progresses. The observation suggests that, without careful design, models may be prone to ignore input cues and instead over-rely on verb–object co-occurrence statistics present in the training data.
D.3 Temporal/Static split of Sth-com
Table 10: Ablation study on the EK100-com. We provide the results of the ablation study on EK100-com presented in the main paper. In every experiment, we use C2C [16] with CLIP-B/16 [28] as our baseline. We report performance on both seen and unseen compositions, along with their harmonic mean (H.M.).
(a) Effects of VOCAMix and TORC. VOCAMix TORC Verb Object Composition (OW) Seen Unseen Seen Unseen Seen Unseen H.M. 66.19 49.71 56.61 47.48 42.72 22.38 29.38 65.55 52.15 57.09 46.87 43.01 23.81 30.65 65.90 52.31 57.28 46.74 43.48 23.67 30.65 66.21 53.51 56.97 47.52 43.53 25.02 31.78
(b) Effects of each TORC loss term. $L_{\cos}$ $L_{\text{ent}}$ Verb Object Composition (OW) Seen Unseen Seen Unseen Seen Unseen H.M. 66.19 49.71 56.61 47.48 42.72 22.38 29.38 65.43 52.11 56.54 47.44 42.67 24.24 30.92 66.00 51.06 56.64 47.37 42.82 22.53 29.53 65.90 52.31 57.28 46.74 43.48 23.67 30.65
We present extended results corresponding to the experiment shown in Figure 5 (c) of the main paper, including all versions of the Temporal/Static split. Inspired by prior work [31, 39], we rigorously evaluate the temporal modeling capabilities of both the baseline (C2C [16]) and our RCORE by partitioning the Sth-com [16] dataset into Temporal and Static splits. Sevilla et al. [31] originally defined these splits based on cognitive science experiments on Something-Something V2 [12] verb classes: a Temporal split (18 verbs) where temporal information matters, and a Static split (18 verbs) where such information is redundant. However, noting that this Temporal split includes verbs requiring complex temporal reasoning such as ‘Folding’, we reconfigure the splits based on verb semantics. Specifically, we categorized verbs based on this criterion: those requiring temporal information for discrimination are assigned to the Temporal split (62 verbs), while verbs distinguishable by momentary keyframes are assigned to the Static split (80 verbs). The indices of the verbs in the Temporal split are as follows: 0, 1, 5, 6, 14, 25, 26, 29, 30, 32, 33, 34, 35, 38, 39, 40, 41, 42, 43, 44, 45, 46, 58, 59, 60, 61, 65, 66, 67, 68, 72, 73, 74, 75, 76, 78, 79, 85, 86, 90, 91, 92, 98, 100, 103, 109, 122, 123, 131, 134, 136, 137, 139, 140, 145, 146, 147, 148, 149, 150, 153, 154. The indices of the verbs in the Static split are as follows: 2, 3, 8, 9, 10, 11, 12, 13, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 27, 28, 31, 36, 37, 47, 48, 49, 50, 51, 52, 53, 54, 55, 77, 80, 81, 87, 88, 93, 94, 95, 96, 97, 99, 101, 102, 104, 105, 106, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 126, 127, 128, 129, 130, 132, 133, 135, 138, 141, 142, 143, 151, 152, 155, 156, 157, 158, 159, 160. To balance the sample sizes between the Temporal and Static splits, we excluded a few verbs with the fewest samples from the Temporal split. We then utilize both original and temporally shuffled inputs to assess the model’s temporal modeling capability and its reliance on static cues. In Figure 10, we present the results for the Temporal/Static splits across seen/unseen compositions. We also provide results applying the original Temporal/Static verb split [31] to Sth-com [16]. In all cases, RCORE exhibits lower performance on shuffled inputs but outperforms the baseline on original inputs. This indicates that RCORE predicts verbs without relying on static cues, leading to robust improvements in verb performance regardless of the split.
D.4 Ablation studies on EK100-com
In Table 10, we provide the ablation study results conducted on the EK100-com dataset, complementing the results discussed with the Sth-com [16] dataset in the main paper. In Table 10 (a), we demonstrate that our framework, integrating VOCAMix for sparsity resolution and TORC for explicit temporal modeling, outperforms the baseline on EK100-com, achieving 3.8 points higher verb accuracy in unseen compositions than the baseline. In Table 10 (b), using only $L_{cos}$ in TORC yields the highest harmonic mean of composition accuracy on EK100-com.

### v2

@@
 Abstract
-We study Compositional Video Understanding (CVU), where models must recognize verbs and objects and compose them to generalize to unseen combinations.
-We find that existing Zero-Shot Compositional Action Recognition (ZS-CAR) models fail primarily due to an overlooked failure mode: object-driven verb shortcuts.
-Through systematic analysis, we show that this behavior arises from two intertwined factors—severe sparsity and skewness of compositional supervision, and the asymmetric learning difficulty between verbs and objects.
-As training progresses, the existing ZS-CAR model increasingly ignores visual evidence and overfits to co-occurrence statistics.
-Consequently, the existing model does not gain the benefit of compositional recognition in unseen verb-object compositions.
-To address this, we propose RCORE, a simple and effective framework that enforces temporally grounded verb learning.
-RCORE introduces (i) a composition-aware augmentation that diversifies verb–object combinations without corrupting motion cues, and (ii) a temporal order regularization loss that penalizes shortcut behaviors by explicitly modeling temporal structure.
-Across two benchmarks—Sth-com and our newly constructed EK100-com—RCORE significantly improves unseen composition accuracy, reduces reliance on co-occurrence bias, and achieves consistently positive compositional gaps.
-Our findings reveal object-driven shortcuts as a critical limiting factor in ZS-CAR and demonstrate that addressing them is essential for robust compositional video understanding.
-The code is available at https://github.com/KHU-VLL/RCORE.
-(a) Co-occurrence bias
-(b) Asymmetric learning difficulty
+Zero-Shot Compositional Action Recognition (ZS-CAR) requires recognizing novel verb–object combinations composed of previously observed primitives.
+In this work, we tackle a key failure mode: models predict verbs via object-driven shortcuts (i.e., relying on the labeled object class) rather than temporal evidence.
+We argue that sparse compositional supervision and verb–object learning asymmetry can promote object-driven shortcut learning.
+Our analysis with proposed diagnostic metrics shows that existing methods overfit to training co-occurrence patterns and underuse temporal verb cues, resulting in weak generalization to unseen compositions.
+To address object-driven shortcuts, we propose Robust COmpositional REpresentations (RCORE) with two components.
+Co-occurrence Prior Regularization (CPR) adds explicit supervision for unseen compositions and regularizes the model against frequent co-occurrence priors by treating them as hard negatives.
+Temporal Order Regularization for Composition (TORC) enforces temporal-order sensitivity to learn temporally grounded verb representations.
+Across Sth-com and EK100-com, RCORE reduces shortcut diagnostics and consequently improves compositional generalization.
+This work was done during an internship at NAVER Cloud.Corresponding author.
+1 Introduction
+Many human actions can be decomposed into two semantic primitives—verbs and objects—and robust video understanding requires recognizing each component and reasoning over their interaction as a composition [goyal2017something, mat2020sthelse, damen2022ek100, diba2020hvu, bae2024devias].
+Zero-Shot Compositional Action Recognition (ZS-CAR) formalizes this goal: a model must recognize unseen verb–object pairs while keeping the verb and object vocabularies shared across splits [li2024c2c, jiang2025dhd, jung2025crr, ye2025logic].
+In this setting, a key failure mode emerges: models often predict verbs via object-driven shortcuts rather than temporal evidence.
+Throughout, object refers to the labeled noun class in the verb–object vocabulary (not unlabeled background entities).
+We argue that object-driven shortcuts arise from two issues inherent to ZS-CAR.
 Figure 1: Why object-driven shortcuts emerge in compositional video understanding?
@@
 Datasets are intrinsically sparse and highly skewed in their verb–object combinations, creating strong co-occurrence priors.
-Models exploit these priors as a shortcut: once the object is recognized, the model often predicts the most frequent verb paired with it, ignoring temporal evidence.
 (b) Asymmetric learning difficulty.
@@
 This imbalance makes object features dominate training signals.
+(c) Object-driven shortcuts.
 Together, these two factors drive models to adopt object-driven shortcuts, hindering genuine compositional generalization.
-Best viewed with zoom and color.
-11footnotetext: This work was done during an internship at NAVER Cloud.22footnotetext: Corresponding author.
-1 Introduction
-Many human actions can be decomposed into two semantic primitives—verbs and objects—and robust video understanding requires recognizing each component and reasoning over their interaction as a composition [12, 22, 7, 8, 2].
-Zero-Shot Compositional Action Recognition (ZS-CAR) formalizes this goal: a model must correctly identify unseen verb–object combinations despite only being exposed to a sparse subset of compositions during training [16].
-However, existing ZS-CAR models exhibit a characteristic and persistent failure: they mispredict verbs in unseen compositions by relying on object-driven shortcuts.
-In other words, rather than learning verb semantics from temporal evidence, models exploit objects as the dominant cue whenever such shortcuts correlate with training labels.
-To the best of our knowledge, this failure mode has not been explicitly identified or systematically addressed in prior ZS-CAR approaches [16, 14].
-We show that this shortcut failure mode arises from two intertwined issues of the ZS-CAR setting: (i) Sparsity and skewness of compositional supervision.
-The combinatorial nature of verb–object pairs yields extreme under-coverage of the compositional space, e.g.
-only $\sim 14.0\%$ label coverage ratio in the Something-Else [22] and EPIC-KITCHENS-100 [7] datasets.
-This induces strong co-occurrence statistics to the deep learning model as depicted in Figure 1 (a).
-(ii) Asymmetric learning difficulty between verbs and objects.
-Objects are recognizable from single frames, while verbs require multi-frame temporal reasoning, as shown in Figure 1 (b).
-This imbalance naturally encourages models to treat objects as easy predictive shortcuts [25, 17, 32, 11, 30], especially under sparse data regimes.
-Our analysis reveals that the existing ZS-CAR model amplifies these issues during training.
-They increasingly bias predictions toward frequently co-occurring compositions, exhibit growing false-seen and false-co-occurrence prediction ratios, and ultimately yield negative compositional gaps—indicating that they fail to extract any benefit from compositional reasoning beyond independent verb and object recognition.
-(See Section 3 for details.) These observations call into question the robustness of current ZS-CAR approaches and motivate the need for techniques that explicitly target shortcut mitigation.
-We propose RCORE (Robust COmpositional REpresentations), a simple and effective framework designed to mitigate object-driven shortcut and promote temporally grounded verb representations.
-RCORE consists of two components: (i) a composition-aware augmentation that synthesizes plausible unseen verb–object combinations while preserving the temporal structure of the original video; and (ii) a temporal order regularization objective that penalizes reliance on static cues by contrasting original and perturbed temporal representations.
-Together, these components address the two root causes of failure in ZS-CAR: compositional sparsity and asymmetric learning difficulty.
-Across both datasets, RCORE significantly strengthens verb recognition, reduces co-occurrence bias, produces positive compositional gaps, and improves unseen composition accuracy over strong conditional baselines.
-Our findings indicate that the primary obstacle in compositional video understanding is not model capacity but model behavior: ZS-CAR models fail because they resort to object-driven shortcuts.
-By directly tackling this failure mode, RCORE provides a simple path toward more reliable and generalizable ZS-CAR.
-Figure 2: Controlled experiments demonstrate object-driven shortcut learning in ZS-CAR.
-We empirically identify a key failure mode in ZS-CAR—object-driven shortcuts.
-(a) Objects are easier to learn than verbs.
-We train a randomly initialized ViT [10] on a balanced 10 $\times$ 10 verb-object subset from Sth-com [16].
-The learning curves show that object accuracy increases much faster than verb accuracy, indicating that objects are easier to learn.
-(b) Under a strong co-occurrence bias, models exploit object cues as shortcuts, failing to learn robust verb representations.
-We prepare a perfectly biased training set and two kinds of test splits utilizing the Sth-com [16] dataset: seen compositions (bias-aligned) and unseen compositions (bias-conflict).
-Both the randomly initialized ViT [10] and pretrained CLIP [28] exhibit a consistent trend: while unseen object accuracy increases alongside seen performance, unseen verb accuracy concurrently drops below random chance.
-Best viewed with zoom and color.
+Once the object is recognized, the model often predicts the most frequent verb paired with it, ignoring temporal evidence.
+First, compositional supervision is sparse and skewed.
+As illustrated in Figure 1 (a), datasets cover only a small fraction of the combinatorial verb–object space, and training labels concentrate on a limited set of frequent pairs [mat2020sthelse, damen2022ek100].
+This skew induces a strong co-occurrence prior in the training data—the empirical dominance of frequent verb–object pairs—which can be amplified into a model shortcut at training time.
+Under such skewed supervision, predictions for unseen compositions tend to drift toward seen (especially frequent) pairs.
+Second, verbs are harder to learn than objects.
+As shown in Figure 1 (b), an object often becomes recognizable from a single frame, whereas a verb typically requires multi-frame temporal reasoning.
+Prior work on shortcut learning suggests that models tend to rely on easier-to-learn cues rather than the intended evidence [nam2020lff, li2018resound, singh2020dont, geirhos2018cuecon, scimeca2022shortcut].
+In ZS-CAR, this verb–object difficulty gap makes object cues an especially attractive shortcut for verb prediction.
+In this work, we posit that these two intertwined issues in verb-object compositional learning promote object-driven shortcuts, which we diagnose as the key factor hindering ZS-CAR.
+As Figure 1 (c) illustrates, a model can match a verb once it recognizes its frequently co-occurring object during training, without modeling the temporal dynamics essential to correctly recognize the verb.
+This behavior weakens generalization to unseen compositions because it ties verb prediction to object-dependent priors rather than temporally grounded evidence.
+We observe that models [li2024c2c], even with a modern video-pretrained backbone [wang2024iv2], indeed confuse verbs with opposite temporal order (e.g., opening vs.
+closing) on unseen compositions (see Section 5.2).
+To dissect how object-driven shortcuts hinder unseen generalization, we perform a comprehensive quantitative diagnosis.
+First, through controlled experiments, we show that the asymmetric learning difficulty between verbs and objects, together with skewed co-occurrence priors, promotes verb shortcut learning driven by object cues.
+Next, we quantify shortcut-driven failures in existing ZS-CAR models by two diagnostic ratios: False Seen Prediction (FSP) and False Co-occurrence Prediction (FCP), which measure how often an unseen input collapses to a seen—and especially frequent—training composition.
+Using FSP/FCP, we show that a strong ZS-CAR baseline [li2024c2c] remains heavily biased toward seen compositions, resulting in poor generalization to unseen compositions.
+Driven by the diagnosis, we propose Robust COmpositional REpresentations (RCORE), targeting the two root causes of object-driven shortcuts.
+First, Co-occurrence Prior Regularization (CPR) expands supervision over originally absent compositions and suppresses the dominance of frequently seen pairs by treating them as hard negatives.
+Second, Temporal Order Regularization for Composition (TORC) enforces temporal-order sensitivity so that the model learns temporally grounded verb representations instead of relying on static cues.
+We validate RCORE on Sth-com [li2024c2c] under a realistic open-world evaluation protocol that evaluates all possible verb-object pairs, without requiring test-set ground truth labels.
+We also introduce EK100-com, a ZS-CAR dataset repurposed from EPIC-KITCHENS-100 [damen2022ek100], which exhibits more severe compositional sparsity than Sth-com [li2024c2c].
+Across datasets and backbones, RCORE reduces shortcut diagnostics (FSP/FCP) and improves unseen compositional generalization.
 In this work, we make the following major contributions:
-  • Identifying the core failure mode of ZS-CAR.
-We show that existing models rely on object-driven verb shortcuts, caused by the sparsity of compositional supervision and the asymmetric learning difficulty between verbs and objects—an issue not explicitly recognized in prior work.
-  • Diagnosing limitations of the ZS-CAR evaluation protocol.
-We reveal that common practices, such as test-set bias calibration, obscure true compositional generalization.
-To address this, we adopt an open-world unbiased evaluation setting and introduce the compositional gap metric to measure the benefit of compositional reasoning.
-  • A simple yet effective framework that mitigates shortcut learning.
-We propose RCORE, which comprises a composition-aware augmentation and a temporal order regularization loss (TORC) to promote temporally grounded verb representations under compositional sparsity.
-Through experiments on Sth-com and our new EK100-com dataset, we demonstrate that RCORE shows consistently favorable performance improvements across diverse evaluation scenarios, metrics, and datasets.
+  • Shortcut Diagnosis: We provide a diagnostic view of object-driven shortcuts in ZS-CAR.
+We empirically identify strong co-occurrence priors and asymmetric learning difficulty as the root causes of the shortcuts.
+Furthermore, we expose the limitations of baselines by quantifying how unseen inputs collapse into seen-especially frequent-training pairs, using FSP/FCP.
+  • Diagnosis-Driven Framework: We propose RCORE comprising CPR to suppress co-occurrence priors by expanding supervision over originally absent compositions and TORC to promote temporally grounded verb learning.
+  • Effectiveness & Robustness: We demonstrate that RCORE is effective, achieving superior performance on unseen compositions over baselines on Sth-com and EK100-com datasets, with multiple VLM backbones.
+Moreover, through extensive quantitative analyses, we prove that our approach mitigates co-occurrence biases and learns temporally grounded verb features.
 2 Related Work
-Zero-shot compositional action recognition.
-Zero-Shot Compositional Action Recognition (ZS-CAR) [16] aims to generalize to novel verb–object combinations not seen during training.
-Unlike zero-shot (or open-set) action recognition [5, 4, 40], ZS-CAR retains the same verb and object classes across train and test splits and evaluates a model’s ability to compose known primitives.
-It addresses the performance degradation issue on novel unseen combinations of known concepts, which arises from the model’s lack of compositional generalization ability.
-Conditional learning [16, 13] has been effective for estimating joint verb–object probabilities, while other work [14] pursues improved feature disentanglement.
-However, as our analysis reveals in Section 3 and Section 5.2, these approaches remain susceptible to object-driven shortcut behavior and progressively overfit co-occurrence statistics.
-Compositional zero-shot Learning in image understanding.
-Compositional Zero-Shot Learning (CZSL) has been widely studied in the image domain [23, 24, 21, 34, 36, 26].
-CZSL enables models to recognize novel combinations of previously learned primitive concepts, such as objects and their states.
-Prior works have approached this problem through feature disentanglement [29, 42], modular architectures [27], and CLIP-based prompt learning [26, 18].
-Unlike image-based CZSL, ZS-CAR requires modeling temporal dynamics for verbs, making the problem fundamentally different.
-Our work focuses specifically on the video domain and highlights a failure mode—object-driven verb shortcuts—that does not arise in image-only CZSL.
+Compositional Action Recognition (CAR).
+CAR aims to recognize actions by factoring them into verbs and objects.
+A recurring challenge is learning verb representations that are robust to static cues (e.g.
+objects/scenes) and capture temporal dynamics [mat2020sthelse, sun2021cdn, chatterjee2023oap].
+Prior work addresses this by modeling object-level interactions [mat2020sthelse, sun2021cdn] or by training verb representations to be less sensitive to object appearance [chatterjee2023oap].
+Related directions include open-vocabulary CAR, which expands the action label space beyond a fixed taxonomy (often via text-driven semantics) [luo2022dark, chatterjee2023oap].
+In this work, we study Zero-Shot Compositional Action Recognition (ZS-CAR) under a fixed verb/object vocabulary, where verb and object label sets are shared across splits but some verb–object pairs are held out for testing [li2024c2c, jiang2025dhd, jung2025crr, ye2025logic].
+In ZS-CAR, conditional learning [li2024c2c, jiang2025dhd] and disentanglement-oriented designs [jung2025crr] improve joint modeling, yet our study highlights that these models can still exhibit object-driven shortcut behavior under sparse and skewed compositional supervision (see Section 3, Section 5.2).
+Compositional Zero-Shot Learning (CZSL) for image understanding.
+In the image domain, CZSL focuses on recognizing novel compositions of seen primitives (e.g., attribute–object) under sparse compositional supervision [misra2017red, naeem2021czsl, mancini2021compcos, kim2023hierarchical, wu2025cond, nayak2023csp].
+Representative approaches include feature disentanglement [saini2022disen, zhang2024disen], modular factorization [purushwalkam2019taskdriven], and CLIP-based prompt learning [nayak2023csp, li2024context].
+Compared to image CZSL, ZS-CAR additionally requires temporal grounding of verbs, which changes both the failure modes and the regularization opportunities.
+In contrast, we reveal how sparse compositional supervision, coupled with the higher learning difficulty of verbs, amplifies shortcut reliance in videos, and propose training-time regularizers tailored to ZS-CAR.
 Shortcut learning.
-Shortcut learning is a pervasive issue in visual recognition.
-Several studies show that models tend to rely on easy-example biases [35, 15, 30] or easier-to-learn cues—such as color or texture—when these cues correlate with labels [25, 11].
-A line of work attributes this behavior to co-occurrence statistics in the training data [1, 32].
-In video understanding, similar shortcut behaviors occur when action models rely on backgrounds rather than motion [17, 6, 33, 9, 2].
-Building on these observations, we argue that ZS-CAR introduces an additional structural imbalance: verbs require temporal reasoning, whereas objects can be recognized from a single frame.
-This learning asymmetry amplifies shortcut tendencies and leads models to predict verbs primarily from object cues—a failure mode we diagnose and directly address in this work.
+Models can learn shortcuts by exploiting spurious correlations that are easier to fit than the intended evidence [geirhos2018cuecon, ilyas2019adversarial, nam2020lff, scimeca2022shortcut].
+Such shortcuts are closely tied to skewed training distributions and co-occurrence statistics, which can induce reliance on dataset priors [agrawal2018dont, singh2020dont].
+In compositional settings, prior work reweights rare combinations to mitigate co-occurrence bias [kim2023hierarchical].
+In video understanding, several works report reliance on static cues such as scenes [whycantchoi, wang2021be, ding2022fame, bae2024devias] and objects [bahng2019rebias, chatterjee2023oap] instead of motion-sensitive evidence.
+Building on these observations, we study a ZS-CAR-specific manifestation: object-driven shortcuts where object cues dominate verb prediction under sparse verb–object supervision, and we quantify this behavior with dedicated diagnostics.
+Vision-Language Model (VLM).
+Recent VLMs broadly follow two paradigms: encoder-based and decoder-based.
+CLIP [radford2021clip] is a representative encoder-only VLM pretrained via large-scale contrastive learning, and it exhibits strong zero-shot recognition [radford2021clip] as well as non-trivial compositional behavior [abbasi2024decipher].
+Following CLIP, recent works pretrain encoder-based video VLMs on large-scale video-text data [ma2022xclip, li2023unmasked, zhao2024videoprism, wang2024iv2].
+Decoder-based VLMs (e.g., LLaVA [liu2023llava], Qwen-VL [bai2025qwen25vl]) couple a vision encoder with an LLM decoder and excel at open-ended generation.
+In ZS-CAR, however, inference and evaluation operate over a fixed verb–object label space with explicit compositional scores (i.e., logits over $\mathbb{Y}^{V}\times\mathbb{Y}^{O}$), which provides a controlled interface for diagnosis and for designing training-time regularizers that are transferable across backbones.
+Such decoder-based VLM interfaces often do not expose, in a standardized and reproducible way, explicit scores over the fixed label space, as they primarily provide language outputs at inference time.
+Therefore, we focus on encoder-based VLMs in this controlled setting, and show that object-driven shortcuts persist even with a video-pretrained backbone [wang2024iv2] when paired with a strong ZS-CAR baseline [li2024c2c]; we then propose RCORE to regularize compositional scoring to mitigate these shortcuts.
 3 Diagnosis: Why ZS-CAR Models Fail?
 3.1 Zero-Shot Compositional Action Recognition
-We study Zero-Shot Compositional Action Recognition (ZS-CAR) [16], where the goal is to recognize verb–object compositions that might never appear in the training set.
-Let $\mathbb{Y}^{V}$ and $\mathbb{Y}^{O}$ denote the sets of verb and object labels, and let each action label be a pair $\mathbf{y}=(\mathbf{y}^{V},\mathbf{y}^{O})\in\mathbb{Y}^{V}\times\mathbb{Y}^{O}$.
-The training set is $\mathbb{D}_{\text{train}}=\{(\mathbf{X}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$, and its set of observed (seen) compositions is $\mathbb{Y}_{\text{seen}}=\{\mathbf{y}_{i}\mid(\mathbf{X}_{i},\mathbf{y}_{i})\in\mathbb{D}_{\text{train}}\}$.
+We study Zero-Shot Compositional Action Recognition (ZS-CAR) [li2024c2c], where the goal is to recognize unseen verb–object compositions that are held out from the training set, while the verb and object vocabularies are fixed and shared across train/validation/test splits.
+Thus, unlike open-set or open-vocabulary settings, ZS-CAR does not introduce unseen verbs or objects at test time—only unseen compositions (verb–object pairs).
+Let $\mathbb{Y}^{V}$ and $\mathbb{Y}^{O}$ denote the sets of verb and object labels, and let each composition label be a pair $\mathbf{y}^{C}=(\mathbf{y}^{V},\mathbf{y}^{O})\in\mathbb{Y}^{V}\times\mathbb{Y}^{O}$.
+The training set is $\mathbb{D}_{\text{train}}=\{(\mathbf{X}_{i},\mathbf{y}^{C}_{i})\}_{i=1}^{N}$, and its set of observed (seen) compositions is $\mathbb{Y}_{\text{seen}}=\{\mathbf{y}^{C}_{i}\mid(\mathbf{X}_{i},\mathbf{y}^{C}_{i})\in\mathbb{D}_{\text{train}}\}$.
 The space of unseen compositions is $\mathbb{Y}_{\text{unseen}}=(\mathbb{Y}^{V}\times\mathbb{Y}^{O})\setminus\mathbb{Y}_{\text{seen}}$.
-We define the validation and test sets analogously.
 The validation and test sets contain both seen and unseen compositions, enabling evaluation of compositional generalization.
-ZS-CAR therefore measures whether a model can correctly compose familiar verbs and objects into novel pairs at test time.
+Figure 2: Controlled experiments demonstrate object-driven shortcut.
+(a) Objects are easier than verbs.
+A randomly initialized ViT trained on a balanced $10{\times}10$ Sth-com subset learns objects substantially faster than verbs.
+(b) Co-occurrence bias induces object-driven shortcuts.
+On a perfectly biased training split, CLIP (with AIM [yang2023aim]) achieves high object accuracy, but verb accuracy on bias-conflict unseen compositions drops below chance, indicating shortcut-driven verb failures.
 3.2 Diagnostic Metrics
-For better understanding the underlying behavior of ZS-CAR models, we present two diagnostic metrics: (1) training bias metrics to measure a model’s overfitting and (2) Compositional Gap to assess the benefits of compositional modeling.
-Training bias metrics
-To measure how the model overfits to the training set statistics, we introduce two types of misclassification ratios: False Seen Prediction and False Co-occurrence Prediction ratios.
-False Seen Prediction (FSP) ratio denotes the ratio of unseen composition samples incorrectly predicted to seen composition labels.
-Out of these mispredicted samples, False Co-occurrence Prediction (FCP) ratio is likewise defined as the proportion of those assigned to frequently co-occurring labels.
-Here, a composition label is considered ‘frequently co-occurring’ if its occurrence is higher than one standard deviation above the mean frequency in the training data.
-Compositional Gap
-We introduce Compositional Gap to examine whether a model genuinely benefits from understanding videos in a compositional manner. Compositional Gap ($\Delta_{\text{CG}}$) is defined as
-$\Delta_{\text{CG}}=\mathrm{Acc}^{C}-(\mathrm{Acc}^{V}\times\mathrm{Acc}^{O}),$ (1)
-where $\text{Acc}^{C}$ is the composition accuracy (%) computed from the logits over the composition space, whereas $\text{Acc}^{V}$ and $\text{Acc}^{O}$ are the verb and object accuracies (%) derived from the logits over their individual component spaces.
-A positive $\Delta_{\text{CG}}$ indicates that the model truly gains an advantage from its compositional architecture.
-Figure 3: Learning curve of the SOTA model with our diagnostic metrics.
-We plot the learning curve of C2C [16] trained on Sth-com [16].
-We measure the False Seen Prediction (FSP) and False Co-occurrence Prediction (FCP) ratios, and observe that the seen–unseen accuracy gap ($\Delta_{SU}$) correlates strongly with both metrics.
-There observations indicate that the current SOTA model exhibits overfitting to seen compositions.
+To better understand ZS-CAR model behavior, we introduce two diagnostics: (1) training-bias metrics that quantify over-reliance on training co-occurrence statistics, and (2) the Compositional Gap that measures the benefit of modeling verb–object dependence beyond independent recognition.
+Training-bias metrics.
+To quantify reliance on co-occurrence priors, we define two misclassification ratios on the unseen subset of an evaluation split $\mathbb{Y}_{\text{unseen}}$.
+False Seen Prediction (FSP) is the fraction of misclassified unseen composition samples whose predictions $\hat{\mathbf{y}}^{C}$ fall into the seen composition categories $\mathbb{Y}_{\text{seen}}$.
+False Co-occurrence Prediction (FCP) is the fraction of misclassified unseen composition samples whose prediction $\hat{\mathbf{y}}^{C}$ falls into the frequent compositions $\mathbb{Y}_{\text{freq}}$.
+Let $f(\mathbf{y}^{C})$ denote the training frequency of a seen composition $\mathbf{y}^{C}\in\mathbb{Y}_{\text{seen}}$.
+We define the set of frequent seen compositions $\mathbb{Y}_{\text{freq}}$ as
+$\vskip-3.00003pt\mathbb{Y}_{\text{freq}}=\{\mathbf{y}^{C}\in\mathbb{Y}_{\text{seen}}\mid f(\mathbf{y}^{C})>\mu_{f}+\sigma_{f}\},\vskip-3.00003pt$ (1)
+where $\mu_{f}$ and $\sigma_{f}$ are the mean and standard deviation of $\{f(\mathbf{y}^{C})\}_{\mathbf{y}^{C}\in\mathbb{Y}_{\text{seen}}}$.
+To further analyze shortcut patterns, we decompose FSP/FCP into three component-wise cases based on prediction errors: (i) Verb-collapse, where the predicted object is correct but the verb is incorrect; (ii) Object-collapse, where the predicted verb is correct but the object is incorrect; and (iii) Dual-collapse, where both verb and object are incorrect.
+Using the metrics, we can identify whether co-occurrence bias primarily harms verb or object learning.
+Compositional Gap.
+We introduce Compositional Gap ($\Delta_{\text{CG}}$) to examine whether a model benefits from predicting the joint verb–object composition beyond getting the two components correct independently.
+Given joint logits over $\mathbb{Y}^{V}\times\mathbb{Y}^{O}$, we obtain the predicted composition $\hat{\mathbf{y}}^{C}=(\hat{\mathbf{y}}^{V},\hat{\mathbf{y}}^{O})$ by selecting the top-1 composition label.
+We then compute top-1 accuracies $\mathrm{Acc}^{V}$, $\mathrm{Acc}^{O}$, and $\mathrm{Acc}^{C}$ respectively.
+Then we define $\Delta_{\text{CG}}$ as
+$\vskip-3.00003pt\Delta_{\text{CG}}=\mathrm{Acc}^{C}-(\mathrm{Acc}^{V}\times\mathrm{Acc}^{O}).\vskip-3.00003pt$ (2)
+We report $\Delta_{\text{CG}}$ alongside FSP/FCP and unseen composition accuracy as a diagnostic metric, rather than interpreting it as a standalone evaluation criterion.
+(a) CLIP
+(b) InternVideo2
+(c) FSP/FCP breakdown
+Figure 3: Shortcut reliance correlates with poor unseen generalization.
+We show learning curves of C2C [li2024c2c] on Sth-com with CLIP [radford2021clip] and InternVideo2 [wang2024iv2].
+(a) With CLIP, the seen–unseen accuracy gap grows together with FSP/FCP, indicating strong overfitting to seen compositions.
+(b) Even with the video-pretrained InternVideo2 backbone, the model still shows a notable seen–unseen gap and high FSP.
+(c) FSP/FCP breakdown at the final epoch: CLIP is dominated by verb-collapse due to object-driven shortcuts, while InternVideo2 exhibits milder but persistent shortcut behavior.
 Best viewed with zoom and color.
+3.3 Empirical Diagnosis of Object-driven Shortcuts in ZS-CAR
+In this section, we empirically diagnose object-driven shortcut behavior in ZS-CAR and analyze how it correlates with weak verb learning and poor generalization to unseen compositions.
+We show that this behavior persists even with a modern video-pretrained VLM backbone [wang2024iv2], and provide additional evidence in Section 5 and Section 0.D.
+Objects are easier to learn than verbs.
+To compare the learning difficulty of verbs and objects, we train a randomly initialized ViT [dosovitskiy202vit] on a controlled subset of Sth-com [li2024c2c].
+Specifically, we construct a balanced $10\times 10$ subset containing all 100 verb–object pairs to isolate learning dynamics from compositional sparsity.
+Based on prior observations that models tend to fit easier cues earlier under limited supervision [nam2020lff, scimeca2022shortcut], we compare the learning curves of verb vs.
+object prediction.
+In Figure 2 (a), the model learns objects faster and reaches higher accuracy than verbs under this controlled setting, suggesting that object prediction is easier than verb prediction in our ZS-CAR setup.
+Object-driven shortcuts indeed exist in ZS-CAR.
+Under skewed co-occurrence statistics, models may rely on easier cues as shortcuts instead of learning the intended evidence [nam2020lff, scimeca2022shortcut, geirhos2018cuecon, bahng2019rebias, li2018resound, whycantchoi].
+We test whether CLIP exhibits such object-driven shortcuts by constructing a bias-controlled training split where each verb is strongly correlated with a particular object (see construction details in Section 0.C), and evaluating on two splits: (i) bias-aligned seen compositions and (ii) bias-conflict unseen compositions.
+As shown in Figure 2 (b), CLIP achieves high object accuracy on unseen compositions while verb accuracy on unseen compositions drops below the uniform-chance level ($1/|\mathbb{Y}^{V}|$), indicating that object cues dominate verb prediction under this biased supervision.
+We observe this trend even with random initialization, implying shortcuts stem from the training distribution and objective, not just pretraining.
+See Section 0.C for more details.
+Shortcut reliance correlates with poor unseen generalization.
+Next, we test whether shortcut reliance correlates with weak unseen generalization in a standard SOTA pipeline.
+Figure 3 plots validation accuracies together with FSP/FCP for C2C [li2024c2c] on Sth-com [li2024c2c].
+As shown in Figure 3 (a), with a CLIP backbone [radford2021clip], increasing FSP/FCP accompanies a widening seen–unseen gap, consistent with growing reliance on co-occurrence statistics.
+Our error breakdown in Figure 3 (c) further shows that, among unseen samples that collapse to seen pairs, a large fraction corresponds to verb-collapse, supporting the presence of object-driven shortcut behavior.
+As shown in Figure 3 (b) and (c), with an InternVideo2 [wang2024iv2] backbone, the trends are more stable; however, FSP remains high, and the FCP breakdown still exhibits substantial verb-collapse, suggesting that stronger video pre-training alone is insufficient to eliminate the tendency to overfit to seen pairs via co-occurrence-driven shortcuts.
+Table 1: $\Delta_{\mathrm{CG}}{}$ on Sth-com.
+Backbone Split $\Delta_{\text{CG}}$
+CLIP Seen +3.24
+Unseen -0.42
+InternVideo2 Seen +2.24
+Unseen -0.63
+$\Delta_{\text{CG}}$ indicates degraded compositional behavior on unseen compositions.
+We report $\Delta_{\text{CG}}$ of C2C [li2024c2c] with CLIP [radford2021clip] and InternVideo2 [wang2024iv2] backbones.
+As shown in Table 1, the model exhibits negative $\Delta_{\text{CG}}$ on unseen compositions even with the video-pretrained backbone, indicating that $\mathrm{Acc}^{C}$ falls below the independent reference $\mathrm{Acc}^{V}\times\mathrm{Acc}^{O}$.
+Together with the training-bias metrics (FSP/FCP), these results support that shortcut reliance is closely associated with weak unseen compositional generalization in current ZS-CAR pipelines.
+4 RCORE
+We introduce RCORE, a diagnosis-driven learning framework that mitigates object-driven shortcuts in ZS-CAR.
+RCORE improves compositional generalization by strengthening verb–object representations under sparse supervision through two complementary components: (i) Co-occurrence Prior Regularization (CPR) expands supervision over novel compositions and leverages frequently seen pairs as hard negatives to suppress co-occurrence priors, and (ii) Temporal Order Regularization for Composition (TORC) loss promotes temporally grounded verb learning over static object cues.
+Following prior work [li2024c2c, jung2025crr], we employ an adapter-based tuning, AIM [yang2023aim] for CLIP [radford2021clip] and LoRA [hu2022lora] for InternVideo2 [wang2024iv2], as our backbone.
+Figure 4 (a) gives an overview of RCORE.
+Feature extraction.
+Given a video $\mathbf{X}$ with $T$ frames, the backbone encoder outputs frame-level features $\mathbf{F}\in\mathbb{R}^{T\times D}$.
+We transform them into verb features $\mathbf{F}^{V}\in\mathbb{R}^{T\times D}$ and object features $\mathbf{f}^{O}\in\mathbb{R}^{D}$ using dedicated encoders.
+We obtain text embeddings for verbs and objects, $\mathbf{E}^{V}\in\mathbb{R}^{|\mathbb{Y}^{V}|\times D}$ and $\mathbf{E}^{O}\in\mathbb{R}^{|\mathbb{Y}^{O}|\times D}$, by feeding class-specific prompts into the text encoder.
+4.1 CPR: Co-occurrence Prior Regularization
+To mitigate object-driven shortcuts under sparse and skewed compositional supervision, we propose CPR, a training strategy that injects supervision for synthesized verb–object pairs while preserving the stability of closed-world optimization.
+CPR combines (i) synthesized composition supervision, (ii) a margin-based regularizer that down-weights frequently seen hard negatives, and (iii) a batch-adaptive expansion of the composition label space to make these new pairs trainable without optimizing over the full $\mathbb{Y}^{V}\times\mathbb{Y}^{O}$ space.
 Figure 4: Overview of RCORE.
 (a) Overview of our proposed RCORE framework.
-(b) VOCAMix synthesizes plausible yet unseen verb–object compositions while preserving the temporal structure of the primary video.
+(b) CPR synthesizes plausible yet unseen verb–object compositions and penalizes frequently seen hard negatives by a margin-based regularizer.
 (c) TORC penalizes alignment between original and temporally perturbed feature vectors, enforcing explicit temporal order modeling and reducing object-driven shortcuts.
-3.3 Empirical Diagnosis of Object-driven Shortcuts in ZS-CAR
-In this section, we systematically diagnose the key failure mode in ZS-CAR—object-driven shortcuts.
-We demonstrate that these shortcuts prevent ZS-CAR models from learning robust verb representations, consequently leading to poor generalization to unseen compositions.
-We present additional evidence of object-driven shortcut learning in ZS-CAR models in Section 5 and the supplementary material.
-Objects are easier to learn than verbs.
-To investigate the learning difficulty of verbs and objects, we train a randomly initialized model on a subset of Sth-com [16].
-Based on the observation that the easier-to-learn cue is captured earlier and exhibits higher accuracy [25], we can analyze the learning difficulty of each component.
-In Figure 2 (a), the results show a clear trend: the model learns objects faster and consistently achieves higher accuracy for objects than for verbs.
-While intuitively expected, these results empirically verify that objects are inherently easier to learn than verbs in compositional action recognition.
-Object-driven shortcuts in ZS-CAR
-Under strong co-occurrence priors, models choose shortcuts emerged from easier-to-learn cues, rather than learn robust representations [25, 30, 11, 3, 17, 6].
-We argue this shortcut learning behavior accelerates in ZS-CAR due to the asymmetric learning difficulty.
-To verify this, inspired by prior work [30], we train models on a perfectly biased training set and test on two distinct splits—bias-aligned seen compositions and bias-conflict unseen compositions— utilizing Sth-com [16].
-As shown in Figure 2 (b), in the both cases of the randomly initialized ViT [10] and CLIP [28], the object accuracies for unseen compositions increased as the seen composition accuracies also increased, while the verb accuracies for unseen compositions drop under the random accuracy.
-Paradoxically, CLIP [16], well-known for its superior representational power, exhibits the highest object accuracy (94.0%) and the lowest verb accuracy (1.9%) concurrently on unseen compositions.
-These results demonstrate models tend to fall into object-driven shortcuts to exploit co-occurrence priors in the training data, resulting in poor verb representations.
-Object-driven shortcuts directly harm unseen generalization.
-Next, we investigate whether the state-of-the-art (SOTA) model also fails to generalize to unseen compositions as a consequence of these object-driven shortcuts.
-As shown in Figure 3, we trace the validation accuracies along with FSP and FCP ratios—metrics designed to quantify the model’s reliance on the training statistics—of C2C [16] throughout the training progress on Sth-com [16].
-The clear correlation between increased FSP/FCP ratios and widen seen-unseen accuracy gap indicates that the model overfits to the training statistics.
-Table 1: The SOTA model exhibits a negative Compositional Gap on unseen compositions.
-We measure Compositional Gap ($\Delta_{\mathrm{CG}}$) of the SOTA ZS-CAR model.
-We employ C2C [16] trained on Sth-com [16] with 30 epochs and evaluate the model in open-world setting on the Sth-com [16] test set.
-The model shows a negative $\Delta_{\mathrm{CG}}$ on unseen compositions, indicating its composition accuracy falls below the level expected from independent verb–object predictions.
-Split $\mathrm{Acc}^{V}$ $\mathrm{Acc}^{O}$ $\mathrm{Acc}^{V}\times\mathrm{Acc}^{O}$ $\mathrm{Acc}^{C}$ $\Delta_{\text{CG}}$
-Seen 63.60 67.72 43.07 46.31 +3.24
-Unseen 54.36 56.10 30.50 30.08 -0.42
-Compositional Gap reveals the SOTA model fails to benefit from compositional modeling.
-We evaluate the Compositional Gap ($\Delta_{\text{CG}}$) performance of current SOTA ZS-CAR model [16].
-As shown in Table 1, the model shows a negative $\Delta_{\text{CG}}$ on unseen compositions in Sth-com [16].
-This demonstrates that, without addressing these shortcuts, the SOTA model gains no advantage from compositional modeling compared to independent modeling for unseen compositions.
-4 RCORE
-We introduce RCORE, a learning framework that mitigates object-driven shortcut learning in zero-shot compositional action recognition.
-RCORE improves compositional generalization by strengthening verb–object representations under sparse supervision through two complementary components: (i) VOCAMix, a composition-aware augmentation that expands the variety of verb–object combinations, and (ii) TORC, a temporal order regularization loss that promotes temporally grounded verb learning over static object cues.
-Following prior work [16], we use AIM [37], a CLIP-based [28] parameter-efficient spatio-temporal encoder, as our backbone.
-Figure 4 (a) gives an overview of RCORE.
-Feature extraction.
-Given an augmented video $\hat{\mathbf{X}}$ with $T$ frames, the backbone encoder outputs frame-level features $\mathbf{f}_{t}\in\mathbb{R}^{D}$.
-We then transform them into verb features $\mathbf{F}^{V}\in\mathbb{R}^{T\times D}$ and an object features $\mathbf{f}^{O}\in\mathbb{R}^{D}$ using dedicated verb and object encoders.
-We obtain text embeddings for verbs and objects, $\mathbf{e}^{V}$ and $\mathbf{e}^{O}$, by feeding class-specific prompts (e.g., “a photo of {verb/object}) into the text encoder.
-4.1 Composition-Aware Augmentation
-To alleviate the sparsity of compositional training data, we propose Verb–Object Composition-Aware Augmentation (VOCAMix), which synthesizes plausible but unseen combinations while preserving the temporal structure of the primary video.
-As shown in Figure 4 (b), VOCAMix replaces the motion region of a video with a static object region from another sample in the mini-batch.
-Given a primary video $\mathbf{X}_{i}$ with label $(\mathbf{y}_{i}^{V},\mathbf{y}_{i}^{O})$ and a randomly selected video $\mathbf{X}_{j}$ with object label $\mathbf{y}_{j}^{O}$ ($i\neq j$), we construct an augmented video $\hat{\mathbf{X}}_{i}$ and a soft object label:
-$h(\hat{\mathbf{X}}_{i}(k,:))$ $=(1-\lambda)\,h(\mathbf{X}_{i}(k,:))+\lambda\,h(\mathbf{X}_{j}(\lfloor T/2\rfloor,:)),$ (2)
-$\tilde{\mathbf{y}}^{O}_{i}$ $=(1-\lambda)\,\mathbf{y}_{i}^{O}+\lambda\,\mathbf{y}_{j}^{O},$ (3)
-where $h(\cdot)$ is a high motion-region estimation function, $k\in\{1,\ldots,T\}$ is the frame index, and $\lambda\in[0,1]$ controls the strength of object injection.
-We adopt a simple, learning-free estimator [9] for $h(\cdot)$.
-We keep the verb label $\mathbf{y}_{i}^{V}$ unchanged to preserve temporal semantics.
-By injecting a consistent static object across the temporal axis, VOCAMix preventing disruption of temporal cues [33] while expanding compositional diversity.
-while synthesizing novel compositions.
-4.2 Temporal Order Regularization Loss
-To counter object-driven shortcuts in verb learning, we introduce the Temporal Order Regularization for Composition (TORC) loss.
-TORC loss forces the model to rely on temporal structure: it penalizes alignment with temporally incorrect feature sequences and suppresses confident verb predictions when temporal ordering is corrupted.
-This regularization steers the model away from static object cues and strengthens temporally grounded verb representations.
-Figure 5: Analysis on the effects of RCORE on the Sth-com [16] dataset.
-(a) RCORE prevents the False Co-occurrence Prediction (FCP) ratio from increasing during training, whereas the baseline shows a clear rise in FCP.
-As a result, RCORE consistently maintains a smaller seen–unseen accuracy gap ($\Delta_{SU}$) throughout training.
-(b) The cosine similarity between the original and reversed verb features becomes strongly negative for RCORE as training progresses, indicating improved temporal discriminative capability.
-In contrast, the baseline maintains a high similarity (0.91), revealing limited temporal sensitivity.
-(c) On the Temporal subset, RCORE exhibits a substantially larger performance gap between original and temporally shuffled features compared to the baseline, demonstrating that RCORE learns verb representations that depend on temporal dynamics rather than static object cues.
-Best viewed with zoom and color.
-Temporal perturbation.
-Given frame-level features $\mathbf{F}=(\mathbf{f}_{1},\dots,\mathbf{f}_{T})$, we construct two perturbed variants: (i) a reversed sequence $\mathbf{F}_{\text{rev}}=(\mathbf{f}_{T},\dots,\mathbf{f}_{1})$, and (ii) a temporally shuffled sequence $\mathbf{F}_{\text{shuffled}}=\pi(\mathbf{F})$, where $\pi$ is a random permutation sampled from $\mathcal{P}$.
+Synthesized composition supervision.
+To provide explicit supervision for compositions absent from the training set, we first synthesize videos that represent new composition labels.
+Given a training sample $\mathbf{X}_{i}$ with composition label $\mathbf{y}_{i}^{C}=(\mathbf{y}_{i}^{V},\mathbf{y}_{i}^{O})$, we construct a new video $\tilde{\mathbf{X}}_{i}$ by injecting the static object cue from another video $\mathbf{X}_{j}$ (with $\mathbf{y}_{j}^{O}\neq\mathbf{y}_{i}^{O}$) into the high-motion regions of $\mathbf{X}_{i}$, as shown in Figure 4 (b).
+Specifically, for each frame $k\in\{1,\ldots,T\}$ we obtain
+$\vskip-3.00003pt\tilde{\mathbf{X}}_{i}^{k}$ $=\bigl(\mathbf{1}-\lambda\,\mathbf{M}_{i}^{k}\bigr)\odot\mathbf{X}_{i}^{k}\;+\;\bigl(\lambda\,\mathbf{M}_{i}^{k}\bigr)\odot\mathbf{X}_{j}^{\lfloor T/2\rfloor},\vskip-3.00003pt$ (3)
+where $\lambda\in[0,1]$ controls the injection strength and $\mathbf{M}_{i}^{(k)}\in\{0,1\}^{H\times W}$ is the high-motion region mask for frame $k$, extracted by a learning-free estimator [ding2022fame].
+The operator $\odot$ denotes element-wise multiplication, with $\mathbf{M}_{i}^{k}$ broadcast across channels.
+Accordingly, we use soft labels for the synthesized video as follows:
+$\vskip-3.00003pt\tilde{\mathbf{y}}^{O}_{i}$ $=(1-\lambda)\,\mathbf{y}_{i}^{O}+\lambda\,\mathbf{y}_{j}^{O},$ (4)
+$\tilde{\mathbf{y}}^{C}_{i}$ $=(1-\lambda)\,\phi(\mathbf{y}_{i}^{V},\mathbf{y}_{i}^{O})+\lambda\,\phi(\mathbf{y}_{i}^{V},\mathbf{y}_{j}^{O}),\vskip-3.00003pt$ (5)
+where $\phi:\mathbb{Y}^{V}\times\mathbb{Y}^{O}\rightarrow\{0,1\}^{|\mathbb{Y}^{V}\times\mathbb{Y}^{O}|}$ maps a composition $(\mathbf{y}^{V},\mathbf{y}^{O})$ to its one-hot vector in the flattened joint label space.
+Co-occurrence prior regularization loss.
+Even after adding supervision for synthesized compositions, skewed co-occurrence statistics can still make the model over-score a small set of frequently seen pairs.
+To counter this score collapse, we enforce a margin constraint that makes the target synthesized composition $\mathbf{y}^{C}_{i,\text{new}}=(\mathbf{y}_{i}^{V},\mathbf{y}_{j}^{O})$ outrank frequent seen hard negatives by at least $m$.
+We define the hard-negative label set for $\mathbf{y}^{C}_{i,\text{new}}$ as follows:
+$\vskip-5.0pt\mathbb{Y}_{\text{hard},i}=\left\{(\mathbf{y}^{V},\mathbf{y}^{O})~\middle|~\big[(\mathbf{y}^{V}=\mathbf{y}_{i}^{V})\oplus(\mathbf{y}^{O}=\mathbf{y}_{j}^{O})\big]\setminus(\mathbf{y}^{V},\mathbf{y}^{O})\neq(\mathbf{y}_{i}^{V},\mathbf{y}_{i}^{O})\right\},\vskip-5.0pt$ (6)
+where $\oplus$ denotes the exclusive or operator.
+Let $\mathbb{Y}_{\text{freq}}\subset\mathbb{Y}_{\text{seen}}$ be the set of frequently seen compositions (details about the criterion in Section 0.C.2).
+Then we define the CPR loss with margin $m$ over the penalty set $\mathbb{Y}_{\text{pen}}=\mathbb{Y}_{\text{hard},i}\cap\mathbb{Y}_{\text{freq}}$ as:
+$\vskip-20.00003ptL_{\text{CPR}}=\sum_{\mathbf{y}^{C}\in\mathbb{Y}_{\text{pen}}}\max\Big(0,\;s(\mathbf{y}^{C})-s(\mathbf{y}^{C}_{i,\text{new}})+m\Big),\vskip-20.00003pt$ (7)
+where $s(\mathbf{y}^{C})$ denotes the logit assigned to the composition label $\mathbf{y}^{C}$.
+Batch-adaptive label-space expansion.
+Conventional ZS-CAR pipelines optimize classification only over seen compositions $\mathbb{Y}_{\text{seen}}$ (closed-world over compositions), which makes it non-trivial to directly supervise novel verb–object pairs during training.
+On the other hand, naively applying cross-entropy over the full $\mathbb{Y}^{V}\times\mathbb{Y}^{O}$ space can harm unseen accuracy by treating most unseen compositions as negatives throughout training (see Section 5.4).
+To provide an effective supervision for novel compositions, we introduce a batch-adaptive label space expansion strategy.
+For each mini-batch, we construct an expanded label set as $\mathbb{Y}_{\text{exp}}=\mathbb{Y}_{\text{seen}}\cup\mathbb{Y}_{\text{new}}$, where $\mathbb{Y}_{\text{new}}$ denotes the label set for the newly synthesized videos.
+We compute cross-entropy only over $\mathbb{Y}_{\text{exp}}$.
+This preserves the optimization stability of closed-world training while progressively injecting supervision for new compositions.
+4.2 TORC: Temporal Order Regularization for Composition
+To counter object-driven shortcuts in verb learning, we introduce TORC, a temporal order-sensitivity regularizer.
+While CPR expands supervision over compositions, verb recognition can still collapse to static object cues due to the asymmetric learning difficulty between verbs and objects.
+As shown in Figure 4 (c), TORC directly regularizes the verb representation to depend on temporal structure: it (i) discourages invariance to temporal reversal by separating forward and reversed features, and (ii) penalizes over-confident verb predictions when temporal order is disrupted by encouraging high-entropy outputs.
+Together, these constraints reduce reliance on static cues and promote temporally grounded verb representations.
+Temporal perturbation for regularization.
+Given frame-level features $\mathbf{F}=(\mathbf{f}_{1},\dots,\mathbf{f}_{T})$, we form two regularization views: (i) a reversed sequence $\mathbf{F}_{\text{rev}}=(\mathbf{f}_{T},\dots,\mathbf{f}_{1})$, and (ii) a temporally shuffled sequence $\mathbf{F}_{\text{shuffled}}=\pi(\mathbf{F})$, where $\pi$ is a random permutation sampled from $\mathcal{P}$.
 We then feed $\mathbf{F}_{\text{rev}}$ and $\mathbf{F}_{\text{shuffled}}$ into the verb encoder to obtain perturbed verb features $\mathbf{f}^{V}_{\text{rev}}\in\mathbb{R}^{D}$ and $\mathbf{f}^{V}_{\text{shuffled}}\in\mathbb{R}^{D}$.
@@
 We first push the model to distinguish forward and reversed temporal semantics.
-We minimize the cosine similarity between the original verb feature $\mathbf{f}^{V}$ and the reversed feature $\mathbf{f}^{V}_{\text{rev}}$:
-$L_{\cos}=\cos(\mathbf{f}^{V},\mathbf{f}^{V}_{\text{rev}}),\vskip-5.0pt$ (4)
-where $\cos(\mathbf{a},\mathbf{b})=\frac{\mathbf{a}^{\top}\mathbf{b}}{\|\mathbf{a}\|\|\mathbf{b}\|}$.
+We minimize the cosine similarity between the original verb $\mathbf{f}^{V}$ and the reversed feature vectors $\mathbf{f}^{V}_{\text{rev}}$: $L_{\cos}=\frac{\mathbf{f^{V}}^{\top}\mathbf{f}^{V}_{\text{rev}}}{\|\mathbf{f}^{V}\|\|\mathbf{f}^{V}_{\text{rev}}\|}$.
 Temporal reversal often flips action meaning (e.g., opening vs.
-closing), yet the state-of-the-art model [16] yields a high cosine similarity ($0.91$; see Figure 5), indicating weak temporal discrimination.
+closing), yet the state-of-the-art model [li2024c2c] yields a high cosine similarity ($0.91$; see Figure 13), indicating weak temporal discrimination.
 Next, we regularize the model to avoid confident verb predictions when temporal structure is disrupted.
-Given the set of verb text embeddings $\mathbb{E}^{V}=\{\mathbf{e}^{V}_{i}\}_{i=1}^{|\mathbb{Y}^{V}|}\in\mathbb{R}^{|\mathbb{Y}^{V}|\times D}$, we define the negative entropy loss $L_{\text{ent}}$ as:
-$L_{\text{ent}}$ $=\sum_{i=1}^{|\mathbb{Y}^{V}|}p_{i}\log p_{i},\vskip-5.0pt$ (5)
-where $p_{i}$ denotes the predicted probability of the $i$-th verb class based on the cosine similarity between visual and textual features, and $\tau$ is a temperature parameter.
+Given the verb text embeddings $\mathbf{E}^{V}=\{\mathbf{e}^{V}_{m}\}_{m=1}^{|\mathbb{Y}^{V}|}$, we define the negative entropy loss $L_{\text{ent}}=\sum_{m=1}^{|\mathbb{Y}^{V}|}p_{m}\log p_{m}$ with the predicted probability $p_{m}$ of the $m$-th verb class based on the cosine similarity between visual and textual features as:
+$\vskip-20.00003ptp_{m}$ $=\frac{\exp(\cos(\mathbf{f}_{\text{shuffled}},\mathbf{e}^{V}_{m})/\tau)}{\sum_{n=1}^{|\mathbb{Y}^{V}|}\exp(\cos(\mathbf{f}_{\text{shuffled}},\mathbf{e}^{V}_{n})/\tau)},\vskip-20.00003pt$ (8)
+where $\tau$ is a temperature parameter.
 Maximizing entropy prevents the model from inferring verbs from static spatial cues alone, enforcing reliance on temporal evidence.
 We define the TORC loss as: $L_{\text{TORC}}=L_{\cos}+L_{\text{ent}}$.
+4.3 Training
 Component loss.
-We compute verb/object logits using cosine similarity between visual features and their corresponding text features.
+We compute verb logits $\mathbf{S}_{V}\in\mathbb{R}^{|\mathbb{Y}^{V}|}$ and object logits $\mathbf{S}_{O}\in\mathbb{R}^{|\mathbb{Y}^{O}|}$ based on the cosine similarity scores.
 We apply standard cross-entropy losses for verb and object prediction: $L_{\text{com}}=L_{V}+L_{O}$.
 Composition loss.
-Following prior work [16], we combine marginal scores $(\mathbf{p}_{V},\mathbf{p}_{O})$ with conditional logits $(\mathbf{p}_{V|O},\mathbf{p}_{O|V})$ to obtain the composition prediction $\hat{\mathbf{y}}$ as $\hat{\mathbf{y}}=\texttt{Softmax}(\mathbf{p}_{O}\mathbf{p}_{V|O}+\mathbf{p}_{V}\mathbf{p}_{O|V})$.
-We employ the cross-entropy loss: $L_{\text{comp}}=L_{\text{CE}}(\hat{\mathbf{y}},\mathbf{y})$.
-Margin loss: mitigating co-occurrence bias.
-Highly frequent verb–object pairs often dominate predictions, even for unseen compositions.
-To counter this, we introduce a margin loss that penalizes frequent-but-incorrect compositions whose logits approach the ground-truth logit:
-$L_{M}=\sum_{c\in\mathcal{C}_{\text{margin}}}\max\left(0,\log\hat{\mathbf{y}}(c)-\log\hat{\mathbf{y}}(g)+m\right),$ (6)
-where $g\in\{1,\dots,|\mathbb{Y}^{V}|\times|\mathbb{Y}^{O}|\}$ denotes the ground-truth index, and $\mathcal{C}_{\text{margin}}$ is the intersection of top-$K$ incorrect predictions and the set of frequently co-occurring compositions.
-This term explicitly discourages reliance on dataset co-occurrence priors.
+Following prior work [li2024c2c], we obtain the composition logits $\mathbf{S}_{C}\in\mathbb{R}^{|\mathbb{Y}^{V}\times\mathbb{Y}^{O}|}$ by aggregating factorized logits: $\mathbf{S}_{C}=\mathbf{S}_{V}\boxplus\mathbf{S}_{O|V}+\mathbf{S}_{O}\boxplus\mathbf{S}_{V|O}$, where $\mathbf{S}_{O|V}$ and $\mathbf{S}_{V|O}$ represent the conditional logits and $\boxplus$ denotes broadcasted addition into the joint verb–object space.
+Then we employ the cross-entropy loss integrated with our batch-adaptive label space expansion (Section 4.1).
+Given $\mathbf{s}_{c}\in\mathbf{S}_{C}$, the final composition loss utilizing the dynamically constructed denominator $\mathbb{Y}_{\text{exp}}$ is formulated as: $L_{\text{comp}}=-\log(\exp(\hat{s}_{c})/\sum_{c=1}^{|\mathbb{Y}_{\text{exp}}|}\exp(s_{c}))$.
 Total loss.
-We train RCORE with the combined objective $L_{\text{total}}$ defined as $L_{\text{total}}=\alpha L_{\text{com}}+\beta L_{\text{comp}}+\gamma L_{\text{TORC}}+\delta L_{M}$ , where $\alpha$, $\beta$, $\gamma$, and $\delta$ control the contribution of each term.
+We train RCORE with the combined objective $L_{\text{total}}$ defined as $L_{\text{total}}=\alpha L_{\text{com}}+\beta L_{\text{comp}}+\gamma L_{\text{TORC}},+\delta L_{\text{CPR}},$ where $\alpha$, $\beta$, $\gamma$ and $\delta$ are hyperparameters.
 5 Experimental Results
-In this section, we conduct rigorous experiments to address the following research questions: (1) Do conventional ZS-CAR evaluation protocols faithfully reveal true model behavior?
-If not, what is a more appropriate protocol for exposing real generalization performance?
-(Section 5.1) (2) Does the proposed compositional gap (CG) metric reveal whether a model actually benefits from compositional understanding beyond independent verb/object predictions?
-(Section 5.2 and Section 5.3) (3) Does RCORE mitigate the two key failure modes—co-occurrence overfitting and object-driven verb shortcuts—identified in our diagnosis?
-(Section 5.2) (4) Do these improvements translate to consistent gains across datasets?
-(Section 5.3) (5) How does each proposed component contribute to robust compositional learning?
-(Section 5.4) We describe the experimental setup and then report results aligned with these questions.
-Table 2: Sth-com [16] results.
-We show the top-1 verb, object, and composition classification accuracy (%).
-All models use CLIP-B/16 [28] as the backbone.
-We report performance on both seen and unseen compositions, along with their harmonic mean (H.M.).
-Model Bias Calib. Verb Object Composition
-Seen Unseen H.M. Seen Unseen H.M. Open-world Closed-world
-Seen $(\Delta_{\mathrm{CG}}{})$ Unseen $(\Delta_{\mathrm{CG}}{})$ H.M. H.M. AUC
-AIM [37] $\times$ 54.99 39.19 45.76 67.19 53.43 59.53 40.24 (+3.29) 18.50 (-2.44) 25.35 - -
-C2C w/o conditioning $\times$ 62.53 48.67 54.74 66.90 52.66 58.93 44.95 (+3.12) 23.94 (-1.69) 25.35 - -
-C2C [16] $\times$ 63.60 54.36 58.62 67.72 56.10 61.36 46.31 (+3.24) 30.08 (-0.42) 36.47 - -
-RCORE $\times$ 65.93 57.29 61.31 67.82 55.56 61.08 48.56 (+3.85) 32.23 (+0.40) 38.75 - -
-AIM [37] 50.32 43.19 46.48 64.02 54.78 59.04 33.40 (+1.19) 24.60 (+0.94) 28.33 34.44 16.99
-C2C w/o conditioning 57.08 50.31 53.48 65.07 53.32 58.61 39.32 (+2.18) 26.70 (+1.89) 31.80 37.62 19.74
-C2C [16] 60.87 56.81 58.77 65.83 57.37 58.93 42.27 (+2.20) 34.22 (+1.30) 37.82 42.75 23.77
-Jung et al. [14] - - - - - - - - - 42.60 23.50
-RCORE 63.57 58.77 61.08 66.73 56.78 61.35 45.34 (+2.91) 35.26 (+1.89) 39.67 45.45 26.52
-C2C (Baseline)
-RCORE (Ours)
-Figure 6: RCORE mitigates object-driven shortcuts in verb learning.
-We visualize confusion matrices for six representative verbs to compare the ability of RCORE and C2C to distinguish opposite temporal semantics on unseen compositions of the Sth-com [16] test set.
-All values in the confusion matrices are normalized frequencies across the entire verb classes in the dataset.
+Building on our diagnosis, we evaluate whether RCORE reduces shortcut reliance and improves compositional generalization.
+We first describe the experimental setup (Section 5.1), then present diagnostic analyses (Section 5.2), followed by quantitative comparisons (Section 5.3), and finally ablation studies (Section 5.4).
 5.1 Experimental setup
-Limitations of conventional ZS-CAR evaluation.
-Existing ZS-CAR works [16, 13, 14] rely on the closed-world protocol, where inference is restricted to a small set of compositions that appear in the validation or test split.
-This protocol hides a model’s incorrect tendency to over-predict seen compositions and obscures co-occurrence-driven behaviors.
-A more serious methodological flaw is the use of test-set–tuned bias calibration.
-Selecting the bias term with access to test labels systematically inflates unseen accuracy and prevents meaningful comparison across methods.
+Limitations of conventional evaluation.
+Existing ZS-CAR works [li2024c2c, jiang2025dhd, jung2025crr] rely on the closed-world protocol, where inference is restricted to compositions that appear in the validation/test split.
+This protocol hides a model’s tendency to over-predict seen compositions and obscures co-occurrence-driven behaviors.
+More critically, test-set–tuned bias calibration selects a bias term using test labels, inflating unseen accuracy and undermining fair comparison.
 Our evaluation setup.
 To expose genuine generalization, we adopt an open-world, unbiased protocol by default: (i) inference spans the full space $\mathbb{Y}^{V}\times\mathbb{Y}^{O}$, and (ii) no test labels are used to tune biases.
-We also report: (a) open-world biased results using a realistic validation set-tuned bias, and (b) closed-world H.M.
+We also report closed-world H.M.
 and AUC for comparison with existing methods.
-All results in Section 3 also follow the default open-world, unbiased protocol.
+We show open-world biased results using a realistic validation set-tuned bias in Section 0.E.4.
 Datasets.
-We evaluate on two ZS-CAR benchmarks: Sth-com [16] and our curated EK100-com.
-Sth-com is derived from Something-Something V2 [12] and Something-Else [22], containing 79K videos with 161 verbs and 248 objects.
-Following prior work [39, 31], we also evaluate on the Temporal subset consisting of verbs requiring temporal reasoning.
-We additionally introduce the EPIC-KITCHENS-100-composition (EK100-com) benchmark, constructed by repurposing EPIC-KITCHENS-100 [7] following the Sth-com protocol.
+We evaluate on two ZS-CAR benchmarks: Sth-com [li2024c2c] and our curated EK100-com.
+Sth-com is derived from Something-Something V2 [goyal2017something] and Something-Else [mat2020sthelse], containing 79K videos with 161 verbs and 248 objects.
+Following prior work [yun2022time, laura2021temporal], we also evaluate on the Temporal subset consisting of verbs requiring temporal reasoning.
+We additionally introduce the EPIC-KITCHENS-100-composition (EK100-com) dataset, constructed by repurposing EK100 [damen2022ek100] following the Sth-com protocol.
 EK100-com includes 71K egocentric videos, 82 verbs, and 228 objects.
-Full construction details are provided in the Section A.
+Due to the long-tailed label distribution of EK100, it exhibits severe compositional sparsity.
+Its label coverage ratio is only $7.5\%$, which is considerably lower than the $12.8\%$ observed in Sth-com [li2024c2c].
+Full details are in Section 0.B.
+Figure 5: Analysis on the effects of RCORE on the Sth-com [li2024c2c] dataset.
+(a) RCORE suppresses the growth of FSP/FCP during training, unlike the baseline.
+(b) This reduces the seen–unseen gap and improves unseen composition accuracy.
+(c) On the temporal subset, RCORE shows a larger drop under temporal shuffling, indicating stronger temporal dependence over static cues.
+Best viewed with zoom and color.
 Evaluation metrics.
-We report top-1 composition, verb, and object accuracy on seen and unseen compositions, along with their harmonic mean (H.M.).
-We also measure the compositional gap $\Delta_{\mathrm{CG}}$ (1) which indicates whether a model benefits from compositional reasoning beyond independent verb/object predictions.
-(see Section 3) A higher $\Delta_{\mathrm{CG}}$ reflects stronger compositional synergy.
+We report top-1 composition, verb and object accuracies.
+Verb and object accuracies are conditioned on the seen/unseen composition: verb@seen-comp and object@seen-comp are computed on samples with $\mathbf{y}^{C}\in\mathbb{Y}_{\text{seen}}$, while verb@unseen-comp and object@unseen-comp are computed on samples with $\mathbf{y}^{C}\in\mathbb{Y}_{\text{unseen}}$.
+We report the corresponding harmonic mean (H.M.) between seen-comp and unseen-comp accuracies for composition, verb, and object.
+We also report the compositional gap $\Delta_{\mathrm{CG}}$ (2).
+For a detailed illustration of our evaluation setting, please refer Section 0.A.
 Baselines.
-We primarily compare RCORE against C2C [16], the current SOTA ZS-CAR method.
-To isolate the effect of conditional learning, we also evaluate two independent baselines.
-AIM [37] is the CLIP [28]-based spatio-temporal adapter model used as our backbone.
-C2C w/o conditioning augments AIM with all architectural components of C2C except the conditional modeling itself.
-We additionally include closed-world results of a recent ZS-CAR approach [14].
+We employ two encoder-based VLMs, CLIP-B/16 [radford2021clip] and InternVi-deo2-CLIP-B/14 [wang2024iv2], as backbones for ZS-CAR methods.
+We put the InternVid-eo2-1B[wang2024iv2] backbone results in the supplementary material.
+We primarily compare RCORE against C2C [li2024c2c], the current SOTA ZS-CAR method.
+To isolate the effect of conditional learning, we also evaluate the independent modeling baselines with CLIP [radford2021clip] and InternVideo2 [wang2024iv2] backbones which are denoted as ‘AIM’ and ‘LoRA’, named after their respective adapter tuning methods, in Section 5.3.
+AIM [yang2023aim] is the CLIP [radford2021clip]-based spatio-temporal adapter model.
+We additionally include closed-world results of a recent ZS-CAR approach [jung2025crr].
+5.2 Analysis
+We empirically show that RCORE reduces reliance on co-occurrence priors and mitigates object-driven shortcuts in verb learning, yielding more robust compositional behavior than the baseline [li2024c2c].
+Here, we present results with the CLIP [radford2021clip] backbone, and we put InternVideo2 [wang2024iv2] results in the supplementary.
+C2C (Baseline)      RCORE (Ours)
+Figure 6: RCORE mitigates object-driven shortcuts in verb learning.
+We visualize confusion matrices for six representative verbs on unseen compositions of the Sth-com [li2024c2c] test set.
+All values are normalized frequencies across the six verb classes.
+RCORE less relies on co-occurrence prior.
+In Figure 5 (a), we present the False Seen Prediction (FSP) and False Co-occurrence Prediction (FCP) curves of RCORE and the baseline [li2024c2c] on Sth-com.
+While the baseline’s FSP and FCP increase substantially during training (53% $\rightarrow$ 63% and 26% $\rightarrow$ 29%), those of RCORE remain consistently lower and even decrease (47% $\rightarrow$ 44% and 23% $\rightarrow$ 23%), indicating markedly less dependence on co-occurrence statistics.
+Improved unseen generalization.
+In Figure 5 (b), we present the learning curve of our RCORE and the baseline [li2024c2c] on Sth-com.
+RCORE significantly improves unseen composition accuracy upon the baseline (34% vs.
+30%), and reduces the seen–unseen accuracy gap (9 points vs.
+17 points).
+The learned verb representations are indeed effective.
+Inspired by prior work [yun2022time], we evaluate models on the Temporal subset of Sth-com [li2024c2c], as shown in Figure 5 (c).
+Specifically, to measure a model’s dependence on static cues, we compare its performance when using temporally shuffled verb features $\mathbf{f}^{V}_{\text{shuffled}}$ versus the original features $\mathbf{f}^{V}$.
+Across both seen and unseen compositions, RCORE exhibits a substantially larger performance gap between original and shuffled features.
+This demonstrates that its verb representations are effectively grounded in temporal dynamics.
+In Figure 6, we visualize verb confusion matrices on unseen compositions of the Sth-com test set.
+RCORE discriminates pairs of opposite verbs, while C2C [li2024c2c] frequently confuses them, e.g.
+misclassifying ‘unfolding’ as ‘folding’.
+These results collectively demonstrate that C2C fails to overcome the challenge of object-driven shortcuts in verb learning, whereas RCORE effectively mitigates it.
+Table 2: Sth-com [li2024c2c] results.
+We show the top-1 verb, object, and composition classification accuracy (%).
+We report performance on both seen and unseen compositions, along with their harmonic mean (H.M.).
+Backbone Method Verb Object Composition
+@Seen Comp @Unseen Comp H.M. @Seen Comp @Unseen Comp H.M. Seen ($\Delta_{\text{CG}}$) Unseen ($\Delta_{\text{CG}}$) H.M.
+CLIP AIM [yang2023aim] 54.99 39.19 45.76 67.19 53.43 59.53 40.24 (+3.29) 18.50 (-2.44) 25.35
+C2C [li2024c2c] 63.60 54.36 58.62 67.72 56.10 61.36 46.31 (+3.24) 30.08 (-0.42) 36.47
+RCORE 65.73 59.00 62.18 64.79 56.34 60.27 44.99 (+2.40) 33.90 (+0.66) 38.67
+InternVideo2 LoRA [hu2022lora] 48.83 36.50 41.78 64.79 58.61 61.55 32.83 (+1.19) 19.71 (-1.68) 24.63
+C2C [li2024c2c] 70.35 63.29 66.63 70.03 63.44 66.58 51.50 (+2.24) 39.53 (-0.63) 44.73
+RCORE 71.65 66.65 69.06 67.96 64.56 66.22 50.20 (+1.51) 43.98 (+0.95) 46.88
+5.3 Comparisons with State-of-the-Art
+Sth-com.
+As shown in Table 2, RCORE achieves clear gains on Sth-com [li2024c2c] under a rigorous open-world setting.
+Compared to the C2C [li2024c2c] baseline, RCORE improves the verb@unseen-comp by 4.6 points and the unseen composition accuracy by 3.8 points with the CLIP [radford2021clip] backbone, and by 3.4 points and 4.5 points with InternVideo2 (IV2) [wang2024iv2], respectively.
+These gains are consistent with stronger verb recognition observed in our analysis (Section 5.2).
+Notably, RCORE yields a positive compositional gap ($\Delta_{\text{CG}}$) on unseen compositions, whereas all baselines remain negative, indicating improved compositional behavior on unseen pairs.
+This improvement coincides with the highest composition H.M.
+under the open-world protocol.
+In Table 3, RCORE also records the highest H.M.
+and AUC under the closed-world, test-set-tuned protocol used in prior work [jung2025crr].
+Table 3: Closed-world results on Sth-com[li2024c2c].
+Method Closed-world
+H.M. AUC
+AIM [yang2023aim] 34.44 16.99
+C2C [li2024c2c] 42.75 23.77
+Jung et al. [jung2025crr] 42.60 23.50
+RCORE 43.21 25.27
+EK100-com.
+Despite the strong co-occurrence prior in the EK100-com training compositions (see Section 5.1), RCORE achieves higher unseen composition accuracy and the best H.M.
+As shown in Table 4, compared to C2C [li2024c2c], RCORE substantially improves unseen composition accuracy—by 6.9 points with the CLIP [radford2021clip] backbone and 7.7 points with the IV2 [wang2024iv2] backbone—despite a decrease in seen composition accuracy.
+This pattern is consistent with reduced reliance on seen co-occurrence statistics and improved recognition of unseen compositions.
+Moreover, RCORE yields a positive compositional gap ($\Delta_{\text{CG}}$) on unseen compositions (+1.32 with CLIP and +3.68 with IV2), which is consistent with improved compositional behavior under highly sparse supervision.
 Table 4: EK100-com results.
 We show the top-1 verb, object, and composition classification accuracy (%).
-All models use CLIP-B/16 [28] as the backbone.
 We report performance on both seen and unseen compositions, along with their harmonic mean (H.M.).
-Model Bias Calib. Verb Object Open-world Composition
-Seen Unseen H.M. Seen Unseen H.M. Seen $(\Delta_{\mathrm{CG}}{})$ Unseen $(\Delta_{\mathrm{CG}}{})$ H.M.
-AIM [37] $\times$ 60.04 39.78 47.85 56.04 46.33 50.72 38.60 (+4.95) 14.74 (-3.69) 21.34
-C2C [16] $\times$ 66.19 49.71 56.78 56.61 47.48 51.64 42.72 (+5.25) 22.38 (-1.22) 29.38
-RCORE $\times$ 66.07 52.07 58.24 57.76 47.11 51.89 43.86 (+5.70) 24.24 (-0.29) 31.23
-AIM [37] 52.42 49.54 50.94 50.70 51.05 50.87 28.00 (+1.42) 32.07 (+6.78) 29.89
-C2C [16] 62.21 57.74 59.89 51.66 54.16 52.88 36.05 (+3.91) 37.96 (+6.69) 36.98
-RCORE 62.47 59.77 61.09 53.42 54.79 54.10 36.53 (+3.16) 39.78 (+7.03) 38.09
-5.2 Analysis
-In this section, we validate that our RCORE framework effectively addresses the two key challenges in ZS-CAR—(i) reliance on co-occurrence statistics and (ii) object-driven shortcuts in verb learning—and achieves improved compositional representations compared to the baseline [16].
-Overcoming reliance on co-occurrence statistics.
-In Figure 5 (a), we present the False Co-occurrence Prediction (FCP) curves of RCORE and the baseline [16] on Sth-com.
-While the baseline’s FCP ratio increases substantially during training (11% $\rightarrow$ 17%), the FCP of RCORE decreases (9% $\rightarrow$ 8%), indicating that RCORE is markedly less dependent on co-occurrence statistics.
-Addressing object-driven shortcuts.
-In Figure 5 (b), we compare the cosine similarity between the original verb features $\mathbf{f}^{V}$ and the reversed verb features $\mathbf{f}^{V}_{\text{rev}}$ for RCORE and the baseline [16] over the course of training.
-While the baseline [16] maintains a high cosine similarity ($0.91$) due to its reliance on object-driven shortcuts, RCORE drives this value to a strongly negative (0.73-0.73), successfully distinguishing opposite temporal semantics such as ’opening’ vs.’closing’.
-In Figure 6, we further visualize this behavior through verb confusion matrices on the Sth-com [16] test set.
-RCORE cleanly separates pairs of opposite verbs, while C2C [16] frequently confuses them, e.g.
-misclassifying ‘unfolding’ as ‘folding’.
-These results collectively demonstrate that C2C fails to overcome the critical challenge of object-driven shortcuts in verb learning, whereas RCORE directly targets this failure mode and learns verb representations that generalize robustly to unseen compositions.
-The learned verb representations are indeed effective.
-We further assess the quality of the learned verb representations by evaluating temporal modeling ability on the Temporal subset of Sth-com [16], as shown in Figure 5 (c).
-To measure a model’s dependence on static cues, we compare its performance when using temporally shuffled verb features $\mathbf{f}^{V}_{\text{shuffled}}$ versus the original features $\mathbf{f}^{V}$.
-Across both seen and unseen compositions, RCORE exhibits a substantially larger performance gap between original and shuffled features.
-This demonstrates that its verb representations are effectively grounded in the temporal dynamics of the input.
-Improved unseen generalization.
-In Figure 5 (a), we present the learning curve of our RCORE and the baseline [16] on Sth-com.
-Addressing the two key challenges in ZS-CAR, RCORE significantly improves unseen composition accuracy upon the baseline.
-While the baseline significantly increases FCP ratio from 10.8% at 5 epochs to 16.8% at 25 epochs, we observe our RCORE reports FCP ratio values of 10.8% at 5 epochs and 8.7% at 25 epochs, demonstrating stable mitigation of the co-occurrence bias.
-Notably, we also see a reduction in the seen-unseen accuracy gap.
-Table 4: Ablation study on the Sth-com [16].
-In every experiment, we use C2C [16] as our baseline.
-We report performance on both seen and unseen compositions, along with their harmonic mean (H.M.).
-We report the performances using the unbiased open-world setting.
-The best H.M.
-numbers are highlighted.
-(a) Effects of RCORE components.
-VOCAMix TORC Margin Verb Object Composition Seen Unseen Seen Unseen Seen Unseen H.M.
-63.60 54.36 67.72 56.10 46.31 30.08 36.47 64.26 54.65 67.32 56.61 46.86 30.35 36.84 65.65 56.80 68.59 55.19 48.65 31.35 38.13 64.60 54.42 66.99 55.68 46.71 29.80 36.38 65.72 56.60 67.91 55.81 47.68 31.96 38.27 65.93 57.29 67.82 55.56 48.56 32.23 38.75
-(b) Effects of TORC components.
-$L_{\cos}$ $L_{\text{ent}}$ Verb Object Composition (OW) Seen Unseen Seen Unseen Seen Unseen H.M.
-63.60 54.36 67.72 56.10 46.31 30.08 36.47 64.54 55.01 67.17 55.68 47.06 29.98 36.62 64.84 54.82 67.57 56.23 47.46 30.84 37.39 65.65 56.80 68.59 55.19 48.65 31.35 38.13
-(c) Effects of VOCAMix.
-Method Verb Object Composition Seen Unseen Seen Unseen Seen Unseen H.M.
-Baseline 63.60 54.36 67.72 56.10 46.31 30.08 36.47 CutMix [38] 63.87 52.40 68.08 57.01 46.63 29.00 35.76 Mixup [41] 62.79 52.42 67.84 56.34 45.82 29.37 35.79 VOCAMix w/o F.E.
-64.26 54.35 67.61 56.19 46.82 30.07 36.62 VOCAMix 64.26 54.65 67.32 56.61 46.86 30.35 36.84
-(d) Effects of Margin loss.
-Method Verb Object Composition FCP Seen Unseen Seen Unseen Seen Unseen H.M.
-Baseline (CE) 63.60 54.36 67.72 56.10 46.31 30.08 36.47 15.30 L-softmax [19] 63.39 52.87 67.56 55.96 46.59 29.11 35.84 10.93 Margin (Ours) 64.60 54.42 66.99 55.68 46.71 29.80 36.38 8.81
-5.3 Comparisons with State-of-the-Art
-Sth-com results.
-As shown in Table 2, RCORE delivers consistent improvements across all metrics under both unbiased and biased evaluation on the Sth-com [16] dataset.
-Importantly, RCORE achieves a positive compositional gap $(\Delta_{\mathrm{CG}})$ on unseen compositions under the open-world unbiased setting, whereas all baselines exhibit negative gaps.
-This highlights the strong compositional generalization ability of RCORE.
-The superior verb representations learned by RCORE—validated in Section 5.2—translate directly into substantial gains on the ZS-CAR dataset.
-EK100-com results.
-In Table 3, RCORE consistently outperforms C2C [16] on both seen and unseen compositions under both unbiased and biased evaluation settings, despite the egocentric challenges.
-Notably, RCORE achieves a substantial improvement in unseen verb accuracy compared to C2C (52.07% vs.
-49.71%), underscoring its stronger temporal modeling ability and reduced object-driven shortcuts.
+Backbone Method Verb Object Composition
+@Seen Comp @Unseen Comp H.M. @Seen Comp @Unseen Comp H.M. Seen ($\Delta_{\text{CG}}$) Unseen ($\Delta_{\text{CG}}$) H.M.
+CLIP AIM [yang2023aim] 59.97 39.99 47.98 56.23 46.76 51.06 38.58 (+4.86) 14.98 (-3.71) 21.58
+C2C [li2024c2c] 65.52 49.57 56.44 56.54 46.99 51.32 42.70 (+5.65) 21.56 (-1.73) 28.65
+RCORE 66.19 54.31 59.66 54.35 49.88 52.02 39.70 (+3.73) 28.41 (+1.32) 33.12
+InternVideo2 LoRA [hu2022lora] 59.73 32.99 42.51 51.08 39.34 44.45 36.43 (+5.92) 5.63 (-7.35) 9.75
+C2C [li2024c2c] 68.95 57.01 62.41 60.54 52.31 56.12 47.30 (+5.55) 30.33 (+0.51) 36.96
+RCORE 69.00 60.98 64.74 56.59 56.40 56.49 42.32 (+3.27) 38.07 (+3.68) 40.08
 5.4 Ablation Studies
 Effects of RCORE components.
-In Table 4 (a), we investigate the effect of each component in our RCORE framework.
-When used individually, TORC loss is the most effective by providing explicit temporal modeling objectives.
-Notably, it significantly improves verb recognition performances across both seen and unseen compositions (63.6% vs.
-65.7% on seen compositions and 54.4% vs.
-56.8% on unseen compositions).
-Importantly, we can achieve the highest ZS-CAR performance by employing all three components together during training.
-This synergistic improvement is achieved by effectively tackling dataset sparsity with VOCAMix, reinforcing temporal order modeling with TORC, and alleviating co-occurrence bias with Margin loss.
+In Table 5 (a), using TORC alone yields the largest gain in verb generalization, improving verb@unseen-comp by 3.5 points, and gain in unseen composition accuracy by 3.8 points, compared to the baseline without CPR and TORC (C2C [li2024c2c]).
+Using CPR alone primarily improves unseen composition accuracy by 3.1 points, while reducing seen composition accuracy 4.2 points, indicating a seen–unseen trade-off.
+Combining CPR and TORC achieves the best overall results, improving unseen composition accuracy by +5.0 points and composition H.M.
+by +2.7 points, suggesting complementary benefits of mitigating co-occurrence bias and enforcing temporal sensitivity.
 Effects of TORC components.
 Table 5 (b) evaluates the roles of $L_{\cos}$, which enforces discrimination of opposite temporal semantics, and $L_{\text{ent}}$, which prevents reliance on static cues.
-We show the each component of TORC can encourage performance improvement.
-Each term individually improves performance, while using both together yields the largest gains in unseen verb accuracy.
-Effects of VOCAMix.
-In Table 4 (c), we validate the effectiveness of our proposed composition-aware VOCAMix by comparing it with widely adopted mixing augmentations.
-We argue that conventional image-based augmentations are insufficient for composition verb-object learning.
-Specifically, CutMix [38] allows for mixing backgrounds with corresponding labels and Mixup [41] blends verbs and objects concurrently, failing to preserve the temporal difference information that is essential for verb recognition.
-As a result, these methods yield a lower H.M.
-score than the baseline.
-In contrast, our VOCAMix, aiming to generate realistic unseen composition samples, shows modest performance gains over the baseline.
-Furthermore, we observe that mixing a static frame into the estimated foreground region shows additional performance boost (+0.2 points in the H.M.) over applying it to the entire background (denoted as ‘VOCAMix w/o F.E.’).
-Effects of Margin loss.
-In Table 4 (d), we validate the effectiveness of our proposed Margin loss.
-To mitigate the model’s reliance on dataset co-occurrence priors while maintaining baseline performance, we employ our method and validate its effectiveness by comparing it with the widely-used large-margin softmax loss.
-The large-margin softmax (L-softmax) [19] loss is implemented by multiplying the angle $\theta$ between the input feature vector and the target classifier by a margin $m$ (e.g., $\cos(m\theta)$), thereby forcing the model to learn stricter decision boundaries with narrower angular margins.
-We apply this L-softmax loss into composition classification loss, consistent with the implementation schedule of our Margin loss.
-Compared to the L-softmax [19] approach, our proposed Margin loss successfully reduces FCP ratio while preserving performance.
-This demonstrates the effectiveness of our strategy, which selectively targets frequent-but-incorrect logits not the entire logits.
+Each term individually improves performance, while using both terms yields the largest gains in the verb@unseen-comp (+3.5 points) and unseen composition accuracy (+3.8 points).
+Table 5: Ablation study on the Sth-com [li2024c2c] validation set.
+We use C2C [li2024c2c] with the CLIP [radford2021clip] backbone as our baseline, using the unbiased open-world setting.
+We report performance on both seen and unseen compositions, along with their harmonic mean (H.M.).
+V@S and V@U denotes verb@seen-comp and verb@unseen-comp, while O@S and O@U denotes object@seen-comp and object@unseen-comp accuracies.
+(a) Effects of RCORE components.
+CPR TORC Verb Object Composition
+V@S V@U O@S O@U Seen Unseen H.M.
+65.26 54.02 67.21 56.81 47.31 30.14 36.83
+63.95 54.05 63.97 59.52 43.09 33.22 37.51
+64.93 57.49 66.20 57.23 46.28 33.92 39.15
+67.06 58.13 63.67 58.66 45.00 35.16 39.48
+(b) Effects of TORC components.
+$L_{\cos}$ $L_{\text{ent}}$ Verb Object Composition (OW) V@S V@U O@S O@U Seen Unseen H.M.
+65.26 54.02 67.21 56.81 47.31 30.14 36.83 64.15 55.96 67.71 55.97 47.13 31.11 37.48 64.84 54.82 67.57 56.23 47.46 30.84 37.39 64.93 57.49 66.20 57.23 46.28 33.92 39.15
+(c) Effects of label space in CPR.
+Label Space Verb Object Composition V@S V@U O@S O@U Seen Unseen H.M.
+Closed-world ($\mathbb{Y}_{\text{seen}}$) 64.61 54.26 67.21 57.46 46.78 30.80 37.14 Full Open-world ($\mathbb{Y}^{V}\times\mathbb{Y}^{O}$) 67.36 45.01 68.49 52.29 50.72 17.34 25.85 Batch-adaptive ($\mathbb{Y}_{\text{exp}}$) 64.17 54.55 63.60 58.70 42.74 33.05 37.28
+(d) Effects of penalty set in CPR.
+Penalty Set $\mathbb{Y}_{\text{pen}}$ Choice Verb Object Composition V@S V@U O@S O@U Seen Unseen H.M.
+None (w/o $L_{\text{CPR}}$) 64.17 54.55 63.60 58.70 42.74 33.05 37.28 All Hard Negatives ($\mathbb{Y}_{\text{hard}}$) 61.95 57.22 47.99 49.05 31.85 30.82 31.33 Frequent Hard Negatives ($\mathbb{Y}_{\text{hard}}\cap\mathbb{Y}_{\text{freq}}$) 65.36 55.86 63.20 57.62 43.60 33.16 37.67
+Effects of label space in CPR.
+In Table 5 (c), we validate the batch-adaptive label space expansion in CPR.
+Compared to closed-world training over $\mathbb{Y}_{\text{seen}}$, our batch-adaptive label space improves unseen composition accuracy (+2.3 points) and H.M.
+by injecting supervision for synthesized compositions while avoiding unstable optimization.
+In contrast, naively optimizing cross-entropy over the full $\mathbb{Y}^{V}\times\mathbb{Y}^{O}$ space severely hurts unseen composition accuracy by treating most unseen compositions as negatives throughout training.
+Effects of penalty set in CPR.
+In Table 5 (d), we study the effect of the penalty set $\mathbb{Y}_{\text{pen}}$.
+Compared to disabling $L_{\text{CPR}}$, penalizing frequent hard negatives ($\mathbb{Y}_{\text{hard}}\cap\mathbb{Y}_{\text{freq}}$) yields the best trade-off, improving unseen composition accuracy and achieving the highest H.M.
+of 37.7%.
+In contrast, penalizing all hard negatives degrades object prediction and hurts overall composition performance, suggesting that restricting penalties to frequent confounders is critical.
 6 Conclusions
-We revisit zero-shot compositional action recognition and show that existing methods fail mainly due to object-driven shortcuts, stemming from compositional sparsity and the asymmetric difficulty of learning verbs versus objects.
-We also find that the widely used closed-world, test-tuned evaluation protocol conceals this failure, overstating model generalization.
-We introduce RCORE, a simple framework that targets these issues directly: VOCAMix enlarges the effective compositional support, and TORC loss enforces temporal order sensitivity to reduce reliance on static object cues.
-Across two benchmarks and under a realistic open-world, unbiased evaluation setting, RCORE consistently improves unseen composition recognition performances, indicating genuine compositional reasoning.
-Our results highlight that both proper evaluation and shortcut-aware training are essential for progress in compositional video understanding.
+In this work, we revisit ZS-CAR and identify object-driven shortcuts as a key failure mode, driven by compositional sparsity and asymmetric verb–object learning difficulty.
+Using our diagnostic metrics, we show that existing models overfit to co-occurrence priors and consequently generalize poorly to unseen compositions.
+We propose RCORE with CPR (synthesized supervision with frequent hard-negative regularization) and TORC (temporal-order sensitivity), and demonstrate consistent gains on two benchmarks under open-world evaluation without test-tuned calibration.
 Acknowledgments
@@
 Appendix
-In this appendix, we provide comprehensive implementation/dataset/method details and quantitative/qualitative results to complement the main paper. We organize the appendix as follows:
+In this appendix, we provide comprehensive evaluation-protocol/ dataset/implementation details and additional results to complement the main paper. We organize the appendix as follows:
   1. 1.
-    Details on EK100-com dataset (Section A).
+    Overview of our evaluation setting (Section 0.A).
   2. 2.
-    Complete implementation details (Section B).
+    Details on EK100-com dataset (Section 0.B).
   3. 3.
-    Additional evidence of object-driven shortcuts (Section C).
+    Complete implementation details (Section 0.C).
   4. 4.
-    Additional results (Section D).
-Appendix A Details on EK100-com dataset
+    Additional evidence of object-driven shortcuts (Section 0.D).
+  5. 5.
+    Additional results (Section 0.E).
+Appendix 0.A Overview of our evaluation setting
+In Figure 7 and Figure 8, we briefly describe our evaluation setting.
+In Zero-Shot Compositional Action Recognition (ZS-CAR), the compositional label space spans the Cartesian product of the verb and object primitives ($|\mathbb{Y}^{V}|\times|\mathbb{Y}^{O}|$).
+At test time, samples are categorized into seen and unseen compositions depending on whether they were observed during training, as shown in Figure 7 (a).
+Under the closed-world evaluation setting, the evaluation is restricted to compositions present within the dataset (i.e., the union of the training and test sets).
+Consequently, the model computes compositional logits exclusively for these predefined closed-world pairs, as shown in Figure 7 (b).
+However, this closed-world setting suffers from the impractical assumption of requiring prior knowledge of the test set’s ground-truth labels.
+Therefore, we adopt the open-world evaluation setting as our primary protocol.
+By evaluating across the entire compositional space and computing logits for all possible combinations, as shown in Figure 7 (b), the open-world setting enables the evaluation of various possible-but-unseen compositions encountered in real-world applications, where the test distribution is unknown.
+(a) Closed-world and open-world evaluation settings
+(b) An example of closed/open-world evaluation
+Figure 7: Closed-world and open-world evaluation protocols in ZS-CAR.
+(a) (Left) Primitives are learned from a subset of training compositions (Train Pairs).
+(Middle) Closed-world evaluation restricts inference to dataset-present pairs (Train Pairs and Test Seen/Unseen Pairs), explicitly ignoring others as Masked Pairs.
+(Right) Open-world evaluation spans the entire compositional space, requiring the model to navigate all Possible Unseen Pairs without artificial masking.
+(b) In the closed-world setting, the Possible Unseen Pair—not appearing in the dataset—(Close, Window) becomes a Masked Pair, and its logit is ignored.
+Consequently, the next valid class (Open, Window) is selected as the top-1 prediction.
+This closed-world setting constrains the evaluation of numerous possible-but-unseen compositions.
+However, in the open-world setting, the model treats all pairs as valid, thus predicting (Close, Window) as the top-1 prediction.
+As shown in Figure 8, we measure verb, object, and composition accuracies across seen and unseen splits to evaluate not only the model’s compositional recognition capability, but also how primitive recognition degrades on unseen compositions.
+Specifically, based on the model’s composition predictions, we verify whether the composition, verb, or object is predicted correctly, yielding three sets of metrics: (1) seen/unseen composition accuracy, (2) verb@seen/unseen-comp, and (3) object@seen/unseen-comp.
+Figure 8: Evaluation metrics used in this work.
+We evaluate accuracies for verbs, objects, and compositions on seen and unseen composition splits separately.
+Based on the model’s composition predictions, we count the correct matches at each level to report (1) seen/unseen composition accuracy, (2) verb@seen/unseen-comp, and (3) object@seen/unseen-comp.
+For instance, predicting an unseen composition (Lift up, Spoon) as (Lift up, Pen) updates only the correct count for verb@unseen-comp (orange arrow).
+Appendix 0.B Details on EK100-com dataset
 In this section, we provide details about our curated ZS-CAR benchmark, EPIC-KITCHENS-100-composition (EK100-com).
-We construct EK100-com by repurposing EPIC-KITCHENS-100 (EK100) [7] following the same protocol of constructing Sth-com [16].
-In particular, we use the original training (67217 samples) and validation (9668 samples) split of EK100 [7] and split them as follows: (1) we filter the data to include only compositions with more than five samples each.
+We construct EK100-com by repurposing EPIC-KITCHENS-100 (EK100) [damen2022ek100] following the same protocol of constructing Sth-com [li2024c2c].
+In particular, we use the original training (67,217 samples) and validation (9,668 samples) split of EK100 [damen2022ek100] and split them as follows: (1) we filter the data to include only compositions with more than five samples each.
 (2) we ensure that all compositions present in the validation set also exist in the training set.
@@
 Beyond expanding the pool of publicly available ZS-CAR benchmarks, our repurposed EK100-com exhibits two key properties that make it a compelling ZS-CAR benchmark.
-First, unlike Something-Something V2 [12], where videos may involve multiple objects associated with a verb, videos in EK100 [7] depict actions consisting of a single verb and a single object.
+First, unlike Something-Something V2 [goyal2017something], where videos may involve multiple objects associated with a verb, videos in EK100 [damen2022ek100] depict actions consisting of a single verb and a single object.
 Therefore, the training data of EK100-com provides explicit object supervision for the models.
-Second, since EK100 [7] has long-tailed distribution, the label coverage ratio of EK100-com is 7.5%, which is much lower than that of Sth-com [16] (12.8%).
+Second, since EK100 [damen2022ek100] has long-tailed distribution, the label coverage ratio—defined as the number of unique verb–object pairs present in the dataset relative to all possible combinations—–of EK100-com is 7.5%, which is much lower than that of Sth-com [li2024c2c] (12.8%).
 This makes EK100-com a challenging benchmark, suitable for evaluating whether a model can overcome the risk of shortcut learning induced by data sparsity.
-(a) Top-30 frequent compositions
-(b) Bottom-30 frequent compositions
-Figure 7: Top/Bottom-30 frequent compositions in the EK100-com training set.
 Table 6: Details of the our introduced EK100-com benchmark. The label coverage ratio of EK100-com is $7.5\%=1320/(81\times 216)$.
@@
 Total 81 216 1320 71238
-(a) Seen composition examples
-(b) Unseen composition examples
+Top-30 frequent compositions
+Bottom-30 frequent compositions
+Figure 9: Top/Bottom-30 frequent compositions in the EK100-com training set.
+Seen composition examples
+Unseen composition examples
 Figure 10: Example of seen/unseen composition samples in the EK100-com dataset.
-Appendix B Complete Implementation Details
-B.1 Implementation details on our diagnosis
-In this section, we provide comprehensive details of experimental setup and implementation details of our diagnosis for Zero-Shot Compositional Action Recognition (ZS-CAR) in Section 3 in the main paper.
-All notations used in this supplementary document are consistent with those defined in the main paper.
-As mentioned in the main paper, we adopt the state-of-the-arts model, C2C [16] which uses a conditional learning approach, as the existing ZS-CAR baseline for diagnosis.
+Appendix 0.C Complete Implementation Details
+0.C.1 Implementation details on our diagnosis
+In this section, we provide comprehensive details of experimental setup and implementation details of our diagnosis for ZS-CAR in Section 3.
+All notations used in this appendix are consistent with those defined in the main paper.
+As mentioned in Section 3, we adopt the state-of-the-art model, C2C [li2024c2c], as the existing ZS-CAR baseline for diagnosis.
 Details about our diagnosis in Figure 2 (a).
 In Figure 2 (a) in the main paper, we empirically demonstrate the asymmetric learning difficulty between verbs and objects.
-We train a randomly initialized ViT [10] with joint spatio-temporal attention on our selected subset of the Sth-com [16] dataset.
-We aim to select a subset that maximizes the number of samples while maintaining a sufficient number of distinct labels.
+We train a randomly initialized ViT [dosovitskiy202vit] with joint Spatio-Temporal (ST) attention on our selected subset of the Sth-com [li2024c2c] dataset.
+We aim to select a dense subset that eliminates sparsity by ensuring samples exist for every pair, maximizing the number of samples while maintaining a sufficient number of distinct labels.
 As a result, the selected subset is composed of 10 verb classes: ‘Moving sth and sth away from each other’, ‘Moving sth and sth closer to each other’, ‘Moving sth away from sth’, ‘Moving sth closer to sth’, ‘Moving sth down’, ‘Moving sth up’, ‘Pushing sth from left to right’, ‘Pushing sth from right to left’, ‘Pushing sth so that it falls off the table’, and ‘Pushing sth so that it slightly moves’; and 10 object classes: ‘Ball’, ‘Battery’, ‘Book’, ‘Bottle’, ‘Coin’, ‘Glass’, ‘Knife’, ‘Paper’, ‘Remote’, and ‘Spoon’.
-We construct the training set consisting of 2,555 samples from the original Sth-Com [16] training and validation samples, while the validation set comprises 260 samples from the original Sth-Com [16] test set.
+We construct the training set consisting of 2,555 samples from the original Sth-Com [li2024c2c] training and validation samples, while the validation set comprises 260 samples from the original Sth-Com [li2024c2c] test set.
 For training, we employ a multi-task learning approach using separate linear classifiers for verbs and objects, taking the ViT’s CLS token as input.
-In the main paper, we show that objects are easier to learn than verbs by monitoring the learning curves of a randomly initialized ViT.
-Furthermore, we substantiate in Section C demonstrating that this asymmetric learning difficulty leads the model to exploit object-driven shortcuts in compositional video understanding.
+Figure 11: Object-driven shortcuts with the random initialized model.
+Even with a randomly initialized model, a ViT-B [dosovitskiy202vit] with joint ST attention, object-driven shortcuts emerge: the model achieves high object accuracy, but verb accuracy on bias-conflicting unseen compositions drops below chance, indicating shortcut-driven verb failures.
 Details about our diagnosis in Figure 2 (b).
-To precisely control the sample counts for each verb-object pair, we select a $4\times 4$ toy set from the Sth-Com [16] dataset.
+To precisely control the sample counts for each verb-object pair, we select a $4\times 4$ toy set from the Sth-Com [li2024c2c] dataset.
 This set contains four verb labels (‘Moving sth down’, ‘Moving sth up’, ‘Pushing sth from left to right’, and ‘Pushing sth from right to left’) and four object labels (Book, Remote, Glass, Bottle).
-The training set consists of the four diagonal pairs (refer to Figure 2 (b)) with 40 samples per pair.
+As shown in Figure 2 and Figure 11, the training set consists of the four diagonal pairs with 40 samples per pair.
 The test set comprises two distinct splits: seen compositions (25 samples per pair) and unseen compositions (30 samples per pair).
 We use 8 frames for input videos.
+In Figure 11, we demonstrate that the object-driven verb shortcut learning observed with CLIP [radford2021clip] in the main paper emerges identically even in a randomly initialized model.
+This indicates that this phenomenon is not a side effect of pre-trained knowledge, but rather a structural issue inherent to ZS-CAR.
 Details about our diagnosis in Figure 3.
-In Figure 3 in the main paper, we present the learning curves of the baseline model [16] trained on the Sth-com [16].
+In Figure 3, we present the learning curves of the baseline model [li2024c2c] trained on the Sth-com [li2024c2c].
 We further present diagnostic results for both the baseline and RCORE on EK100-com in Figure 12.
-We define two diagnostic metrics, False Seen Prediction (FSP) ratio and False Co-occurrence Prediction (FCP) ratio.
-FSP measures the proportion of unseen validation samples incorrectly classified as seen compositions:
-$\textit{False Seen Prediction}=\frac{1}{|\mathbb{D}_{\text{val}}^{\text{unseen}}|}{\sum_{i=1}^{|\mathbb{D}_{\text{val}}^{\text{unseen}}|}\mathbb{I}(\hat{\mathbf{y}}_{i}\in\mathbb{Y}_{\text{seen}})},$ (7)
+We define two misclassification ratios, False Seen Prediction (FSP) and False Co-occurrence Prediction (FCP).
+FSP measures the proportion of unseen composition validation samples incorrectly classified as seen compositions:
+$\vskip-30.00005pt\textit{False Seen Prediction}=\frac{1}{|\mathbb{D}_{\text{val}}^{\text{unseen}}|}{\sum_{i=1}^{|\mathbb{D}_{\text{val}}^{\text{unseen}}|}\mathbb{I}(\hat{\mathbf{y}}^{C}_{i}\in\mathbb{Y}_{\text{seen}})},\vskip-5.0pt$ (9)
 where $\mathbb{I}(\cdot)$ is the indicator function.
-FCP further quantifies the fraction of these false-seen predictions that correspond to frequently co-occurring verb–object pairs in the training data:
-$\textit{False Co-oc.
-Prediction}=\frac{\sum_{i=1}^{|\mathbb{D}_{\text{val}}^{\text{unseen}}|}\mathbb{I}(\hat{\mathbf{y}}_{i}\in\mathbb{Y}_{\text{freq}})}{\sum_{i=1}^{|\mathbb{D}_{\text{val}}^{\text{unseen}}|}\mathbb{I}(\hat{\mathbf{y}}_{i}\in\mathbb{Y}_{\text{seen}})},$ (8)
-where $\mathbb{Y}_{\text{freq}}$ denotes the set of frequently observed verb–object pairs in the training set.
-$\mathbb{Y}_{\text{freq}}$ is defined with the frequency distribution of the training set $P_{\mathbb{D}_{train}}$ as follows:
-$\mathbb{Y}_{\text{freq}}$ $=\mathbb{Y}_{\text{freq}}^{O\mid V}\cap\mathbb{Y}_{\text{freq}}^{V\mid O},$ (9)
-where $\mathbb{Y}_{\text{freq}}^{O\mid V}$ is defined as the set of pairs where the conditional probability of an object given a verb in the training data is greater than $\mu+\sigma$.
-Here, $\mu$ and $\sigma$ represent the mean and standard deviation of the conditional probabilities calculated on the training set.
-The $\mathbb{Y}_{\text{freq}}^{V\mid O}$ is defined similarly for the conditional probability of a verb given an object.
-$(\mu_{O|V}+\sigma_{O|V},\mu_{V|O}+\sigma_{V|O})$ are calculated as $(0.13,0.19)$ for the Sth-com [16] dataset and $(0.23,0.40)$ for EK100-com.
-Figure 9: Learning curve of the baseline and RCORE on the EK100-com dataset.
-RCORE suppresses the increase of the FCP ratio during training, effectively narrowing the performance gap between seen and unseen composition validation accuracies.
-B.2 Implementation details on RCORE
-Experimental Setup.
-We implement our method using PyTorch and conduct all experiments on 8 NVIDIA Tesla V100 GPUs.
-For video preprocessing, we uniformly sample frames from each video to generate clips of size $3\times 16\times 224\times 224$ (channels $\times$ frames $\times$ height $\times$ width).
-Model Architecture.
-We utilize CLIP-B/16 [28] as our visual/text backbone.
-Following prior work [16], we adopt AIM [37], a parameter-efficient spatio-temporal adapter for CLIP [28], as the video encoder.
-We also construct the verb encoder with two temporal convolution layers with ReLU activations, while the object encoder employs temporal average pooling followed by a two-layer MLP with ReLU activations.
-We use CoOp-style [43] learnable text prompts for all models.
-Training Strategy.
-We train all models, including baselines and RCORE, for 30 epochs with a total batch size of 128.
-The base learning rate is set to $5\times 10^{-4}$ for Sth-com [16] and $1\times 10^{-4}$ for EK100-com, initialized with a 3-epoch warmup.
-Following prior work [16], we adopt closed-world training setting.
-We empirically find that setting the softmax temperature to 0.07, same as the implementation of CLIP [28], results in better ZS-CAR performance than the value of 0.01, used in prior work [16].
-Therefore, we adopt this temperature value for both C2C [16] and RCORE, except for the case of AIM [37] with independent modeling.
-For VOCAMix, we employ the unsupervised foreground estimation algorithm FAME [9] to identify high-motion regions.
-FAME [9] isolates the moving foreground from background regions based on frame differences and color statistics.
-To generate the mixed samples, we first sample the mixing coefficient from a $\text{Beta}(2.0,2.0)$ distribution, following prior work [38, 41].
-We then scale it by 0.2 for Sth-com [16] and 0.5 for EK100-com to obtain the final blending coefficient $\lambda$.
-We summarize the hyperparameters used in this work in Table 6.
-Loss configuration.
-For the total loss, we set the weights for the verb and object components to $\alpha=0.2$ and for the composition loss to $\beta=1.0$, following prior work [16].
-As demonstrated in Table 9 (g) and (h), we empirically find it beneficial to linearly warm up the weights for the auxiliary losses: $\gamma$ is increased from 0.0 to 1.0 over 5 epoch to 10 epoch, and $\delta$ from 0.0 to 0.8 over epochs 15 epoch to 20 epoch.
-Inference.
-During inference, we adopt a single-view, single-crop protocol. For bias calibration, we select the bias term that maximizes the harmonic mean of the accuracies for seen and unseen compositions.
-Table 6: Hyperparameters used for training RCORE on each dataset.
-Config Sth-com EK100-com
-Optimizer AdamW [20]
-Base learning rate (visual) 5e-4 1e-4
-Base learning rate (text) 1e-4 1e-5
+FCP further quantifies the fraction of these false-seen predictions that correspond to frequently co-occurring verb–object pairs $\mathbb{Y}_{\text{freq}}$ in the training data:
+$\vskip-10.00002pt\textit{False Co-oc.
+Prediction}=\frac{1}{|\mathbb{D}_{\text{val}}^{\text{unseen}}|}{\sum_{i=1}^{|\mathbb{D}_{\text{val}}^{\text{unseen}}|}\mathbb{I}(\hat{\mathbf{y}}^{C}_{i}\in\mathbb{Y}_{\text{freq}})}.\vskip-10.00002pt$ (10)
+We further break down these misclassification ratios into the three cases: (i) Verb-collapse, where the predicted object is correct but the verb is incorrect; (ii) Object-collapse, where the predicted verb is correct but the object is incorrect; and (iii) Dual-collapse, where both verb and object are incorrect.
+We define them in FSP as follows:
+Verb-collapse $=\frac{\sum_{i=1}^{|\mathbb{D}_{\text{val}}^{\text{unseen}}|}\mathbb{I}(\hat{\mathbf{y}}^{C}_{i}\in\mathbb{Y}_{\text{freq}})\cdot\mathbb{I}(\hat{\mathbf{y}}^{V}_{i}\neq\mathbf{y}^{V}_{i}\land\hat{\mathbf{y}}^{O}_{i}=\mathbf{y}^{O}_{i})}{\sum_{i=1}^{|\mathbb{D}_{\text{val}}^{\text{unseen}}|}\mathbb{I}(\hat{\mathbf{y}}^{C}_{i}\in\mathbb{Y}_{\text{freq}})}$ (11)
+Object-collapse $=\frac{\sum_{i=1}^{|\mathbb{D}_{\text{val}}^{\text{unseen}}|}\mathbb{I}(\hat{\mathbf{y}}^{C}_{i}\in\mathbb{Y}_{\text{freq}})\cdot\mathbb{I}(\hat{\mathbf{y}}^{V}_{i}=\mathbf{y}^{V}_{i}\land\hat{\mathbf{y}}^{O}_{i}\neq\mathbf{y}^{O}_{i})}{\sum_{i=1}^{|\mathbb{D}_{\text{val}}^{\text{unseen}}|}\mathbb{I}(\hat{\mathbf{y}}^{C}_{i}\in\mathbb{Y}_{\text{freq}})}$ (12)
+Dual-collapse $=\frac{\sum_{i=1}^{|\mathbb{D}_{\text{val}}^{\text{unseen}}|}\mathbb{I}(\hat{\mathbf{y}}^{C}_{i}\in\mathbb{Y}_{\text{freq}})\cdot\mathbb{I}(\hat{\mathbf{y}}^{V}_{i}\neq\mathbf{y}^{V}_{i}\land\hat{\mathbf{y}}^{O}_{i}\neq\mathbf{y}^{O}_{i})}{\sum_{i=1}^{|\mathbb{D}_{\text{val}}^{\text{unseen}}|}\mathbb{I}(\hat{\mathbf{y}}^{C}_{i}\in\mathbb{Y}_{\text{freq}})},$ (13)
+where $\mathbf{y}^{C}_{i}=(\mathbf{y}^{V}_{i},\mathbf{y}^{O}_{i})$. We apply the identical formulation for FCP as well.
+(a) FSP/FCP
+(b) Composition accuracy
+Figure 12: Learning curve of the baseline and RCORE on EK100-com.
+RCORE (a) suppresses the increase of FCP during training, (b) effectively narrowing the performance gap between seen and unseen composition validation accuracies, compared to the baseline (C2C [li2024c2c]).
+0.C.2 Implementation details on RCORE
+Table 7: Hyperparameters used for training RCORE on each dataset. We denote InternVideo2 [wang2024iv2] as ‘IV2’.
+Config Sth-com [li2024c2c] EK100-com
+CLIP [radford2021clip] IV2 [wang2024iv2] CLIP [radford2021clip] IV2 [wang2024iv2]
+Optimizer AdamW [adamw]
+Base learning rate (visual) 5e-4 5e-5 1e-4 5e-5
+Base learning rate (text) 1e-4 5e-5 1e-5 5e-5
 Weight decay (visual) 1e-4
@@
 Warmup epochs 3
-$\alpha,\beta,\gamma,\delta$ 0.2, 1.0, 1.0, 0.8
-$p_{aug}$ 0.5 0.3
+$\alpha,\beta,\gamma,\delta$ 0.2, 1.0, 1.0, 0.1
+$p_{\text{CPR}}$ 0.95 0.8 0.95 0.8
 (start, end) epoch for scaling $\gamma$ (5, 10)
-(start, end) epoch for scaling $\delta$ (15, 20)
-Scale factor for $\lambda$ in VOCAMix 0.2 0.5
-K for top-k selection in Margin loss 10
+(start, end) epoch for scaling $\delta$ (15, 15) (30, 35) (35, 40) (30, 35)
+Scale factor for $\lambda$ in CPR 0.5
 Softmax temperature 0.07
-Training epochs 30
+Training epochs 40 50 40 40
+Experimental Setup.
+We implement our method using PyTorch and conduct all experiments on 8 NVIDIA Tesla V100 GPUs.
+For video preprocessing, we uniformly sample frames from each video to generate clips of size $3\times 16\times 224\times 224$ (channels $\times$ frames $\times$ height $\times$ width) for the CLIP [radford2021clip] backbone and $3\times 8\times 224\times 224$ for the InternVideo2 [wang2024iv2] backbone.
+Model Architecture.
+We utilize CLIP-B/16 [radford2021clip] as our visual/text backbone.
+Following prior work [li2024c2c], we adopt AIM [yang2023aim], a parameter-efficient spatio-temporal adapter for CLIP [radford2021clip], as the video encoder.
+We also employ InternVideo2-CLIP-B/14 [wang2024iv2] as our modern video-pretrained VLM backbone and utilize LoRA [hu2022lora] for parameter-efficient fine-tuning.
+We use CoOp-style [zhou2022coop] learnable text prompts for all models.
+We also construct the verb encoder with two temporal convolution layers with ReLU activations, while the object encoder employs temporal average pooling followed by a two-layer MLP with ReLU activations.
+To extract $T$ frame-level features with the InternVideo2 [wang2024iv2] backbone, we modify its attention pooling mechanism.
+Originally, the model concatenates a CLS token with video tokens in $\mathbb{R}^{NT\times D}$, where $N$ is the number of spatial patches and $D$ is the feature dimension, and applies attention pooling to produce a single video-level representation in $\mathbb{R}^{1\times D}$.
+In our approach, we adapt this to operate at the frame level.
+Specifically, for each frame, we concatenate the CLS token with the corresponding spatial tokens in $\mathbb{R}^{N\times D}$.
+This creates $T$ separate feature sets, each in $\mathbb{R}^{(N+1)\times D}$.
+Finally, we apply attention pooling independently to each of these sets, resulting in the frame-level video features in $\mathbb{R}^{T\times D}$.
+Training Strategy.
+We train all models, including the baselines and RCORE, following the hyperparameters detailed in Table 7, using a total batch size of 128.
+For CPR, we employ the unsupervised foreground estimation algorithm FAME [ding2022fame] to identify high-motion regions.
+FAME [ding2022fame] isolates the moving foreground from background regions based on frame differences and color statistics.
+To generate the mixed samples, we first sample the mixing coefficient from a $\text{Beta}(2.0,2.0)$ distribution, following prior work [yun2019cutmix, zhang2018mixup].
+We then scale it by 0.5 to obtain the final blending coefficient $\lambda$.
+We apply this CPR’s augmentation to the samples within a batch with a probability of $p_{\text{aug}}$.
+For the frequent verb-object pairs to apply the $L_{\text{CPR}}$ loss term, we select pairs that possess a strong co-occurrence prior from both verb and object perspectives as the frequent pairs $\mathbb{Y}_{\text{freq}}$, which is defined as $\mathbb{Y}_{\text{freq}}=\mathbb{Y}_{\text{freq}}^{O\mid V}\cap\mathbb{Y}_{\text{freq}}^{V\mid O}$.
+$\mathbb{Y}_{\text{freq}}^{O\mid V}$ is the set of pairs where the conditional probability of an object given a verb in the training data is greater than $\mu+\sigma$.
+Here, $\mu$ and $\sigma$ represent the mean and standard deviation of the conditional probabilities calculated on the training set.
+The $\mathbb{Y}_{\text{freq}}^{V\mid O}$ is defined similarly for the conditional probability of a verb given an object.
+$(\mu_{O|V}+\sigma_{O|V},\mu_{V|O}+\sigma_{V|O})$ are calculated as $(0.13,0.19)$ for the Sth-com [li2024c2c] dataset and $(0.23,0.40)$ for EK100-com.
+Loss configuration.
+For the total loss, we set the weights for the verb and object components to $\alpha=0.2$ and for the composition loss to $\beta=1.0$, following prior work [li2024c2c].
+As demonstrated in Table 11 (d), we empirically find it beneficial to linearly warm up the weights for the TORC loss: $\gamma$ is increased from 0.0 to 1.0 over 5 epoch to 10 epoch.
+We use $\delta$, the loss scale of $L_{\text{CPR}}$, differently depending on the dataset and backbone, as detailed in Table 7.
+Inference.
+During inference, we adopt a single-view, single-crop protocol.
+To select the best model during training and for bias calibration, we use the harmonic mean of the accuracies for seen and unseen compositions as the model selection criterion.
 Table 8: Top-10 failure cases of the existing ZS-CAR model on unseen compositions of the Sth-com test set.
-We employ C2C [16] trained on Sth-com [16] as baseline for analysis.
+We employ C2C [li2024c2c] trained on Sth-com [li2024c2c] as baseline for analysis.
 We use ‘sth’ instead of ‘something’ to shorten.
@@
 Table 9: Top-10 failure cases of the existing ZS-CAR model on unseen compositions of the EK100-com test set.
-We employ C2C [16] trained on EK100-com as baseline for analysis.
+We employ C2C [li2024c2c] trained on EK100-com as baseline for analysis.
 We report the results with unbiased open-world inference scores.
@@
 The top-10 cases listed below account for $\sim$ 39% of the total mispredictions on unseen compositions.
-Rank Ground Truth (Verb, Object) Error Top-2 Misprediction (seen/unseen) Prediction Most Frequent Co-occurring Component During Training
+Rank Ground Truth (Verb, Object) Error Top-2 Misprediction (seen/unseen) Prediction Most Frequent Co-occurring Component
 Count / Rate (%) Count / Rate (%) Verb-side Object-side
@@
 (‘pour’, ‘salt’) (seen) 7 / 9.5% w/ ‘water’ (25.2%) w/ ‘put’ (37.9%)
-Appendix C Additional Evidence of Object-Driven Shortcuts
+Appendix 0.D Additional evidence of object-driven shortcuts
 In this section, we provide further evidence demonstrating that existing ZS-CAR learning approaches are susceptible to object-driven shortcut learning.
 Failure case analysis of the existing ZS-CAR model.
-In Table 7 and Table 8, we investigate top-10 failure cases of the model trained with the existing ZS-CAR learning approach [16].
-The top-10 failure cases shown in the Table 7 constitute 9.5% of the misclassifications on unseen compositions for Sth-com [16], compared to 49% for EK100-com as shown in Table 8.
-For the analysis of Sth-com [16], we exclude ambiguous cases arising from annotation issues to obtain meaningful results.
+In Table 8 and Table 9, we investigate top-10 failure cases of the model trained with the state-of-the-art method [li2024c2c].
+The top-10 failure cases shown in the Table 8 constitute 9.5% of the misclassifications on unseen compositions for Sth-com [li2024c2c], compared to 49% for EK100-com as shown in Table 9.
+For the analysis of Sth-com [li2024c2c], we exclude ambiguous cases arising from annotation issues to obtain meaningful results.
 A prime example is the top-1 failure case, (‘Covering [something] with [something]’, ‘Paper’), where 66% of the errors are misclassified as ‘Pencil’.
@@
 These findings suggest that the model has failed to learn robust verb representations that generalize to unseen compositions, largely due to object-driven shortcut learning.
-Table 9: Additional ablation study of RCORE on the Sth-com.
-We provide the additional ablation study results on Sth-com [16].
-In every experiment, we use C2C [16] with CLIP-B/16 [28] as our baseline.
+Figure 13: Similarity analysis between original and reversed verb features.
+We plot the cosine similarity between the original and reversed verb features for the baseline (C2C [li2024c2c]) and RCORE with the CLIP [radford2021clip] backbone on the Sth-com [li2024c2c] dataset.
+The baseline maintains a high similarity (+0.92) during entire training, revealing limited temporal sensitivity.
+In contrast, the similarity becomes strongly negative for RCORE as training progresses, indicating improved temporal discriminative capability.
+Similarity analysis of verb features.
+In Figure 13, we track the cosine similarity between the original verb features $\mathbf{f}^{V}$ and the reversed verb features $\mathbf{f}^{V}_{\text{rev}}$ for the baseline [li2024c2c] over the course of training.
+The baseline [li2024c2c] maintains a high cosine similarity (+0.92+0.92), revealing its reliance on object-driven shortcuts rather than temporally grounded verb representations.
+Additionally, we plot the similarity of RCORE in Figure 13.
+RCORE drives this similarity to a strongly negative value (0.79-0.79), indicating that the model successfully distinguishes opposite temporal semantics, e.g.
+‘opening’ vs.
+‘closing’.
+Appendix 0.E Additional Results
+0.E.1 Macro Compositional Gap
+Table 10: Macro Compositional Gap results.
+We show Macro Compositional Gap ($\Delta_{\text{CG}}^{\text{macro}}$) results of C2C [li2024c2c] and RCORE on the Sth-com [li2024c2c] and EK100-com datasets.
+The best numbers are highlighted.
+Backbone Method Sth-com EK100-com
+Seen Unseen Seen Unseen
+CLIP AIM [yang2023aim] +1.18 -0.47 +0.95 -1.10
+C2C [li2024c2c] +0.78 +0.16 +1.56 -0.38
+RCORE +0.53 +0.64 -0.03 +0.11
+InternVideo2 LoRA [hu2022lora] 0.00 -0.16 +0.60 -0.53
+C2C [li2024c2c] +0.59 +0.14 +0.83 -0.38
+RCORE -0.33 +0.76 +0.02 +0.85
+In this section, we define Macro Compositional Gap ($\Delta_{\text{CG}}^{\text{macro}}$).
+$\Delta_{\text{CG}}^{\text{macro}}$ is the macro version of Compositional Gap ($\Delta_{\text{CG}}$), which averages $\Delta_{\text{CG}}$ per composition class, ensuring robustness to class imbalance.
+Let $\mathbb{Y}_{\text{test}}^{C}\subseteq\mathbb{Y}^{V}\times\mathbb{Y}^{O}$ denote the set of unique composition labels that appear in the test split.
+For each $\mathbf{y}^{C}=(\mathbf{y}^{V},\mathbf{y}^{O})\in\mathbb{Y}_{\text{test}}^{C}$, let $\mathrm{Acc}^{C}_{\mathbf{y}^{C}}$ be the top-1 composition accuracy restricted to samples of $\mathbf{y}^{C}$, and let $\mathrm{Acc}^{V}_{\mathbf{y}^{C}}$ and $\mathrm{Acc}^{O}_{\mathbf{y}^{C}}$ be the corresponding verb/object accuracies on the same samples of $\mathbf{y}^{C}$, computed from the joint top-1 composition predictions.
+Then we define
+$\Delta_{\text{CG}}^{\text{macro}}=\frac{1}{|\mathbb{Y}^{C}_{\text{test}}|}\sum_{\mathbf{y}^{C}\in\mathbb{Y}^{C}_{\text{test}}}\left[\text{Acc}^{C}_{\mathbf{y}^{C}}-\left(\text{Acc}^{V}_{\mathbf{y}^{C}}\times\text{Acc}^{O}_{\mathbf{y}^{C}}\right)\right].$ (14)
+In Table 10, we report $\Delta_{\text{CG}}^{\text{macro}}$ performances of baselines and RCORE on the Sth-com [li2024c2c] and EK100-com datasets.
+RCORE consistently achieves the highest $\Delta_{\text{CG}}^{\text{macro}}$ on unseen compositions across different backbones and datasets, demonstrating its robust compositional reasoning capabilities.
+0.E.2 Additional ablation results
+In Table 11, we provide the comprehensive ablation study results to examine each design choice of RCORE, using the Sth-com [li2024c2c] dataset.
+Effects of TORC with different temporal aggregation methods.
+In Table 11 (a), we demonstrate that the weak temporal order modeling caused by object-driven shortcut learning cannot be resolved merely by employing more sophisticated temporal aggregation methods.
+We compare the baseline’s temporal average pooling with attention pooling (Attn.
+pool).
+In attention pooling, a linear layer computes frame-wise attention weights from the input verb feature sequences, aggregating them via a weighted sum to obtain the final verb feature for classification.
+While attention pooling yields a slight improvement in verb accuracy, it fails to reduce the compositional gap ($\Delta_{CG}$) on unseen compositions.
+In contrast, our TORC achieves significant improvements in harmonic mean (H.M.) with both aggregation methods.
+Notably, the combination of temporal average pooling and TORC enables the model to overcome the negative compositional gap.
+Table 11: Additional ablation study of RCORE on Sth-com.
+We provide the additional ablation study results on Sth-com [li2024c2c].
+In every experiment, we use C2C [li2024c2c] with CLIP-B/16 [radford2021clip] as our baseline.
 We report performance on both seen and unseen compositions, along with their harmonic mean (H.M.).
-(a) Effects of TORC with different temporal aggregation.
-(b) Effects top-$K$ selection for Margin loss.
+The best numbers are highlighted.
+(a) Effects of TORC with different
+temporal aggregation.
+(b) Effects of input samples in CPR.
 Method TORC Verb Object Composition
-Seen Unseen Seen Unseen Seen ($\Delta_{\mathrm{CG}}$) Unseen ($\Delta_{\mathrm{CG}}$) H.M.
+V@S V@U O@S O@U Seen ($\Delta_{\mathrm{CG}}$) Unseen ($\Delta_{\mathrm{CG}}$) H.M.
 Average $\times$ 63.60 54.36 67.72 56.10 46.31 (+3.24) 30.08 (-0.42) 36.47
@@
 Attn. pool 65.51 56.50 67.80 55.00 48.07 (+3.65) 30.78 (-0.30) 37.53
-$K$ Verb Object Composition FCP
-Seen Unseen Seen Unseen Seen Unseen H.M.
-10 64.60 54.42 66.99 55.68 46.71 29.80 36.38 8.81
-100 63.63 53.93 67.58 55.66 46.43 29.18 35.84 9.89
-3451 62.75 53.63 66.89 55.20 45.32 29.16 35.49 8.02
-(c) Effects of hyperparameters of VOCAMix.
+Inputs Verb Object Composition
+V@S V@U O@S O@U Seen Unseen H.M.
+Original 62.54 52.97 56.90 57.05 37.43 31.41 34.16
+Mixup [zhang2018mixup] 62.37 52.56 63.50 59.20 41.62 31.45 35.83
+CPR w/o F.E. [ding2022fame] 63.23 54.29 65.15 56.49 43.57 30.86 36.13
+CPR w/ F.E. [ding2022fame] 64.17 54.55 63.60 58.70 42.74 33.05 37.28
+(c) Effects of hyperparameters of CPR.
+$p_{\text{CPR}}$ Verb Object Composition V@S V@U O@S O@U Seen Unseen H.M.
+0.8 64.66 52.89 65.06 58.77 45.20 31.14 36.88 0.9 64.70 55.50 64.73 57.69 44.75 32.60 37.72 0.95 63.95 54.05 63.97 59.52 43.09 33.22 37.51 1.0 64.95 55.50 61.49 57.20 41.82 32.84 36.79
 (d) Effects of hyperparameters of TORC.
-$p_{aug}$ Verb Object Composition
-Seen Unseen Seen Unseen Seen Unseen H.M.
-0.0 63.60 54.36 67.72 56.10 46.31 30.08 36.47
-0.3 64.67 54.83 68.18 55.37 47.77 29.93 36.80
-0.5 64.26 54.65 67.32 56.61 46.86 30.35 36.84
-1.0 64.02 53.80 67.04 55.26 46.30 29.07 35.72
-$\gamma$ scaling scaling Verb Object Composition
-start epoch end epoch Seen Unseen Seen Unseen Seen Unseen H.M.
-0.0 0 0 63.60 54.36 67.72 56.10 46.31 30.08 36.47
-0.5 1 1 65.59 55.63 68.49 54.80 48.33 30.19 37.17
-1.0 1 1 64.88 55.81 67.21 55.12 47.06 30.69 37.15
-1.0 5 5 64.93 56.03 67.67 55.81 47.20 31.38 37.70
-1.0 5 10 65.65 56.80 68.59 55.19 48.65 31.35 38.13
-(e) Effects of hyperparameters of Margin loss.
-$\delta$ scaling scaling Verb Object Composition
-start epoch end epoch Seen Unseen Seen Unseen Seen Unseen H.M.
-0.3 10 10 63.09 52.94 67.90 54.86 46.31 28.48 35.27
-0.3 15 15 63.96 53.70 67.70 55.68 46.57 29.28 35.96
-0.3 15 20 64.01 53.76 67.53 55.61 46.87 29.42 36.15
-0.5 15 20 63.85 53.61 67.90 56.32 46.79 29.49 36.18
-0.8 15 20 64.60 54.42 66.99 55.68 46.71 29.80 36.38
-Appendix D Additional Results
-D.1 Comprehensive ablation results
-In Table 9, we provide the comprehensive ablation study results to examine each design choice of RCORE, using the Sth-com [16] dataset.
-Effects of TORC loss with different temporal aggregation methods.
-In Table 9 (d), we demonstrate that the weak temporal order modeling caused by object-driven shortcut learning cannot be resolved merely by employing more sophisticated temporal aggregation methods.
-We compare the baseline’s temporal average pooling with attention pooling (Attn.
-pool).
-In attention pooling, a linear layer computes frame-wise attention weights from the input verb feature sequences, aggregating them via a weighted sum to obtain the final verb feature for classification.
-While attention pooling yields a slight improvement in verb accuracy, it fails to reduce the compositional gap ($\Delta_{CG}$) on unseen compositions.
-In contrast, our TORC loss achieves significant improvements in H.M.
-with both aggregation methods.
-Notably, the combination of temporal average pooling and TORC loss enables the model to overcome the negative compositional gap.
-Effects of top-$K$ selection for Margin loss.
-In Table 9 (e), we investigate the impact of the hyperparameter $K$, which determines the number of top-ranked logits to which the Margin loss is applied.
-We find that applying the margin selectively to the frequent-but-incorrect logits within the top-10 ($K=10$) reports the highest H.M., outperforming the approach of applying it to all seen composition logits ($K=3451$) (36.4% vs.
-35.5%).
-This suggests that penalizing only the most confusing candidates—specifically those that frequently co-occur with the input components—effectively mitigates co-occurrence bias without compromising the model’s overall performance.
-For instance, given an input like ‘(Pretending to tear, Paper)’, our intention is to penalize plausible but incorrect distractors such as ‘(Tearing, Paper)’ or ‘(Folding, Paper)’, while avoiding penalties on irrelevant classes like ‘(Throwing, Paper)’ that are already ranked low.
+$\gamma$ scaling scaling Verb Object Composition start epoch end epoch V@S V@U O@S O@U Seen Unseen H.M.
+0.5 1 1 65.59 55.63 68.49 54.80 48.33 30.19 37.17 1.0 1 1 64.88 55.81 67.21 55.12 47.06 30.69 37.15 1.0 5 5 64.93 56.03 67.67 55.81 47.20 31.38 37.70 1.0 5 10 65.65 56.80 68.59 55.19 48.65 31.35 38.13
+(e) Effects of hyperparameters of $L_{\text{CPR}}$.
+$m$ $\delta$ scaling scaling Verb Object Composition start epoch end epoch V@S V@U O@S O@U Seen Unseen H.M.
+0.5 0.1 15 15 67.06 58.13 63.67 58.66 45.00 35.16 39.48 0.5 0.1 15 20 66.16 58.78 63.75 57.96 44.65 34.89 39.17 0.5 0.5 15 15 66.16 57.03 62.75 57.56 43.55 33.77 38.04 0.5 0.5 15 20 67.04 56.77 61.74 57.56 43.37 34.30 38.31 1.0 0.1 15 15 66.31 58.56 63.77 57.99 44.47 34.64 38.94 1.0 0.1 15 20 66.88 58.24 63.88 57.01 44.90 34.21 38.83
+Effects of input samples in CPR.
+In Table 11 (b), we validate the effects of input sample types in CPR.
+We argue that conventional image-based augmentations are insufficient for representing newly created unseen composition labels.
+For example, Mixup [zhang2018mixup] blends verbs and objects concurrently, failing to preserve the temporal difference information that is essential for verb recognition.
+As a result, these methods yield a lower verb@unseen-comp than the baseline (52.6% vs.
+53.0%).
+In contrast, our CPR adopts a static frame mixing strategy, aiming to generate novel compositions by mixing only object information while preserving the verbs, shows modest performance gains over the baseline for both verb and object accuracy on unseen compositions, resulting in the best H.M.
+Furthermore, we observe that mixing a static frame into the estimated foreground region shows additional performance boost (+0.2 points in the H.M.) over applying it to the entire background, as it creates more realistic samples.
+We denote the latter as ‘CPR w/o F.E.’, indicating ‘without foreground estimation’.
+Table 12: Ablation study on EK100-com.
+We provide the results of the ablation study on EK100-com presented in the main paper.
+In every experiment, we use C2C [li2024c2c] with CLIP-B/16 [radford2021clip] as our baseline.
+We report performance on both seen and unseen compositions, along with their harmonic mean (H.M.).
+(a) Effects of CPR and TORC.
+CPR TORC Verb Object Composition) V@S V@U O@S O@U Seen Unseen H.M.
+66.19 49.71 56.61 47.48 42.72 22.38 29.38 66.06 51.28 54.23 50.56 38.93 27.10 31.96 65.90 52.31 57.28 46.74 43.48 23.67 30.65 67.52 54.08 54.00 49.04 40.13 27.40 32.56
+(b) Effects of each TORC loss term.
+$L_{\cos}$ $L_{\text{ent}}$ Verb Object Composition V@S V@U O@S O@U Seen Unseen H.M.
+66.19 49.71 56.61 47.48 42.72 22.38 29.38 65.43 52.11 56.54 47.44 42.67 24.24 30.92 66.00 51.06 56.64 47.37 42.82 22.53 29.53 65.90 52.31 57.28 46.74 43.48 23.67 30.65
+(c) Effects of $L_{\text{CPR}}$ in CPR. $L_{\text{CPR}}$ Verb Object Composition V@S V@U O@S O@U Seen Unseen H.M. 66.35 50.79 54.72 50.16 39.25 26.60 31.71 66.06 51.28 54.23 50.56 38.93 27.10 31.96
+Effects of hyperparameters of RCORE.
+In Table 11 (c), we show the effect of $p_{\text{CPR}}$, the per-batch probability for CPR application, revealing a trade-off between seen and unseen composition accuracies.
+Decreasing $p_{\text{CPR}}$ slightly from 1.0 (e.g.
+0.95 or 0.9) yields optimal unseen accuracy improvements with minimal drops in seen performance, whereas lowering it further (e.g.
+$p_{\text{CPR}}=0.8$) degrades the gains on unseen pairs.
+Consequently, we tune $p_{\text{CPR}}$ for each dataset and backbone based on the best validation H.M.
+of seen and unseen accuracies, as shown in Table 7.
+In Table 11 (d), we observe that introducing the explicit temporal order modeling task via TORC starting from 5 epoch, after initial training, is effective.
+In particular, the strategy of linearly increasing the loss weight from epoch 5 to 10 is the most effective.
+Finally, in Table 11 (e), we analyze the impact of the loss weights of $L_{\text{CPR}}$ in CPR.
+Increasing the margin $m$ or loss weight $\delta$ degrades performance, suggesting that excessive training penalties are detrimental.
+We find $m=0.5$ and $\delta=0.1$ to be optimal.
+Contrary to the results in Table 11 (d), maintaining a constant low loss weight ($\delta=0.1$) for $L_{\text{CPR}}$ yields better performance than applying progressive weight scaling.
+0.E.3 Ablation studies on EK100-com
+In Table 12, we provide the ablation study results conducted on the EK100-com dataset, complementing the results discussed with the Sth-com [li2024c2c] dataset in the main paper.
+In Table 12 (a), we demonstrate that while CPR and TORC are individually effective, integrating both components yields the best performance on EK100-com, outperforming the baseline in H.M.
+by 3.2 points.
+In Table 12 (b), using only $L_{cos}$ in TORC yields the highest H.M.
+on EK100-com.
+In Table 12 (c), surpressing co-occurrence priors with $L_{cos}$ improves H.M.
+by 0.3 points, consistent with the results on Sth-com [li2024c2c].
+0.E.4 Results with validation-set-tuned bias calibration
+Table 13: Sth-com results with bias calibration.
+We show the top-1 verb, object, and composition classification accuracies (%) on the Sth-com [li2024c2c] test set.
+We tune the bias term for unseen composition predictions on the validation set and apply the selected bias (‘Best Bias’) to the test set.
+We report performance on both seen and unseen compositions, along with their harmonic mean (H.M.).
+The best numbers are highlighted.
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+In this section, we provide the open-world biased results with validation-set-tuned calibration.
+Bias calibration is a standard technique in compositional zero-shot learning [purushwalkam2019taskdriven, mancini2021compcos, kim2023hierarchical, nayak2023csp, wu2025cond, li2024context] and ZS-CAR [li2024c2c, jiang2025dhd, jung2025crr, ye2025logic].
+It compensates for the inherent issue where logits for unseen compositions—not optimized during training—are naturally lower than those for seen compositions.
+Specifically, it adds a scalar bias term exclusively to the unseen composition logits.
+Previous ZS-CAR works [li2024c2c, jiang2025dhd, jung2025crr, ye2025logic] often tuned this bias using test-set ground truth under a closed-world setting to report the ‘best’ seen and unseen accuracies.
+We argue that this approach is highly unrealistic.
+Instead, we adopt an open-world validation-set-tuned bias calibration.
+We tune the bias on the validation set using open-world logits and then apply it to the test set.
+This approach remains applicable to real-world scenarios while effectively improving unseen composition performance as a post-processing step.
+In Table 13, we present the calibrated results of the baselines and RCORE on the Sth-com [li2024c2c] dataset.
+Notably, the calibrated performance of RCORE is identical to its original unbiased performance.
+This indicates that our training label space expansion strategy via our proposed CPR is effective.
+Unlike baseline methods that overfit solely to seen compositions, it successfully provides meaningful supervision for unseen compositions during training.
+Furthermore, the unbiased performance of RCORE consistently outperforms the best calibrated performance of all baselines.
+This demonstrates that RCORE is highly suitable for real-world applications, delivering strong ZS-CAR performance without the need for any post-processing.
+Table 14: Sth-com results with InternVideo2-1B.
+We show the top-1 verb, object, and composition classification accuracies (%) on the Sth-com [li2024c2c] and EK100-com datasets with the InternVideo2-1B [wang2024iv2] backbone.
+We report performance on both seen and unseen compositions, along with their harmonic mean (H.M.).
+The best numbers are highlighted.
+Sth-com results
+Method Verb Object Composition
+@Seen Comp @Unseen Comp H.M. @Seen Comp @Unseen Comp H.M. Seen ($\Delta_{\text{CG}}$) Unseen ($\Delta_{\text{CG}}$) H.M.
+C2C [li2024c2c] 67.50 58.21 62.51 70.58 61.62 65.80 50.23 (+2.59) 34.96 (-0.91) 41.23
+RCORE 67.49 60.47 63.79 68.83 63.63 66.13 48.27 (+1.82) 38.35 (-0.13) 42.74
+EK100-com results
+Method Verb Object Composition
+@Seen Comp @Unseen Comp H.M. @Seen Comp @Unseen Comp H.M. Seen ($\Delta_{\text{CG}}$) Unseen ($\Delta_{\text{CG}}$) H.M.
+C2C [li2024c2c] 66.95 52.22 58.67 60.57 50.81 55.26 45.91 (+5.36) 26.66 (+0.13) 33.73
+RCORE 66.36 52.88 58.86 56.83 52.10 54.36 41.86 (+4.16) 28.79 (+1.24) 34.12
 (a) Our reconstructed Temporal/Static splits
-(b) Temporal/Static splits from Sevilla et al. [31]
+(b) Temporal/Static splits from Sevilla et al. [laura2021temporal]
 Figure 14: Performances on Temporal/Static split of Sth-com.
-We evaluate the models on Sth-com [16] using both (a) our reconstructed splits and (b) the splits from Sevilla et al [31].
+We evaluate the models on Sth-com [li2024c2c] using both (a) our reconstructed splits and (b) the splits from Sevilla et al [laura2021temporal].
 We utilize both original and temporally shuffled inputs to assess the model’s temporal modeling capability and its reliance on static cues.
 A larger performance gap between original and shuffled inputs indicates that the model predicts verbs more based on temporal dynamics rather than on static cues.
-Effects of hyperparameters of RCORE.
-In Table 9 (f), we demonstrate that applying augmentation with a probability of 0.5 within a batch is the most effective, achieving a 0.4 point improvement in H.M.
-over the baseline.
-In Table 9 (g), we observe that introducing the explicit temporal order modeling task via TORC starting from 5 epoch, after initial training, is effective.
-In particular, the strategy of linearly increasing the loss weight from epoch 5 to 10 is the most effective, resulting in 1.7 point boost in H.M.
-compared to the baseline.
-Finally, in Table 9 (h), we analyze the impact of scaling the weight of our Margin loss, revealing a trade-off between improved unseen composition accuracy and decreased seen composition accuracy.
-We observe that applying Margin loss starting from epoch 15, the point at which the model begins to overfit the training data as shown in Figure 3 of the main paper, is more effective than starting from the earlier epoch.
-We also find that linearly scaling the loss weight over 5 epochs is the most effective strategy to maintain the model’s performance.
-D.2 Analysis of confidence distribution in composition classification
-(a) Sth-com results
-(b) EK100-com results
-Figure 11: Conditional modeling overfits to co-occurrence statistics.
-We track how the composition prediction confidence of C2C [16] evolves during training both on Sth-com [16] and EK100-com datasets.
-As training progresses, C2C [16] increasingly ignores input evidence and misclassifies unseen compositions as seen ones.
-In Figure 11, we track how the composition prediction confidence of C2C [16] evolves during training both on Sth-com [16] and EK100-com datasets.
-Using the test set of both datasets, we compare two cases of the model’s confidence values: Seen $\to$ Seen and Unseen $\to$ Seen.
-Given the seen/unseen composition set $\mathbb{Y}_{\text{seen}}$$\mathbb{Y}_{\text{unseen}}$ and the ground-truth/model-predicted composition label $\mathbf{y}$$\hat{\mathbf{y}}$, Seen $\to$ Seen refers to the case where $\mathbf{y}_{i}\in\mathbb{Y}_{\text{seen}}\land\hat{\mathbf{y}}_{i}\in\mathbb{Y}_{\text{seen}}$ and Unseen $\to$ Seen refers to the case where $\mathbf{y}_{i}\in\mathbb{Y}_{\text{unseen}}\land\hat{\mathbf{y}}_{i}\in\mathbb{Y}_{\text{seen}}$.
-Across both datasets, we observe that the model increasingly classifies unseen compositions as seen compositions as training progresses.
-The observation suggests that, without careful design, models may be prone to ignore input cues and instead over-rely on verb–object co-occurrence statistics present in the training data.
-D.3 Temporal/Static split of Sth-com
-Table 10: Ablation study on the EK100-com.
-We provide the results of the ablation study on EK100-com presented in the main paper.
-In every experiment, we use C2C [16] with CLIP-B/16 [28] as our baseline.
-We report performance on both seen and unseen compositions, along with their harmonic mean (H.M.).
-(a) Effects of VOCAMix and TORC.
-VOCAMix TORC Verb Object Composition (OW) Seen Unseen Seen Unseen Seen Unseen H.M.
-66.19 49.71 56.61 47.48 42.72 22.38 29.38 65.55 52.15 57.09 46.87 43.01 23.81 30.65 65.90 52.31 57.28 46.74 43.48 23.67 30.65 66.21 53.51 56.97 47.52 43.53 25.02 31.78
-(b) Effects of each TORC loss term.
-$L_{\cos}$ $L_{\text{ent}}$ Verb Object Composition (OW) Seen Unseen Seen Unseen Seen Unseen H.M.
-66.19 49.71 56.61 47.48 42.72 22.38 29.38 65.43 52.11 56.54 47.44 42.67 24.24 30.92 66.00 51.06 56.64 47.37 42.82 22.53 29.53 65.90 52.31 57.28 46.74 43.48 23.67 30.65
+0.E.5 Results with large video-pretrained VLM backbone
+We compare RCORE and the baseline (C2C [li2024c2c]) on Sth-com [li2024c2c] and EK100-com using the InternVideo2-1B [wang2024iv2] backbone, whose 1B-parameter vision encoder is approximately $11\times$ larger than the 87M-parameter vision encoders used in the main paper.
+In Table 14, RCORE yields consistent improvements over the baseline on Sth-com [li2024c2c] and EK100-com, boosting unseen composition accuracy (+3.4p., +2.1p.), unseen $\Delta_{\text{CG}}$ (+0.8p., +1.1p.), and H.M.
+(+1.5p., +0.4p.), respectively.
+These results demonstrate that its powerful representation learning capability is not constrained by the backbone’s parameter size.
+0.E.6 Temporal/Static split of Sth-com
 We present extended results corresponding to the experiment shown in Figure 5 (c) of the main paper, including all versions of the Temporal/Static split.
-Inspired by prior work [31, 39], we rigorously evaluate the temporal modeling capabilities of both the baseline (C2C [16]) and our RCORE by partitioning the Sth-com [16] dataset into Temporal and Static splits.
+Inspired by prior work [laura2021temporal, yun2022time], we rigorously evaluate the temporal modeling capabilities by partitioning the Sth-com [li2024c2c] dataset into Temporal and Static splits, as shown in Figure 14.
+We evaluate both the baseline (C2C [li2024c2c]) and RCORE on the two types of splits, (i) our reconstructed splits and (ii) the splits proposed by Sevilla et al.
+[laura2021temporal].
+In all cases, RCORE exhibits lower performance on shuffled inputs but outperforms the baseline on original inputs.
+This indicates that RCORE predicts verbs without relying on static cues, leading to robust improvements in verb performance regardless of the split.
+In this paragraph, we describe the differences between two Temporal/Static splits.
 Sevilla et al.
-[31] originally defined these splits based on cognitive science experiments on Something-Something V2 [12] verb classes: a Temporal split (18 verbs) where temporal information matters, and a Static split (18 verbs) where such information is redundant.
-However, noting that this Temporal split includes verbs requiring complex temporal reasoning such as ‘Folding’, we reconfigure the splits based on verb semantics.
+[laura2021temporal] originally defined these splits based on cognitive science experiments on Something-Something V2 [goyal2017something] verb classes: a Temporal split (18 verbs) where temporal information matters, and a Static split (18 verbs) where such information is redundant.
+However, noting that this Static split includes verbs requiring complex temporal reasoning such as ‘Folding’, we reconfigure the splits based on verb semantics.
 Specifically, we categorized verbs based on this criterion: those requiring temporal information for discrimination are assigned to the Temporal split (62 verbs), while verbs distinguishable by momentary keyframes are assigned to the Static split (80 verbs).
@@
 To balance the sample sizes between the Temporal and Static splits, we excluded a few verbs with the fewest samples from the Temporal split.
-We then utilize both original and temporally shuffled inputs to assess the model’s temporal modeling capability and its reliance on static cues.
-In Figure 10, we present the results for the Temporal/Static splits across seen/unseen compositions.
-We also provide results applying the original Temporal/Static verb split [31] to Sth-com [16].
-In all cases, RCORE exhibits lower performance on shuffled inputs but outperforms the baseline on original inputs.
-This indicates that RCORE predicts verbs without relying on static cues, leading to robust improvements in verb performance regardless of the split.
-D.4 Ablation studies on EK100-com
-In Table 10, we provide the ablation study results conducted on the EK100-com dataset, complementing the results discussed with the Sth-com [16] dataset in the main paper.
-In Table 10 (a), we demonstrate that our framework, integrating VOCAMix for sparsity resolution and TORC for explicit temporal modeling, outperforms the baseline on EK100-com, achieving 3.8 points higher verb accuracy in unseen compositions than the baseline.
-In Table 10 (b), using only $L_{cos}$ in TORC yields the highest harmonic mean of composition accuracy on EK100-com.